### v1

# Information Sifting Funnel: Privacy-preserving Collaborative Inference Against Model Inversion Attacks

Abstract.
The complexity of neural networks and inference tasks, combined with the demand for computational efficiency and real-time feedback, poses significant challenges for resource-constrained edge devices. Collaborative inference addresses this by delegating shallow feature extraction to edge devices and uploading feature to the cloud for subsequent inference tasks, reducing computational pressure. However, the transmitted features remain vulnerable to model inversion attacks (MIAs), which can reconstruct original input data. Existing defenses, including perturbation and information bottleneck techniques, provide explainable protection but are limited by the lack of standardized criteria for evaluating MIA difficulty, challenges in mutual information estimation, and trade-offs among usability, privacy, and deployability.
To overcome these limitations, we establish the first criterion for assessing the difficulty of MIAs in collaborative inference, supported by theoretical analysis of existing attacks and defenses and validated through experiments using the Mutual Information Neural Estimator (MINE). Building on these findings, we propose SiftFunnel, a privacy-preserving framework for collaborative inference. Specifically, we first train the edge model with both linear and non-linear correlation constraints to limit redundant information in transmitted features, enhancing privacy protection. Moreover, label smoothing and a cloud-based upsampling module are incorporated to balance usability and privacy. To further support deployability, the edge model is designed with a funnel-shaped structure and integrates attention mechanisms, preserving both privacy and usability. Extensive experiments show that SiftFunnel outperforms state-of-the-art defenses against MIAs, achieving superior privacy protection with less than 3% accuracy loss and an optimal trade-off among usability, privacy, and practicality.
Collaborative inference, model inversion attacks, deep learning security and privacy, information theory
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference on Computer and Communications Security; October 13–17, 2025; xx, xxisbn: 978-1-4503-XXXX-X/25/10ccs: Security and privacy Domain-specific security and privacy architectures
1. Introduction
The rapid advancement of deep learning has led to increasingly complex and powerful neural network models, enabling remarkable performance in tasks such as image detection and recognition, speech processing, and natural language understanding. However, these advancements also pose significant challenges for edge computing devices (Huckelberry et al., 2024). In fields such as intelligent transportation, autonomous driving, facial recognition, IoT applications involving cameras and sensors, and remote healthcare monitoring, edge devices must process data instantly while handling growing computational demands. The limited resources of edge devices make it difficult for them to efficiently perform complex deep learning inference tasks independently (Shlezinger and Bajić, 2022).
Figure 1. Collaborative inference and the threat of model inversion attacks.
[…]
Collaborative inference and the threat of model inversion attacks.
Cloud computing enables inference by uploading data to cloud-based models, offering significant computational power. However, it faces two key limitations: bandwidth and network latency hinder its ability to meet real-time application requirements, and uploading raw data raises privacy concerns, especially in sensitive scenarios. Collaborative inference (CI) (Shlezinger et al., 2021) has emerged as a viable solution to these challenges. By partitioning a neural network into $f_{\text{edge}}$ and $f_{\text{cloud}}$, the edge device processes raw inputs into feature representations via $f_{\text{edge}}$, which are then uploaded to the cloud for further inference by $f_{\text{cloud}}$. This approach balances real-time performance and computational efficiency. In recent years, this technique has garnered attention in practical applications, such as unmanned aerial vehicles (UAVs) (Qu et al., 2023) and IoT systems (Shlezinger and Bajić, 2022), as well as in solutions like Apple’s private cloud computing (PCC) (Engineering and Architecture, 2024). With advancements in 5G, IoT, and AI, collaborative inference holds promise for broader adoption and future impact.
However, recent studies have revealed significant privacy risks in collaborative inference, where input data on edge devices can be reconstructed through model inversion attacks (MIAs) (He et al., 2019; Yang et al., 2022; Yin et al., 2023), as shown in Figure 1. MIAs are fine-grained privacy attacks that leverage the output information of AI models and deep learning generation techniques to reconstruct input data (He et al., 2019; Yang et al., 2019), recover representative training data (Zhang et al., 2020; Liu et al., 2024), or expose sensitive information (Fredrikson et al., 2014; Mehnaz et al., 2022). These risks are particularly severe in collaborative inference scenarios. The root cause is that, compared to centralized model deployment, edge devices typically employ shallow neural networks to compute and upload feature representations. These features contain far more information than the final inference results, enabling adversaries to reconstruct the input data (Azizian and Bajić, 2024). For instance, in facial recognition applications, while adversaries in centralized deployments might only recover training data through the model’s API (Yuan et al., 2023; Struppek et al., 2022), those in collaborative inference can directly reconstruct the raw facial data collected by edge devices from the transmitted features.
[…]
Firstly, existing defense techniques in collaborative inference lack a thorough and unified analysis of the fundamental nature of MIA. Current mainstream defense methods can be categorized into three types based on their technical approaches: perturbation-based (He et al., 2020; Wang et al., 2022), information bottleneck (IB)-based (Wang et al., 2024; Peng et al., 2022), and neural network depth & information bottleneck (NND & IB)-based defense strategies (Azizian and Bajić, 2024; Ding et al., 2024). Perturbation-based strategies aim to increase the difficulty of attacks by introducing perturbations to the gradients or output information of the trained model. IB-based strategies focus on approximating the reduction of mutual information $I(x,z)$ between the input data $x$ and the feature $z$, thus minimizing redundant information about $x$ in $z$ (Wang et al., 2021). NND & IB-based strategies leverage the observation that increasing the depth of the edge network can induce an information bottleneck effect, thus enhancing defense. These methods provide theoretically interpretable defense mechanisms, with the introduction of information theory concepts offering notable contributions to defense design. However, existing work lacks a comprehensive understanding of MIA, relies on weak assumptions about adversarial capabilities, and fails to establish a clear criterion for evaluating the complexity of implementing MIA. As a result, defenses against MIAs still have room for improvement, including addressing potential future advancements.
Secondly, the trade-off between usability, privacy, and deployability in edge models remains a challenging and underexplored area in current research. Encryption-based defense strategies receive limited attention in collaborative inference due to resource constraints and the need for multitask processing in real time (Mireshghallah et al., 2021). Additionally, encryption does not guarantee complete data security, as modern side-channel attacks leveraging deep learning can recover encrypted data (Hu et al., 2023). The three main defense strategies discussed above also face issues of imbalance. Perturbation-based methods often compromise usability, whereas IB-based defenses aim to reduce mutual information $I(x,z)$, providing theoretical protection against simpler attacks such as MLE-based MIA. However, accurately computing $I(x,z)$ is challenging, as it relies on approximations through sufficient data and specific algorithms, making it less effective against more advanced attacks such as Gen-based MIA. Adversarial training has been introduced as a component to address the aforementioned issues, but its strong specificity and limited interpretability make it difficult to prevent future advanced attacks while significantly increasing the difficulty of training edge models. The NND & IB-based approach aims to balance privacy and usability by increasing model depth, but this approach significantly raises the resource demands on edge devices, including memory, storage, and bandwidth, thereby reducing deployability. As a result, such methods are often less favored in practical applications.
Thus, there is a pressing need for a novel approach grounded in an in-depth analysis of attack theories to provide interpretable and effective privacy protection while maintaining the usability and deployability of edge models. This need gives rise to three key technical challenges. Challenge 1: How to establish an experimentally validated criterion for assessing MIA difficulty in collaborative inference? Prior studies have interpreted MIA from an information-theoretic perspective, but challenges in computing mutual information and conducting an in-depth analysis of attack principles have hindered the establishment of a verifiable criterion, which is essential for guiding defense design and evaluating methods. Challenge 2: How to effectively and interpretably remove redundant information from transmitted features while preserving edge model usability? Although information bottleneck theory provides an interpretable defensive effect, mutual information is challenging to calculate, and current approximation methods are suboptimal. Challenge 3: How to balance usability and privacy while ensuring practical deployability? NND & IB-based methods significantly increase computational overhead on edge devices, making them challenging for practical application. Moreover, when applied to edge models that perform only simple feature extraction, these methods also struggle to resist advanced Gen-based MIAs.
To overcome the aforementioned challenges, we established a criterion $D_{mia}$ for evaluating the difficulty of MIAs in CI and proposed a lightweight privacy-preserving strategy called Information Sifting Funnel (SiftFunnel), incorporating three key features. (1) SiftFunnel provides interpretable defense principles based on $D_{mia}$ We identified factors influencing attack difficulty in collaborative inference, including $I(x,z)$ and the effective information mean $\delta(z)$, i.e., the average number of non-zero elements in $z$. By analyzing these factors, we clarified defense design directions and developed a framework to address potential future threats. Using the Mutual Information Neural Estimator (MINE) (Belghazi et al., 2018), we quantified the changes in $I(x,z)$ at different network depths, validating the criterion. (2) SiftFunnel effectively reduces redundant information in transmitted features while preserving edge model usability. Guided by the proposed criterion, we enhanced the model’s loss function to include distance-correlation and Pearson-correlation measures, alongside an $l_{1}$-norm constraint on $z$. These refinements reduce $H(z)$, $\delta(z)$ while increasing $H(x|z)$. To further filter redundant information, we integrated Squeeze-and-Excitation (SE) Modules (Hu et al., 2018), Triple Attention Mechanisms (Misra et al., 2021), and Convolutional Block Attention Modules (CBAM) (Woo et al., 2018) into the edge model to focus on task-relevant features. Usability was preserved through a combination of label-smoothing and KL-divergence loss. (3) SiftFunnel balances usability and privacy while light-weighting the edge model. A funnel-shaped edge model was designed to progressively reduce transmitted channels, improving efficiency and reducing redundant information while minimizing memory and storage demands. To maintain usability, an upsampling module was added on the cloud side.
Based on prior evaluation metrics, this paper incorporates $I(x,z)$, quantified using MINE, and $\delta(z)$ as key metrics to validate the theoretical analysis and assess the effectiveness of defense methods. To evaluate deployability, we consider the edge model’s parameter count, memory usage, storage consumption, and inference latency. Through extensive experiments, we demonstrate the validity of the proposed difficulty criterion and the effectiveness of SiftFunnel in balancing usability, deployability, and data privacy compared to existing methods.
Our contributions are summarized as follows.
  • We establish the first difficulty criterion for implementing MIA in collaborative inference, providing guidance for defense design and preventing future advanced MIAs.
  • We propose SiftFunnel, a privacy-preserving scheme for collaborative inference that constrains redundant transmitted features through linear and nonlinear correlation, ensuring model usability with label smoothing and loss function regularization.
  • SiftFunnel adopts a funnel-shaped structure with integrated attention mechanisms to balance usability and privacy while maintaining the deployability of edge models.
  • We conduct systematic experiments validating the proposed criterion and demonstrating SiftFunnel’s advantages in balancing usability, privacy, and deployability.
2. Background and Related Work
2.1. Collaborative inference
The rapid development of artificial intelligence has enabled deep learning models to achieve exceptional performance in image processing and other tasks. However, the computational demands of these models often exceed the capabilities of resource-constrained edge devices, which face limitations in processing power, storage, and energy efficiency. To address these challenges, collaborative inference (Shlezinger et al., 2021) has emerged as an effective solution. This approach divides the inference process into two stages: one part is performed locally on the edge device (which may consist of a single device or a chain of interconnected devices), while the remaining computation is uploaded to the cloud. Formally, this can be expressed as
(1) $f(x)=f_{e}(x)\circ f_{c}(f_{e}(x))$
where $f$ is the full trained model, $x$ is the input, $f_{e}$ represents the model deployed on the edge, and $f_{c}$ corresponds to the portion deployed in the cloud.
Collaborative inference typically begins with training $f$ on centralized datasets in the cloud, followed by partitioning and deploying the $f_{e}$ to the edge. Alternatively, methodologies such as Split Learning  (Zhu et al., 2023) enable edge devices to participate in model training locally. Regardless of the training approach, the inference process remains consistent. As depicted in Figure 1, the edge-deployed model first computes an intermediate representation $feature=f_{e}(x)$, which is then transmitted to the cloud for completion of the inference process $result=f_{c}(feature)$. The final output is either sent back to the edge device or retained in the cloud for further analysis.
Compared to the traditional approach of uploading raw data to the cloud for inference, collaborative inference offers distinct advantages. First, it reduces communication overhead by allowing the transmitted features to be processed, potentially making them smaller in size than the original data. Second, it improves privacy protection by limiting the exposure of sensitive information by transmitting processed features instead of raw data, such as physiological signals or facial images. These advantages make collaborative inference a suitable solution for applications in healthcare, facial recognition, and industrial IoT, where efficiency and privacy are critical requirements.
However, existing research highlights that collaborative inference models are susceptible to Model Inversion Attacks (MIA), where adversaries attempt to reconstruct the original input data from transmitted feature. Notably, in edge-based inference, computations can be distributed across multiple devices in a chained architecture (He et al., 2019), expressed as $f=f_{e1}\circ f_{e2}...\circ f_{c}$. Studies show that deeper neural networks exhibit greater resistance to data reconstruction, making devices positioned closer to the cloud less vulnerable. In contrast, the initial edge device, which typically employs a shallow neural network for feature computation and transmission, generates representations that retain more information from the original input, increasing its susceptibility to MIA. Given this vulnerability, this paper focuses on the MIA threat to the initial edge device, which uses a shallow neural network while the rest resides in the cloud. This common scenario in real-world applications poses a significant security risk due to the heightened likelihood of input data reconstruction, underscoring the need to address this challenge.
2.2. Model inversion attack
Model Inversion Attack (MIA) is a fine-grained privacy attack that uses the output of a target model and adversarial knowledge, such as white-box access or training data distribution, to reconstruct training data (Zhang et al., 2020), sensitive attributes (Mehnaz et al., 2022), or input data (Yang et al., 2019). Initially proposed by Fredrikson et al. (Fredrikson et al., 2015), MIA was applied to linear regression models, successfully reconstructing genomic data from warfarin dosage predictions and patient attributes, thereby revealing the privacy risks. Subsequent studies (Fredrikson et al., 2015) extended this concept to deep learning, showing that training data from deep neural networks could also be reconstructed.
Current MIA techniques in collaborative inference can be broadly categorized into Maximum Likelihood Estimation-based (MLE-based) MIAs and Generative Model-based (Gen-based) MIAs, which are further analyzed in the following sections.
MLE-based MIAs. MLE-based MIA originated from the 2015 study by Fredrikson et al. (Fredrikson et al., 2015), which reconstructed training data from a multilayer perceptron (MLP). This method minimizes the difference between target and output class confidence scores as a loss function and employs gradient descent to optimize the input, enabling the reconstruction of the original training data. Subsequent research enhanced this approach by incorporating techniques like Total Variation (TV) (Rudin et al., 1992) loss as a regularization term to improve optimization, with these improved methods later referred to as rMLE-based MIA by He et al. (He et al., 2019) In collaborative inference scenarios, MLE-based methods are particularly effective for reconstructing data from shallow neural networks deployed on edge devices. Unlike the prediction vectors produced by full neural networks, the richer feature representations and lower complexity of shallow networks provide adversaries with more detailed and easily optimizable gradient information.
Gen-based MIAs. This method originated from the 2019 work of Yang et al. (Yang et al., 2019), who reconstructed input data in facial recognition systems by using generative deep learning models $g$ to approximate the inverse mapping function of the target model $f^{-1}$, such that $g\approx f^{-1}$. By inputting prediction vectors into $g$, they successfully generated facial input images. Building on this approach, He et al. (He et al., 2019) extended it to collaborative inference scenarios, demonstrating that input data from edge models is more vulnerable to reconstruction, thereby highlighting the risks associated with collaborative inference. Unlike MLE-based methods, Gen-based MIA does not require white-box access to the edge model. Instead, it trains the attack model by calculating losses based on transmitted features, enabling it to approximate the inverse mapping of non-linear computations. Furthermore, as edge models in collaborative inference are typically shallow neural networks, their outputs often retain redundant information from the original input and features, maintaining a strong correlation between inputs and outputs and facilitating the reconstruction process.
2.3. Related defense work
To protect the input data collected by edge devices in collaborative inference from potential leakage, He et al. (He et al., 2019, 2020) highlighted the necessity of developing defense strategies while identifying these threats. Building on a detailed analysis of attack methods, subsequent defense techniques have gradually evolved in recent years. These defenses can be broadly categorized into three types: perturbation-based defenses, information bottleneck-based defenses, and defenses combining neural network depth with information bottleneck principles. The following sections provide a detailed overview of each of these defense strategies.
Perturbation-based defense. This approach introduces carefully designed noise perturbations into the gradients and output features of the edge model during training to reduce the likelihood of adversaries extracting meaningful information and defend against MIA. He et al. (He et al., 2020) initially proposed using Gaussian noise and random feature dropout to limit the effective information accessible to attackers, and later, Wang et al. (Wang et al., 2022) extended this strategy by dynamically adjusting the depth of the edge model while adding differential privacy noise to transmitted features. Although these methods enhance defenses, they significantly impact model usability, often reducing accuracy by approximately 10%, and remain less effective against Gen-based MIA. This is because sufficiently expressive decoders in attack models can adapt to small noise or feature dropout, leveraging reconstruction capabilities similar to denoising auto-encoders (DAE) (Gondara, 2016), potentially increasing attack robustness. In centralized MIA defenses, Struppek et al. (Struppek et al., 2023) proposed perturbing training labels by setting the label smoothing factor to a negative value, which significantly reduced the effectiveness of MLE-based MIA and offers an alternative perspective for designing defense strategies.
IB-based defense. This approach enhances defenses against inference attacks by limiting the mutual information between input data $x$ and intermediate features $z$ while preserving the mutual information between features and outputs $y$, thereby maintaining predictive performance (Wang et al., 2021). The objective of this defense can be formally expressed as follows:
(2) $\min_{\theta}-{I}(z,y)+{I}(x,z)$
Initially, Wang et al. (Wang et al., 2021) introduced information bottleneck theory into MIA defenses, but their method was not optimized for collaborative inference and required substantial computational resources when applied to intermediate features. To address this, Peng et al. (Peng et al., 2022) proposed BiDO (Bilateral Dependency Optimization), which optimizes the correlation between inputs, features, and outputs using dependency algorithms. BiDO reduces the correlation between inputs and features while strengthening the correlation between features and outputs, balancing defense and performance. Additionally, Wang et al. (Wang et al., 2024) and Duan et al. (Duan et al., 2023) explored direct mutual information estimation between inputs and features using methods based on mutual information definitions and the CLUB (Contrastive Log Upper Bound) (Cheng et al., 2020) technique, incorporating the estimates as constraints in the optimization objective. IB-based defenses provide interpretable protection mechanisms and valuable design insights for future research. However, these methods face limitations: BiDO relies on kernel functions to constrain nonlinear correlations, and its performance is highly sensitive to hyper-parameter choices, such as kernel bandwidth. Poor parameter selection can result in suboptimal optimization. Methods based on mutual information definitions are constrained by sample scale to estimate $I(x,z)$.
NND & IB-based defenses. Building on the natural advantages of deep neural networks in resisting inference attacks and incorporating information bottleneck theory, this approach designs targeted loss functions to further enhance defense effectiveness. Building on this, Ding et al. (Ding et al., 2024) proposed increasing network depth while compressing newly added hidden layers to restrict the transmission of redundant information. Subsequently, Azizian et al. (Azizian and Bajić, 2024) proposed an improved defense method by introducing encoders at the edge and decoders at the cloud, forming an auto-encoder (AE) structure. The encoder at the edge increases network depth and reduces feature dimensionality, achieving the information bottleneck effect, while the decoder at the cloud restores information to maintain predictive performance. Additionally, an $l_{1}$-norm constraint constraint is applied during feature transmission to reduce attack risks, and adversarial training is used to enhance defense resilience against inference attacks.
However, this class of methods has certain limitations. First, increasing the depth of neural networks and adding additional modules inevitably raises computational overhead and hardware requirements on edge devices, limiting the method’s applicability. Second, the introduction of adversarial training not only obscures certain aspects of the original defense strategy’s effectiveness but also introduces additional challenges. It is less effective against future, more advanced attacks, increases the difficulty of fine-tuning edge models, and may degrade inference performance (Zhang et al., 2019). Additionally, adversaries can adapt their attack models using similar adversarial strategies, reducing the overall effectiveness of the defense (Jin et al., 2024). Thus, while these methods enhance defense performance, their implementation requires careful consideration of computational resources, model adaptability, and suitability for specific application scenarios.
3. Threat Model
3.1. Knowledge background of adversary
Attack scenarios. Based on the adversary’s level of knowledge, attack scenarios can be categorized into the following three types:
  1. (1)
    White-box Scenario: The adversary has complete white-box knowledge of the edge model, including its architecture, parameters, and gradient information.
  2. (2)
    Black-box Scenario: The adversary has only black-box access to the model, restricting their capability to observing the intermediate features. Additionally, their understanding of the edge model’s architecture is restricted to prior knowledge or incomplete assumptions.
  3. (3)
    Gray-box Scenario: The adversary has partial knowledge of the edge model’s architecture but remains unaware of defensive modifications, such as increased depth or additional hidden layers.
The gray-box scenario enables the adversary to leverage their partial understanding of the edge model’s architecture to enhance the capability of their inversion model. For instance, if the defender partitions the first three convolutional layers of a ResNet-18 model (He et al., 2016) and further deepens the network, while the adversary is only aware of the original ResNet-18 architecture, the adversary can only construct an approximate attack model based on partial knowledge of the initial layers. However, in the gray-box scenario, the adversary can address this limitation by utilizing their understanding of the edge model’s architecture and reasoning about its mapping capabilities to refine their inversion model.
Knowledge shared by adversaries in different scenarios. First, in attack scenarios, the adversary is assumed to understand the edge model’s core task. To enhance their capability, the attack model in Gen-based MIA is trained on data distributions closely matching the target distribution, as detailed in Section 6. Second, the adversary has access to lossless transmission features and their dimensional distribution, which are essential for attacks, as the features represent high-dimensional input data and the distribution informs the design of the attack model.
This study focuses on image classification tasks, though the proposed defense methods extend to other tasks where the edge model primarily performs feature extraction. Tasks like segmentation, detection, and classification rely on convolutional layers for shallow feature extraction, with task-specific variations having minimal impact on reconstruction objectives. Prior work (Azizian and Bajić, 2024) has also demonstrated input reconstruction effectiveness in object detection using Gen-based MIA.
Moreover, image data, due to its high dimensionality, is especially vulnerable to MIA attacks, as shallow convolutional layers retain redundant information (Zhang et al., 2020; He et al., 2016). This makes defenses in image-based tasks more demonstrable. In contrast, textual data, with its strong contextual dependencies and syntactic complexity, has shown limited success in existing researches for achieving high-quality MIA reconstructions (Dibbo, 2023; Zhang et al., 2022). This work focuses on image data to explore privacy-preserving techniques, leveraging its interpretability and the intuitive nature of defense effectiveness to validate the proposed methods.
3.2. Attack strategy of the adversary
In the white-box attack scenario, the adversary can perform attacks using the rMLE-based MIA method proposed in (He et al., 2019, 2020). To ensure the experiments are both comparative and representative, hyper-parameter adjustments were made to achieve performance closer to optimal.
In the black-box attack scenario, the adversary can employ methods from (Yang et al., 2019; He et al., 2019; Yin et al., 2023), relying solely on output features for reverse mapping without access to the edge model’s internal details. When protected by NND & IB-based defenses, the adversary lacks knowledge of architectural changes and must design the attack model based on prior assumptions.
In the gray-box attack scenario, the adversary can employ black-box strategies while leveraging known architectural information to optimize the attack model. This allows the reverse mapping capability to adapt to changes in the edge model. Detailed designs and training procedures for the attack models are provided in Section 6.
4. Theoretical Analysis of MIA in CI
This section analyzes key factors affecting MIA effectiveness in collaborative inference, introduces a formulaic representation of its principles, and establishes a difficulty criterion from an information-theoretic perspective. We validate this criterion through experiments and use it to guide the design of SiftFunnel.
4.1. Principles analysis of MIA
MIAs are categorized into MLE-based and Gen-based methods, both aiming to reconstruct input data using the target model’s output. MLE-based MIA calculates the loss between transmitted features and the edge model’s output, then iteratively applies gradient descent to optimize the input, achieving reconstruction of the original data. The attack principle is formulated as follows:
(3) $\begin{gathered}x^{(k+1)}=x^{(k)}-\eta\nabla_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right),\\ x^{*}=\arg\min_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right)\end{gathered}$
Where $x^{(k)}$ represents the input value at the $k$-th iteration, typically initialized as Gaussian noise, zeros, or ones. $x^{*}$ denotes the final reconstructed result, $z$ represents the transmitted features of the target input. $f_{\text{edge }}\left(\cdot\right)$ refers to the target edge model, $\nabla_{x}\mathcal{L}$ is the gradient calculated from the loss function and edge model parameters to optimize $x$, and $\eta$ is the step size for each update.
According to (He et al., 2019), the loss function combines the $l_{2}$ distance between $f_{\text{edge }}\left(x^{(k)}\right)$ and $z$ with a TV regularization term for $x$, forming the basis of rMLE-based MIA. The method relies on gradient descent, where the loss values depend on output features and gradients are computed using the target model’s parameters.
To achieve effective attacks, sufficient output feature information $\delta(z)$ and a well-chosen loss function are essential to ensure optimization space. Additionally, white-box knowledge of the edge model is required, as complex model mappings and limited feature information increase the risk of gradient descent converging to local minima (Struppek et al., 2022; Zhang et al., 2020). For instance, if two similar images produce nearly indistinguishable output features, reconstruction becomes challenging.
[…]
(4) $\begin{gathered}G=\arg\min_{G}\mathbb{E}_{x^{\prime}\sim p\left(x^{\prime}\right)}\mathcal{L}\left(G\left(f_{\text{edge }}\left(x^{\prime}\right)\right),x^{\prime}\right),\\ x^{*}=G(z)\end{gathered}$
Where $x^{\prime}$ represents samples from an auxiliary data distribution $p\left(x^{\prime}\right)$. The loss function, typically MSE, measures the distance between $x^{\prime}$ and the reconstructed data $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ or $z=f_{\text{edge }}\left(x^{\prime}\right)$. The trained generator $G$ must achieve the desired mapping $G\approx{f_{\text{edge}}}^{-1}$ to reconstruct $x^{*}$.
The above formula highlights the conditions required for effective attacks:
  1. (1)
    A well-matched training dataset aligned with the target task, ideally with a distribution similar to the target dataset (Yang et al., 2019).
  2. (2)
    Sufficient and highly separable input information for the generator.
The rationale is that if the information in $f_{\text{edge }}\left(x^{\prime}\right)$ is insufficient or cannot independently represent $x^{\prime}$’s features, the loss in (4) will struggle to converge as $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ remains ambiguous, limiting its training. These conditions must be met simultaneously. For instance, even with random feature drops, if the remaining features effectively distinguish inputs, the attack can still achieve satisfactory reconstruction (He et al., 2020). Such frameworks resemble DAE (Gondara, 2016), where the decoder can adapt to noise or feature drops.
Gen-based MIA demonstrates superior reconstruction performance compared to MLE-based MIA in defense scenarios. First, it avoids relying on model parameters for gradient computation, eliminating challenges from non-convex optimization. Second, it depends only on edge model output, offering greater flexibility. Furthermore, existing defenses struggle to effectively remove redundant feature information, allowing Gen-based MIA to overcome defenses if the generator is capable of learning the edge model’s reverse mapping.
4.2. Difficulty criterion for MIA in CI
Above, we summarized the factors influencing MIA effectiveness through a formulaic representation of its principles. However, this alone does not provide a concrete basis for evaluating the difficulty of implementing MIA. To address this, we will formalize and analyze these factors from an information-theoretic perspective to establish a clear criterion.
First, the mutual information between the input $x$ and the feature $z$, $I(x;z)$, can be expressed as:
(5) $I(x;z)=H(x)-H(x\mid z)=H(z)-H(z\mid x)=H(z)$
Where $H(x\mid z)$ represents the conditional entropy of input $x$ given $z$, and $H(z\mid x)$ represents the reverse. Since there is no uncertainty in the mapping from $x$ to $z$ through $f_{\text{edge}}$, $H(z\mid x)$ is zero.
From Equation (5), we observe that complex mappings and information loss during neural network transmission reduce $I(x;z)$ and increase $H(x|z)$, making non-convex optimization more difficult. Consequently, this increases the challenge for MLE-based MIA to reconstruct $x$ from $z$.
An inappropriate loss function or insufficient output information can further increase $H(x|z)$ and decrease $H(z)$. It is important to note that a low $\delta(z)$ does not necessarily imply a low $H(z)$. For instance, even if 80% of $z$’s elements are randomly dropped, $H(z)$ could remain high, as small amounts of information can still exhibit high entropy (He et al., 2020). Therefore, the performance of MLE-based MIA is directly influenced by $H(x|z)$, $H(z)$, and $\delta(z)$.
To further illustrate the influence of mutual information and entropy on the difficulty of MIA, we can derive the following relationship for any $x\to z\to x^{*}$ using Fano’s inequality (Arrow, 1969; Xu et al., 2024):
(6) $H\left(P_{e}\right)+P_{e}\log|\mathbb{X}|\geq H(x\mid x^{*})\geq H(x\mid z)$
where $P_{e}=Prob\{x\neq x^{*}\}$, $x^{*}$ is the reconstruction by MIA, and $\mathbb{X}$ denotes the value space of $x$.
Since $\mathbb{X}\geq 2$ and $H\left(P_{e}\right)\leq 1$. Therefore inequality (6) can be written as
(7) $P_{e}\geq\frac{H(x\mid z)-1}{\log|\mathbb{X}|}=\frac{H(x)-I(x,z)-1}{\log|\mathbb{X}|}=\frac{H(x)-H(z)-1}{\log|\mathbb{X}|}$
Since $H(x)$ and $|\mathbb{X}|$ are influenced by training data, they are typically assumed to be objective and invariant. Therefore, through inequality (7), we observe that increasing the lower bound of $P_{e}$, and thus the difficulty of MIA, is influenced by $H(x|z)$, $H(z)$, and $I(x,z)$.
From Equation (4), Gen-based MIA requires sufficient input information $\delta(z)$ and high $H(z)$ to ensure maximum separability and enable loss convergence, allowing the generator to simulate reverse mapping effectively. Notably, $\delta(z)$ and $H(z)$ are not directly proportional, as high average information does not always guarantee separability. A larger $\delta(z)$ not only facilitates generator training but also directly impacts the reconstruction of $x^{*}$. The relationship in Inequality (7) is also applicable here.
The validity of related research can also be explained by both. For example, (Yang et al., 2019) reduces $\delta(z)$ by pruning output information, disrupting training and reconstruction, while (Zhang et al., 2023) increases entropy via nonlinear functions to enhance generator performance. Thus, $H(z)$ and $\delta(z)$ are critical factors influencing Gen-based MIA.
The analysis indicates that the difficulty of implementing both types of attacks is inversely proportional to mutual information, entropy, and the amount of information. The difficulty of MIA can be formulated as follows, where $k_{1}$ and $k_{2}$ are proportionality constants.
(8) $D_{\text{mia}}\propto\left(\frac{k_{1}}{\{I(x;z);H(x\mid z);H(z)\}}+\frac{k_{2}}{\delta(z)}\right)$
Based on this criterion, existing defenses (Wang et al., 2021; Peng et al., 2022; Duan et al., 2023) using the information bottleneck theory can be understood as approximating a reduction in $I(x;z)$ through various algorithms to increase the difficulty of MIA. Perturbation-based defenses (He et al., 2020) aim to directly increase $H(x|z)$. The defensive effect of deeper neural networks (Ding et al., 2024; Azizian and Bajić, 2024) can be explained by their ability to reduce $I(x;z)$ and $\delta(z)$.
This is primarily because, without skip connections (Hao et al., 2025), mutual information decreases progressively during the forward propagation in neural networks. In contrast, residual networks with skip connections add the output of one layer to its input, which may prevent $\delta(z)$ from decreasing and, in some cases, even cause it to increase.
4.3. Experimental Demonstration of the Criterion and Guidance for Defense Design
To validate $D_{mia}$ and the theoretical analysis, we designed experiments using a CNN that progressively reduces spatial size while increasing channel size, based on the framework in (Yang et al., 2019; He et al., 2016), and ResNet-18 with skip connections, quantifying each block. We trained MINE (Belghazi et al., 2018) on the same dataset used to train the full model $f$ to estimate the lower bound of mutual information. The core idea of MINE is to optimize a neural network estimator using the Donsker-Varadhan representation (DV bound) (Donsker and Varadhan, 1983) as the loss function, providing an approximate estimate of the mutual information between random variables.
Mutual information is defined as follows:
(9) $I(X;Z)=\mathbb{E}_{p(x,z)}\left[\log\frac{p(x,z)}{p(x)p(z)}\right]$
Where $p(x,z)$ represents the joint distribution, existing defense methods approximate it, but limited batch sizes lead to inaccuracies and variations in mutual information estimates. For high-dimensional data (e.g., images), accurate estimation requires many samples.
To avoid direct distribution estimation, the Donsker-Varadhan representation reformulates mutual information as an optimization problem:
(10) $I(X;Z)\geq\sup_{T}\mathbb{E}_{p(x,z)}[T(x,z)]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T(x,z)}\right]$
Where $T$ is a learnable function, this formulation enables mutual information to be estimated by optimizing $T$ to obtain a lower bound.
MINE achieves this by parameterizing $T$ with a neural network $T_{\theta}$ and maximizing the DV bound through gradient descent to approximate mutual information:
(11) $\hat{I}(X;Z)=\mathbb{E}_{p(x,z)}\left[T_{\theta}(x,z)\right]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T_{\theta}(x,z)}\right]$
As shown in Figure 2, we trained target models on the CIFAR-10 (Krizhevsky et al., 2009) and used Equation (11) to train MINE for the corresponding blocks. Figures 2(a, b) show that increasing network depth reduces both mutual information and effective information, leading to a decline in attack performance. Figures 2(c, d) illustrate that while skip connections within two layers do not reduce effective information, they slightly lower mutual information, moderately affecting attack performance.
Figure 2. The impact of neural network depth on $I(x,z)$, $\delta(z)$, as well as MLE & Gen-based MIA.
[…]
2
Previous research (Hao et al., 2025) identified vulnerabilities caused by skip connections in centralized MIA. Our study further highlights their impact on MIA effectiveness in collaborative inference. Transmission between layers alters feature channels and dimensions, requiring a shortcut layer, which leads to a more pronounced decline in mutual information compared to direct stacking within layers. This, in turn, degrades MIA performance and aligns with our criterion.
Notably, while mutual information at ResNet’s fourth layer is higher than that at CNN’s second layer, MLE-based MIA does not perform better. This is because these methods rely heavily on gradient descent, making them sensitive to suboptimal optimizers, hyper-parameter settings, and deeper network structures, which hinder gradient calculations for optimizing $x$. The lack of flexible and efficient methods for MLE-based MIA remains a gap in current research, often leading to their exclusion in defense evaluations—a topic warranting further exploration (Ding et al., 2024).
In our experiments, we still evaluate MLE-based MIA against various defenses and provide evidence demonstrating the effectiveness of our method in defending against such attacks.
In conclusion, through theoretical analysis and experimental validation, we established $D_{mia}$ and determined that effective defense strategies must impact mutual information, entropy, and $\delta(z)$. The following considerations should guide defense design:
    Since $I(x,z)$ is challenging to compute and current algorithms lack flexibility and efficiency, defenses should focus on indirectly influencing $H(x\mid z)$ and $H(z)$.
    The design of edge models should strategically affect $H(z)$ and $\delta(z)$.
    For practical applications, the light-weighting of edge models should balance usability with their impact on the factors outlined in the criterion.
5. SiftFunnel: Information Sifting Funnel
Building on the theoretical analysis, criterion, and experiments in Section 4, this section introduces SiftFunnel and its adherence to $D_{mia}$ in edge model and loss function design.
We first outline the core principles and objectives of SiftFunnel. Then, we detail the edge model’s structure, key modules, and role in the forward propagation of information. Finally, we explain the loss function’s impact on $H(x\mid z)$, $H(z)$ and $\delta(z)$, highlighting its ability to extract useful information while minimizing redundancy.
5.1. Overview of SiftFunnel
Guided by $D_{{mia}}$, SiftFunnel aims to balance usability and privacy in edge models by increasing the uncertainty of inferring $x$ from $z$ while reducing $z$’s information content and separability. To this end, as shown in Figure 3, we introduce two key improvements: First, the edge model is enhanced by modifying the final hidden layer into a funnel-shaped structure and incorporating attention mechanism modules between layers. Second, the loss function is optimized to ensure that model training balances usability with privacy protection.
Specifically, in the edge model design, the final layer adopts a funnel-shaped structure, reducing feature channels while preserving spatial dimensions. This reduces computation, lightens the edge device load, and limits redundant information. Attention mechanisms and multi-receptive field convolutions enhance task-relevant information, while a feature channel supplementation module in the cloud model ensures task accuracy. With the improved edge model, designing an effective loss function is essential. To reduce redundancy and mutual information between $z$ and $x$, we use Distance Correlation (Li et al., 2012) for nonlinear correlations and the Pearson Coefficient (Cohen et al., 2009) with $l_{1}$-norm for linear correlations. Additionally, label smoothing and KL divergence (Donsker and Varadhan, 1983) ensure model accuracy, enhancing usability while reducing the separability of $z$.
5.2. The Design of Shallow Edge Model
Information Funnel. Deepening the neural network enhances feature extraction and reduces $I(x,z)$, as confirmed in the Section 4. Previous work (Ho et al., 2024) also demonstrates through Fisher information (Rissanen, 1996) that deeper networks improve feature relevance to tasks. Inspired by this, we redesign the final layer of the edge model with an Information Funnel module. This module reduces output channels to one-fourth in the first layer and further in the second layer, with channel selection based on the spatial size of feature maps. If the spatial size is reduced significantly, channels can exceed 10 to maintain accuracy. If unchanged, channels can be set to 2 to limit redundancy while preserving accuracy. Adjusting channels based on spatial size is necessary because spatial information loss is irreversible, whereas channel information loss can be compensated in the cloud model. While this compensation increases $I(z,y)$, it does not reverse the reduction in $I(x,z)$, ensuring attack difficulty remains high, as demonstrated in Section 6.3. So, we add the channel compensation module in the cloud model to guarantee usability. Notably, assumption about extracting features from non-independent transmission modules is extremely strong and idealized, requiring the adversary to have full white-box knowledge of the cloud model as if they were the developer.
Figure 3. Overview of SiftFunnel. Specifically, the design is divided into two aspects: the architecture and the loss function for the edge and cloud model.
[…]
3
The Information Funnel’s channel structure controls output but does not limit redundancy or enhance task-relevant information transmission, affecting usability. Inspired by Inception networks (Szegedy et al., 2017), we redesign it as a multi-receptive field convolutional framework with three low-channel layers. A $1*1$ kernel fuses channel information, followed by $3*3$ and $5*5$ kernels to expand the receptive field, extract richer task-relevant features, improve decision-making, and reduce mutual information with $x$, while lowering parameters of the edge model and controlling $\delta(z)$.
Attention Block. To enhance the output of task-relevant information from the Information Funnel, attention mechanism modules are added both before and after it. The first module combines an SE block (Hu et al., 2018) and Triplet Attention (Misra et al., 2021). The SE block evaluates the contribution of information within each channel, suppressing or nullifying channels with low task relevance. Triplet Attention reduces spatial information while maintaining smooth and continuous outputs, preserving fine-grained features suitable for further processing. The second attention mechanism integrates SE and CBAM modules (Woo et al., 2018). Compared to Triplet Attention, CBAM is a more lightweight module that sequentially calculates channel and spatial weights to enhance or suppress input features, improving network performance. Since CBAM’s channel and spatial attention are independently designed, certain regions are suppressed redundantly, making the output features more sparse. This is ideal for final output processing, effectively increasing $H(x|z)$ while suppressing $H(z)$ and $\delta(z)$.
5.3. The Design of Loss Function
To further enhance model usability and data privacy, a appropriately designed loss function is essential to guide the training process.
To reasonably increase $H(x\mid z)$ and suppress $H(z)$ within the edge model structure—thereby increasing the uncertainty of inferring $x$ from $z$—loss constraints must be applied to $x$ and $z$. The key challenge is that, without intervention, neural networks may retain irrelevant input information in their output during forward propagation. This information, which does not significantly impact decision-making, reduces the difficulty of attacks by maintaining $I(x,z)$.
Some defense approaches utilize Constrained Covariance (COCO) (Gretton et al., 2005b) to estimate correlation, focusing primarily on the linear relationship between $x$ and $z$. However, since neural networks predominantly perform nonlinear operations, studies have shown that this constraint is not only ineffective but may also lead to reduced accuracy (Peng et al., 2022). Alternatively, some approaches employ the Hilbert-Schmidt Independency Criterion (HSIC) (Gretton et al., 2005a) to estimate correlation. However, kernel-based calculations in HSIC rely heavily on the appropriate selection of $\sigma$. Improper $\sigma$ value can introduce bias in nonlinear correlation estimation, limiting the generalizability of this method and reducing its effectiveness across various scenarios.
Nonlinear correlation constraint. We observe that shallow neural network features retain significant similarity to the input and are derived through nonlinear operations. To address this, we use distance correlation to measure the nonlinear relationship between $x$ and $z$, incorporating it as a loss constraint. This helps remove redundant input information from the output, effectively increasing $H(x\mid z)$. The calculation of distance correlation is as follows:
(12) $\mathcal{L}_{\text{dCor}}=\mathrm{d}\operatorname{Cor}(x,z)=\frac{\mathrm{d}\operatorname{Cov}(x,z)}{\sqrt{\mathrm{d}\operatorname{Var}(x)\cdot\mathrm{d}\operatorname{Var}(z)}}$
Where $\mathrm{d}\operatorname{Cov}(x,z)$ represents the distance covariance between $x$ and $z$, and the denominator is the product of the distance variances for the two datasets.
To compute the numerator, a distance matrix $A$ is first calculated, where each element is the Euclidean distance between samples, defined as $A=\left\|x_{i}-x_{j}\right\|_{2}$. To remove the influence of sample shifts, the distance matrix is centered as $A_{ij}^{c}=A_{ij}-\overline{A_{i}}-\overline{A_{j}}+\overline{A_{i}}$, where $\overline{A_{i}}$ denotes the mean of the $i$-th column, and similar operations apply to other terms. With the centered distance matrix, the distance covariance is computed as $\mathrm{d}\operatorname{Cov}(x,z)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c}B_{ij}^{c}$, where $B$ represents the matrix of features $z$, and $n$ denotes the batch size.
The distance variances is defined as $\mathrm{d}\operatorname{Var}(x)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c^{2}}$, which normalizes the distance covariance. The value range of distance correlation is $[0,1]$, where $1$ indicates perfect correlation. Since $x$ and $z$ exhibit high distance similarity and are derived through nonlinear operations, distance correlation is well-suited as a constraint term. Moreover, as a distribution-free measure, it does not rely on specific data distributions, making it effective for non-Gaussian distributions, nonlinear relationships, and complex data types such as images and text. For non-stationary and multimodal data, distance correlation provides stable dependency measurements, allowing the constraint to flexibly adapt to different models and datasets.
Linear correlation constraint. In addition to constraining the loss between $x$ and $z$, it is necessary to design loss constraints specifically for the $z$. The primary issue lies in the $z$ containing repetitive or similar information, which is redundant for decision-making and provide additional input for Gen-based MIAs, strengthening local constraints in the input space and improving reconstruction accuracy.
To address this, we propose using the Pearson correlation coefficient as a loss constraint to reduce the linear correlation within features $z$, effectively eliminating redundant information. The calculation is as follows:
(13) $\mathcal{L}_{\text{Pearson }}=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\operatorname{pCor}_{c_{1},c_{2}}\right)^{2}=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\frac{\operatorname{Cov}_{c_{1},c_{2}}}{\sigma_{c_{1}}\cdot\sigma_{c_{2}}}\right)^{2}$
Where $\operatorname{Cov}_{c_{1},c_{2}}$ represents the covariance matrix between channels $c_{1}$ and $c_{2}$, calculated as $\operatorname{Cov}_{c_{1},c_{2}}=\frac{1}{\mathrm{H}\cdot\mathrm{W}}\sum_{\mathrm{i}=1}^{\mathrm{H}\cdot\mathrm{W}}\left(z_{c_{1},i}-\mu_{c_{1}}\right)\cdot\left(z_{c_{2},i}-\mu_{c_{2}}\right)$, where $z_{c_{1},i}$ is the $i$-th element of channel $c_{1}$ and $\mu_{c_{1}}$ is its mean. $H$ and $W$ denote the height and width of the feature space. $\sigma_{c_{1}}$ and $\sigma_{c_{2}}$ are the standard deviations, calculated as $\sigma_{c_{1}}=\sqrt{\frac{1}{\mathrm{H}\cdot\mathrm{W}}\sum_{\mathrm{i}=1}^{\mathrm{H}\cdot\mathrm{W}}\left(z_{c_{1},i}-\mu_{c_{1}}\right)^{2}}$
To compute the scalar $\mathcal{L}_{\text{Pearson }}$, the squared correlations of all channel pairs are averaged, keeping the value range within $[0,1]$. This avoids negative values, which could lead to suboptimal updates during gradient descent. Minimizing the $\mathcal{L}_{\text{Pearson }}$ in the presence of negative values might push the correlation toward complete negative correlation, introducing unnecessary feature redundancy or unsuitable learning patterns, ultimately preventing convergence. Therefore, the goal is to minimize the coefficient toward zero, as lower values indicate reduced linear correlation. Using the squared correlations and their average ensures a smoother and more stable gradient descent process.
Additionally, to prevent extreme values in the feature output weights during the optimization of the above constraints, we include the $l_{1}$-norm as a loss term. This term is added specifically to suppress weights, but excessive emphasis on it could interfere with normal model training. Therefore, the coefficient $\tau$ for this loss term is set below 1e-2 to balance its influence.
Label smoothing. Finally, to ensure model accuracy, we draw inspiration from (Struppek et al., 2023) and apply label smoothing to refine the target loss, as shown in the formula (14). Additionally, KL divergence is used instead of cross-entropy (CE) as the primary loss function for optimization.
(14) $LS\left(\hat{y}_{c}\right)=1-\alpha+\frac{\alpha}{K},\quad LS\left(\hat{y}_{i}\right)=\frac{\alpha}{K}$
Where $LS\left(\hat{y}_{c}\right)$ refers to the label smoothing element under the target class, $LS\left(\hat{y}_{i}\right)$ is the i-th element of the non-target class. $\alpha$ is smooth factor and $K$ is the total number of categories.
This smoothing mechanism ensures that the loss value computed by cross-entropy emphasizes enhancing inter-class feature distinctions, while KL divergence as a loss function supports smoother and more generalized feature extraction between classes (Struppek et al., 2023). Within-class feature extraction focuses on capturing universal characteristics, ensuring the model maintains a certain level of generalization and robustness for the target task. Moreover, this optimization reduces maximum feature separability, increasing the difficulty of attacks, as demonstrated in Section 6.
In summary, the proposed loss function first applies nonlinear correlation constraints between input $x$ and features $z$ to increase $H(x\mid z)$ and suppress $H(z)$. Second, it imposes linear correlation and extreme value constraints on $z$ to further enhance $H(x\mid z)$ and suppress $\delta(z)$. Finally, label smoothing is applied to the target loss to maintain accuracy while slightly suppressing $H(z)$. The final formulation of this optimization is as follows, with specific parameter settings referenced in Section 6.
(15) $\min\mathcal{L}_{\theta}(x,z,\hat{y})=$
$\lambda_{1}\cdot KL(f(x),LS(\hat{y}))+\lambda_{2}\cdot\mathcal{L}_{\text{dCor}}+\lambda_{3}\cdot\mathcal{L}_{\text{Pearson}}+\tau\cdot\|z\|_{1}$
6. Experiments
In this section, we systematically evaluate the resistance to MIA and compare our method with existing approaches. The results are analyzed from three perspectives: the impact of defense methods on $I(x\mid z)$ and $\delta(z)$, the influence of datasets and model architectures, and ablation studies. The specific experimental setup and evaluation metrics are shown below.
6.1. Experimental Setup
1) Datasets. We used three types of image recognition datasets, with processing details outlined as follows. The detailed data allocation is shown in Table 1.
  • CIFAR-10 (Krizhevsky et al., 2009). The CIFAR-10 dataset consists of 60,000 RGB images categorized into ten classes (e.g., airplane, ship, etc.) with an original resolution of $32\times 32$. To make the results clearer and more intuitive, we used an upscaled resolution of $64\times 64$.
  • FaceScrub (Ng and Winkler, 2014). FaceScrub is a URL dataset with 100,000 images of 530 actors, which contains 265 male actors and 265 female actors. However, since not every URL was available during the writing period, we downloaded a total of 43,149 images for 530 individuals and resized the images to $64\times 64$.
  • CelebA (Liu et al., 2015). CelebA is a dataset with 202,599 images of 10,177 celebrities from the Internet. We used the same crop as (Yang et al., 2019) to remove the background of images in this dataset other than faces to reduce the impact on the experiment. In order to eliminate individual overlap, we removed a total of 6,878 images of 296 individuals from CelebA and similarly resized the images to $64\times 64$.
  • ChestX-ray (Sait et al., 2020). This dataset is a curated collection of COVID-19 chest X-ray images compiled from 15 publicly available datasets. It contains 1,281 COVID-19 X-rays, 3,270 normal X-rays, 1,656 viral pneumonia X-rays, and 3,001 bacterial pneumonia X-rays. For our analysis, we converted the images to gray scale with a resolution of $128\times 128$.
Table 1. Data allocation of the target model and attack model.
Classifier Attack Model
Task Data Auxiliary Data
CIFAR-10 (10 classes) Train: 66.6%, Test: 16.7% 16.7% (10 classes of CIFAR-10)
FaceScrub (530 classes) Train: 80%, Test: 20% CelebA (non-individual overlapping)
ChestX-ray (4 classes) Train: 66.6%, Test: 16.7% 16.7% (4 classes of ChestX-ray)
2) Attack Method. The selection of attack methods has been discussed in Section 3.2. Specifically, for MLE-based MIA, this paper has adjusted to the hyperparameter values that yield the best results. For Gen-based MIA, the attack model in this paper is designed as the inverse architecture of the edge model. Moreover, under the gray-box scenario assumption, the attack model can enhance the expressive power of the reverse mapping as the edge model architecture changes. The attack model is also trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ of 0.5, and a ReduceLROnPlateaud scheduler with a factor of 0.5 and patience of 20. The training conditions for the attack models in this paper have been done with strong assumptions and only the attack models where the test Mean Square Error (MSE) reaches the minimum value are saved to highlight the effectiveness of the defense efforts.
3) Target Edge Model and Implementation Details. We employ the same comprehensive model architecture as detailed in (Yang et al., 2019), comprising four Convolutional Neural Network (CNN) blocks succeeded by two fully connected layers. This model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ parameter of 0.5, and a ReduceLROnPlateaud scheduler with a reduction factor of 0.5 and a patience setting of 25. Each CNN block is composed of a convolutional layer, a batch normalization layer, a max-pooling layer, and a ReLU activation function. The target edge model encompasses the initial two CNN blocks. For the training of the edge model protected by SiftFunnel, we have specified the parameters $\alpha$ to 3.5, $\epsilon$ to 0.35, $\beta$ to 0.8, and $\gamma$ to 0.6 in Equation (15). In addition to this, to argue the impact of the model architecture for the defense, we use different CNNs as follows: (1) VGG16 adapted from (Simonyan, 2014); (2) ResNet-18 adapted from (He et al., 2016), the learning rates are 3e-4. The division positions of resnet and vgg are not fixed. For detailed accuracy and how to divide them, see Section 6.3 for analysis. All experiments were performed on two RTX 4090 GPUs and an Intel Core i9-14900KF ×$\times 32$ CPU.
4) Comparison of Defense Methods. We will compare three existing categories of defense methods. For Perturbation-based defenses, we select Adding noise & Dropping features (He et al., 2020), DPSGD (Abadi et al., 2016), and negative label smoothing (NLS) (Struppek et al., 2023). For IB-based defenses, we choose MID (Wang et al., 2021), BiDO (with HSIC) (Peng et al., 2022), VIB (Wang et al., 2024), and InfoSCISSORS (Duan et al., 2023). For NND & IB-based defenses, we opt for AE-based protection (Azizian and Bajić, 2024). Our method falls into the NND & IB-based defense category, which also requires analyzing the defensive effect in a gray-box scenario. Furthermore, all comparison methods have been selected and set with appropriate hyper-parameters for comparison, with detailed settings found in Section 6.3 analysis. Lastly, all methods are trained without adversarial training, focusing solely on the effects of the original defense methods.
6.2. Evaluation Metrics
We have selected 7 evaluation metrics to assess usability, privacy, and deployability. For usability assessment, we employ Test Accuracy (Test ACC), which evaluates the final performance of the model protected by defense techniques on the test dataset. Regarding privacy assessment, we initially utilize MSE, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) to evaluate the effectiveness of MIA on reconstructing the input data of the protected model (Yin et al., 2023). Here, a higher MSE and lower PSNR and SSIM (ranging from 0 to 1) indicate better defense effectiveness. Compared to the former two, SSIM focuses more on the perceptual quality of the image, encompassing luminance, contrast, and structural information. In addition to these, we also employ MINE (Belghazi et al., 2018), as mentioned in Section  4.3, to assess Mutual Information (MI) (ranging from 0 to 1) and the Effective Information Mean $\delta(z)$, which is the mean of the non-zero information output by the edge model on the test dataset. The specific training methods and model architecture for MINE are referenced from xxx. For deployability, we consider the Edge model’s parameter count $|\theta_{\text{edge}}|$ and Inference Latency. The model parameter count directly determines memory or GPU memory occupation, while inference latency is tested on both CPU and GPU.
6.3. Experimental Results
Table 2. Quantitative evaluation of CNN on CIFAR-10 with different defenses against two MIAs. ↑ indicates higher is better, ↓ indicates lower is better. Bolded values are the best. Blue highlights show unprotected baseline performance, red indicates increased edge device load, yellow indicates reduced load, and gray shows gray-box defense performance.
MLE-based Attack (He et al., 2016) Gen-based Attack (He et al., 2020)
Method Class Method Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
Unprotected 88.28% 0.0045 28.3421 0.9883 0.0017 32.5498 0.9987 0.0566 32,549 299,520
Adding noise (He et al., 2020) ($\sigma$ = 0.8)
74.85% (-13.43%) 0.0291 20.1511 0.9135 0.0172 22.4249 0.9299 0.0502 65,535
Dropping (He et al., 2020) (r = 0.8)
21.18% (-67.10%) 0.0127 23.7340 0.9609 0.0135 23.4835 0.9481 0.0501 6,509
DPSGD (Abadi et al., 2016) (20, 1e-5)
61.83% (-26.45%) 0.0218 21.3868 0.9479 0.0016 32.4800 0.994 0.0623 39,829
Perturbation-based
NLS (Struppek et al., 2023) ($\alpha$ = -0.05)
87.60% (-0.68%)
0.0002 42.2684 0.9996 0.0014 33.3732 0.9949 0.0573 47,183 299,520
MID (Wang et al., 2021) (1e-2) 81.53% (-6.75%) 1.74e-5 50.4675 0.9999 0.0013 34.5184 0.9967 0.0584 45,198
BiDO (Peng et al., 2022) (2; 20) 84.12% (-4.16%) 0.6699 6.5111 0.0138 0.035 19.2965 0.8751 0.0429 3,345
VIB (Wang et al., 2024) 88.21% (-0.07%) 0.5552 7.3264 0.0358 0.0042 28.5054 0.9842 0.0551 18,069
IB-based InfoSCISSORS (Duan et al., 2023) 83.46% (-4.83%) 0.5715 7.2007 0.1345 0.0062 26.8799 0.9948 0.0549 43,364 299,520
0.0542 17.4852 0.8151
AE-based (Azizian and Bajić, 2024) 90.56% (+2.28%) 0.6345 6.7472 0.0051 0.01926 21.9259 0.9295 0.0358 2,617 2,365,056
0.1166 14.1112 0.5419
NND & IB-based SiftFunnel 85.49% (-2.79%) 0.6792 6.4510 0.0051 0.0639 16.7314 0.7647 0.0167 1,006 14,911
Figure 4. A visual evaluation of the attack effectiveness of two MIAs on CNN edge models protected by various defense techniques, with NND & IB-based results specifically reflecting the gray-box scenario.
[…]
4
Table 3. Quantitative evaluation of the resistance to MIA across different architectures on CIFAR-10. The edge model divisions for CNN and ResNet-18 follow the configurations in Tables 2 and 4, while for VGG16, the edge model is set at the first ReLU layer after a max-pooling operation.
MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
Architecture Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
0.1166 14.1112 0.5419
CNN 85.49% (-2.79%) 0.6792 6.4510 0.0051 0.0639 16.7314 0.7647 0.0167 1,006 14,911
0.1957 11.8562 0.0793
ResNet-18 88.43% (-2.00%) 0.6538 6.6166 0.0059 0.2790 10.3230 0.3001 0.4772 4,028 78,864
0.0827 15.5928 0.6559
VGG16 88.89% (-2.18%) 0.7801 5.8450 0.0005 0.0833 15.5916 0.7181 0.0128 3,587 52,086
Table 4. Quantitative evaluation of ResNet-18 on CIFAR-10 with different defenses against two MIAs. The target edge model is the first residual block of ResNet-18. ↑ indicates higher is better, ↓ indicates lower is better. Bolded values are the best. The meanings of color-coded regions are consistent with those in Table 2.
MLE-based Attack (He et al., 2016) Gen-based Attack (He et al., 2020)
Method Class Method Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
Unprotected 90.43% 0.0001 71.0996 0.9991 0.0010 34.6599 0.9991 0.9719 182,134 149,824
Adding noise (He et al., 2020) ($\sigma$ = 2)
81.39% (-9.04%)
0.0035 29.3354 0.9872 0.0054 27.4167 0.9794 0.9619 262,144
Dropping (He et al., 2020) (r = 0.8)
14.09% (-76.34%)
0.0069 26.3641 0.9754 0.0077 25.9334 0.9711 0.9514 35,574
DPSGD (Abadi et al., 2016) (20, 1e-5)
61.85% (-28.58%)
0.1901 11.9821 0.7172 0.0060 26.9751 0.9802 0.9515 178,956
Perturbation-based
NLS (Struppek et al., 2023) ($\alpha$ = -0.05)
88.57% (-1.86%)
0.0034 29.5001 0.9894 0.0011 34.4231 0.9961 0.9718 176,456 149,824
MID (Wang et al., 2021) (1e-2)
79.48% (-10.95%)
1.42e-7 73.2595 0.9999 0.0010 34.5893 0.9965 0.9729 173,587
BiDO (Peng et al., 2022) (5; 10)
90.08% (-0.35%)
4.38e-7 68.3603 0.9999 0.0020 31.7621 0.9927 0.9647 219,832
VIB (Wang et al., 2024)
89.18% (-1.25%)
0.5869 7.0853 0.0216 0.0020 31.7621 0.9928 0.9402 54,800
IB-based InfoSCISSORS (Duan et al., 2023)
88.03% (-2.40%)
3.10e-5 49.8548 0.9999 0.0058 27.1608 0.9799 0.8927 216,919 149,824
0.0143 23.2332 0.9488
AE-based (Azizian and Bajić, 2024) 89.19% (-1.24%) 0.6543 6.6189 0.1243 0.0094 25.0633 0.9660 0.9397 74,356 279,136
0.1957 11.8562 0.0793
NND & IB-based SiftFunnel 88.43% (-2.00%) 0.6538 6.6166 0.0059 0.2790 10.3230 0.3001 0.4772 4,028 78,864
Table 5. Quantitative evaluation of different defense methods in protecting CNN against MIAs on the FaceScrub dataset. Changes in data types alter the inference task, requiring adjustments to the parameters of the defense methods.
MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
Method Class Method Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
Unprotected 88.31% 0.0005 37.6115 0.9972 0.0006 32.4108 0.9908 0.0514 45,271 299,520
Perturbation-based
NLS (Struppek et al., 2023) ($\alpha$ = -0.2)
87.97% (-0.34%) 0.0004 38.8904 0.9980 0.0018 32.3274 0.9906 0.0496 44,783 299,520
MID (Wang et al., 2021) (1e-2) 79.45% (-10.98%) 0.0017 32.3435 0.9897 0.0018 32.1285 0.9901 0.0507 39,277
BiDO (Peng et al., 2022) (2; 20) 88.02% (-0.29%) 0.0004 39.0838 0.9978 0.0018 32.2048 0.9904 0.0493 44,915
VIB (Wang et al., 2024) 88.05% (-0.26%) 0.7354 6.1060 0.0032 0.0033 29.5697 0.9823 0.0493 15,649
IB-based InfoSCISSORS (Duan et al., 2023) 81.61% (-6.70%) 0.7449 6.0504 0.0026 0.0060 26.9803 0.9675 0.0408 57,050 299,520
0.0609 16.9260 0.6656
AE-based (Azizian and Bajić, 2024) 83.45 (-4.86%) 0.6474 6.6594 0.0090 0.0220 21.3444 0.8795 0.0238 585 2,365,056
0.0559 17.3018 0.6854
NND & IB-based SiftFunnel ($\lambda_{1}=5.0,\lambda_{2}=0.1,\lambda_{3}=0.1$) 85.39% (-2.92%) 0.7264 6.1593 0.0060 0.0296 20.0576 0.8323 0.0104 637 14,911
1) The effect of various defenses on the MI and $\delta(z)$. the effectiveness of defenses is reflected by MI and $\delta(z)$, aligning with the criterion $D_{\text{mia}}$ and the theoretical analysis presented in Section 4. Specifically, unprotected edge models show significant vulnerability to input reconstruction under both types of MIAs, with MI at 0.0566, MSE below 0.01, and SSIM exceeding 0.98. When protected by Perturbation-based defenses, all methods except NLS exhibit over a 10% drop in Test ACC. Moreover, these methods achieve only an 11% reduction in MI at most, demonstrating clear weaknesses against both MIAs. NLS shows a slight increase in mutual information for shallow neural networks, leading to weaker resistance to both types of attacks. Under IB-based defenses, usability sees a slight decline, but all methods except MID provide strong resistance against MLE-based MIA, with MSE values exceeding 0.5. Among these techniques, only BiDO significantly impacts MI and $\delta(z)$ , providing effective defense, but its parameter selection and kernel width configuration (as shown in the table) are challenging, limiting its generalizability—a limitation further evidenced in subsequent experiments. This experiment adheres to the InfoSCISSORS setup, providing CLUB with a complete dataset and sufficient training iterations to ensure convergence. While CLUB effectively defends against MLE-based MIA, it struggles under the experimental conditions to resist Gen-based MIA and results in a 4.83% accuracy loss. Under AE-based defenses, target model usability improves, likely because the original model lacked sufficient capacity for optimal task classification, and the added model complexity enhanced performance. Additionally, MI indicates notable reductions in mutual information, showcasing reasonable defense effectiveness. However, in gray-box scenarios, AE-based defenses show reduced resistance, and the added complexity imposes a eightfold storage burden on edge models, as revealed by $|\theta_{\text{edge}}|$. SiftFunnel achieves the best defense performance against both MIAs while maintaining an accuracy loss below 3%. For MLE-based MIA, it achieves an MSE of 0.6792, and for Gen-based MIA, an MSE above 0.1. In gray-box scenarios, it maintains an MSE of 0.06 and an SSIM of 0.7647, ensuring robust resistance. Furthermore, MI and $\delta(z)$ demonstrate that SiftFunnel exerts optimal influence on the key elements of $D_{\text{mia}}$. Additionally, $|\theta_{\text{edge}}|$ reveals that SiftFunnel reduces the edge model’s load by nearly 20 times.
2) The effect of model architecture. Changing the model architecture does not affect the parameter settings of the proposed method, which remain consistent with Section 6.1. We adapt the model to a ResNet-18 with skip connections, using the first residual block as the edge model. This setup is challenging for defense techniques, as shallow networks are simple and do not alter spatial dimensions. Table 4 shows minimal MI impact in unprotected models, leading to weak resistance against MIAs, with MLE-based MIA achieving an MSE of $10^{-4}$ and Gen-based MIA $10^{-3}$. Under these conditions, existing perturbation-based defenses fail to balance usability and privacy effectively, resulting in poor defense performance. Among IB-based defenses, only VIB demonstrates resistance to MLE-based MIA by leveraging the estimation of $I(x,z)$ as a loss to optimize the edge model. However, its effectiveness remains limited, particularly against Gen-based MIA. For BiDO, despite following the kernel bandwidth estimation methods from (Peng et al., 2022) and performing extensive experimental tuning, it fails to provide a configuration that balances usability and privacy, ultimately leading to suboptimal defense outcomes. InfoSCISSORS demonstrates some impact on MI under this configuration, but the increase in $\delta(z)$ undermines its defense effectiveness against Gen-based MIA, showing only marginal improvements. AE-based defenses exhibit certain resistance to attacks on low-dimensional features in earlier stages, but under the ResNet configuration, their performance diminishes. While this method reduces $\delta(z)$ by 60%, it has minimal impact on MI, with only a 0.04% decrease. Consequently, AE-based defenses struggle to counter both black-box and gray-box Gen-based MIAs effectively.
Table 6. SiftFunnel Ablation Study on CIFAR-10. The target edge model is the first residual block of ResNet-18, with Gen-based MIA as the attack method. The Red area indicates the strong attack hypothesis from Section 5.2.
Method Test ACC↑ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta$(z)↓
Unprotected 90.43% 0.0010 34.6599 0.9991 0.9719 182,134
Without Funnel (Average training round takes 1 minute, batchsize=64) 90.33% (-0.10%) 0.0594 17.0709 0.7776 0.8945 24,322
Without Attention Blocks 88.45% (-1.98%) 0.0525 17.5709 0.8178 0.8763 3,904
Without $KL$ and $LS$ 88.59% (-1.84%) 0.0850 15.4813 0.6905 0.7935 3,994
Without $\mathcal{L}_{\text{dCor}}$ 88.77% (-1.66%) 0.0268 20.4929 0.9025 0.9505 4,013
Without $\mathcal{L}_{\text{Pearson}}$ 88.83% (-1.60%) 0.0492 17.8576 0.8294 0.8618 4,037
Without $l_{1}$ 88.48% (-1.95%) 0.0789 15.8057 0.7272 0.8158 4,030
Attack SiftFunnel in Cloud 0.0521 17.6159 0.8205 0.8905 176,380
SiftFunnel 88.43% (-2.00%) 0.2790 10.3230 0.3001 0.4772 4,028
In contrast, SiftFunnel reduces MI by 50% and lowers $\delta(z)$ to the millesimal scale, achieving the best resistance against MIAs while maintaining usability and deployability. Notably, it achieves a minimum SSIM of 0.0793, highlighting its superior performance in balancing privacy and usability.
3) The effect of data type. Changes in data types lead to shifts in task requirements. For instance, in this study, the facial recognition task involves 530 classes, while ChestX-ray classification involves 4 classes, necessitating parameter adjustments as shown in Table 5. As analyzed in the previous section on model architectures, to achieve better defense performance, we adjusted the edge model divisions for ResNet-18 and VGG. Specifically, the division points were set after the second residual block for ResNet and after the first ReLU layer following the second max-pooling layer for VGG. The results, summarized in the table 5, show that the defense performance remains unaffected by changes in data types. SiftFunnel continues to be the best defense method, offering an optimal balance between usability, privacy, and deployability.
4) Ablation Study. To validate the effectiveness of SiftFunnel’s model structure and loss function design, as well as its defense capability, we conducted ablation experiments using Gen-based MIA. As shown in Table 6, we separately removed the Information Funnel and Attention Blocks from the model structure. The results indicate that removing the Funnel had minimal impact on accuracy, with only a 0.1% decrease, while removing the attention mechanism led to a significant increase in MI compared to SiftFunnel. For the loss function, we evaluated the impact of individually removing KL divergence, linear and nonlinear constraints, and the $l_{1}$-norm. The results demonstrate that KL divergence and $l_{1}$-norm moderately improved defense, whereas nonlinear constraints played a critical role; removing them increased the MSE to 0.0268. Additionally, we verified the hypothesis from Section 5.2, showing that even when adversaries reconstructed input using features from the cloud’s compensation module, SiftFunnel outperformed existing defenses in protecting transmitted features.
We also evaluated the processing speed of the edge model on both CPU and GPU under this configuration. For single-sample input, the unprotected edge model had a CPU latency of 1.204ms, the AE-based protected model had 2.812ms, and SiftFunnel achieved 2.192ms. On GPU, the processing latency was 3.9ms, 4.3ms, and 6.9ms, respectively. While SiftFunnel showed slightly higher latency on GPU, it required significantly less memory, enabling more parallel processing of inputs under the same resource constraints.
7. Discussion
Systematic experiments validated both the criterion and the effectiveness of the method. However, our analysis revealed that achieving optimal resistance against attacks remains challenging for current defense methods, including SiftFunnel, particularly when applied to shallow edge neural networks. When edge models perform only basic feature processing, even SiftFunnel can still allow Gen-based MIA to reconstruct blurred feature structures, while other methods are more vulnerable to input reconstruction. Enhancing privacy protection for extremely shallow edge neural networks presents a promising direction for future research.
Additionally, using MI as an evaluation metric, we observed that some defense methods result in only a 0.5% reduction in MI between inputs and features, yet current MLE-based MIAs still fail to bypass these defenses. This is directly linked to the lack of advancement in existing attack methods. Based on our experimental findings, we suggest that further exploration of more powerful MIAs in white-box scenarios could be a potential research direction.
8. Conclusion
The increasing complexity of neural networks and inference tasks, alongside the demand for efficiency and real-time feedback, presents significant challenges for edge devices with limited resources. Collaborative inference offers a solution by delegating feature extraction to edge devices and offloading subsequent tasks to the cloud, but this process leaves transmitted features vulnerable to MIAs. To address these vulnerabilities, we established a first criterion $D_{\text{mia}}$ for evaluating MIA difficulty and proposed SiftFunnel, a privacy-preserving framework designed to limit redundant information and enhance usability. By integrating linear and non-linear correlation constraints, label smoothing, and a funnel-shaped edge model with attention mechanisms, SiftFunnel achieves robust privacy protection while maintaining usability and deployability. Experimental results demonstrate its effectiveness, achieving superior defense against MIAs with minimal accuracy loss and a balanced trade-off among usability, privacy, and practicality.

### v4

@@
-Information Sifting Funnel: Privacy-preserving Collaborative Inference Against Model Inversion Attacks
+How Breakable Is Privacy: Probing and Resisting Model Inversion Attacks in Collaborative Inference
 Abstract.
-The complexity of neural networks and inference tasks, combined with the demand for computational efficiency and real-time feedback, poses significant challenges for resource-constrained edge devices.
-Collaborative inference addresses this by delegating shallow feature extraction to edge devices and uploading feature to the cloud for subsequent inference tasks, reducing computational pressure.
-However, the transmitted features remain vulnerable to model inversion attacks (MIAs), which can reconstruct original input data.
-Existing defenses, including perturbation and information bottleneck techniques, provide explainable protection but are limited by the lack of standardized criteria for evaluating MIA difficulty, challenges in mutual information estimation, and trade-offs among usability, privacy, and deployability.
-To overcome these limitations, we establish the first criterion for assessing the difficulty of MIAs in collaborative inference, supported by theoretical analysis of existing attacks and defenses and validated through experiments using the Mutual Information Neural Estimator (MINE).
-Building on these findings, we propose SiftFunnel, a privacy-preserving framework for collaborative inference.
-Specifically, we first train the edge model with both linear and non-linear correlation constraints to limit redundant information in transmitted features, enhancing privacy protection.
-Moreover, label smoothing and a cloud-based upsampling module are incorporated to balance usability and privacy.
-To further support deployability, the edge model is designed with a funnel-shaped structure and integrates attention mechanisms, preserving both privacy and usability.
-Extensive experiments show that SiftFunnel outperforms state-of-the-art defenses against MIAs, achieving superior privacy protection with less than 3% accuracy loss and an optimal trade-off among usability, privacy, and practicality.
+Collaborative inference (CI) improves computational efficiency for edge devices by transmitting intermediate features to cloud models.
+However, this process inevitably exposes feature representations to model inversion attacks (MIAs), enabling unauthorized data reconstruction.
+Despite extensive research, there is no established criterion for assessing the difficulty of MIA implementation, leaving a fundamental question unanswered: What factors truly and verifiably determine the attack’s success in CI?
+Moreover, existing defenses lack the theoretical foundation described above, making it challenging to regulate feature information effectively while ensuring privacy and minimizing computational overhead.
+These shortcomings introduce three key challenges: theoretical gap, methodological limitation, and practical constraint.
+To overcome these challenges, we propose the first theoretical criterion to assess MIA difficulty in CI, identifying mutual information, entropy, and effective information volume as key influencing factors.
+The validity of this criterion is demonstrated by using the mutual information neural estimator.
+Building on this insight, we propose SiftFunnel, a privacy-preserving framework to resist MIA while maintaining usability.
+Specifically, we incorporate linear and non-linear correlation constraints alongside label smoothing to suppress redundant information transmission, effectively balancing privacy and usability.
+To enhance deployability, the edge model adopts a funnel-shaped structure with attention mechanisms, strengthening privacy while reducing computational and storage burdens.
+Experiments show that, compared to state-of-the-art defense, SiftFunnel increases reconstruction error by $\sim$30%, lowers mutual and effective information metrics by $\geq$50%, and reduces edge burdens by almost 20×20$\times$20 ×, while maintaining comparable usability.
 Collaborative inference, model inversion attacks, deep learning security and privacy, information theory
@@
 1. Introduction
-The rapid advancement of deep learning has led to increasingly complex and powerful neural network models, enabling remarkable performance in tasks such as image detection and recognition, speech processing, and natural language understanding.
-However, these advancements also pose significant challenges for edge computing devices (Huckelberry et al., 2024).
-In fields such as intelligent transportation, autonomous driving, facial recognition, IoT applications involving cameras and sensors, and remote healthcare monitoring, edge devices must process data instantly while handling growing computational demands.
-The limited resources of edge devices make it difficult for them to efficiently perform complex deep learning inference tasks independently (Shlezinger and Bajić, 2022).
+The rapid advancement of deep learning has led to increasingly complex neural networks that achieve remarkable performance in tasks like image recognition and object detection.
+However, deploying such models on edge devices remains challenging due to limited computational resources (Huckelberry et al., 2024).
+Applications like autonomous driving, facial recognition, IoT surveillance, and remote healthcare require real-time processing, yet the limited computational capacity of edge devices makes standalone deep learning inference impractical (Shlezinger and Bajić, 2022).
+Cloud computing provides substantial computational power but faces latency, bandwidth, and privacy concerns when transmitting raw data.
+Collaborative inference (CI) (Shlezinger et al., 2021) mitigates these issues by partitioning neural networks into an edge model $f_{\text{edge}}$ and a cloud model $f_{\text{cloud}}$, where $f_{\text{edge}}$ extracts feature representations for $f_{\text{cloud}}$ to inference.
+This method enhances efficiency while maintaining real-time performance and is widely applied in UAVs (Qu et al., 2024), IoT systems (Shlezinger and Bajić, 2022), and private cloud computing (PCC) (Engineering and Architecture, 2024).
+With the rise of 5G, IoT, and AI, CI is poised for broader adoption.
+However, recent studies have revealed significant privacy risks in CI, where input data on edge devices can be reconstructed by model inversion attacks (MIAs) (He et al., 2019; Yang et al., 2022; Yin et al., 2023), as shown in Figure 1.
+Unlike centralized deployments, where adversaries recover training data through model APIs (Yuan et al., 2023; Struppek et al., 2022), CI often relies on shallow neural networks at the edge, producing feature representations that retain significant mutual information (MI) with input data (Azizian and Bajić, 2024).
+This exposes raw inputs to adversaries, who can reconstruct sensitive data, such as facial images directly from transmitted features (Ding et al., 2024).
+MIAs in CI can be categorized into maximum likelihood estimati-on-based (MLE-based) MIAs (He et al., 2019, 2020) and generative model-based (Gen-based) MIAs (Yin et al., 2023; He et al., 2020).
+MLE-based attacks optimize input data to minimize the loss between the edge model’s output and transmitted features but require white-box access, limiting practicality.
+Gen-based MIAs, on the other hand, use generative models to approximate the inverse mapping of the edge model, reconstructing inputs from features.
+Their success depends on data distribution alignment and the model’s ability to learn the inverse function.
+These privacy risks have driven extensive research on attacks and defenses, but existing methods still face notable limitations.
 Figure 1. Collaborative inference and the threat of model inversion attacks.
@@
 Collaborative inference and the threat of model inversion attacks.
-Cloud computing enables inference by uploading data to cloud-based models, offering significant computational power.
-However, it faces two key limitations: bandwidth and network latency hinder its ability to meet real-time application requirements, and uploading raw data raises privacy concerns, especially in sensitive scenarios.
-Collaborative inference (CI) (Shlezinger et al., 2021) has emerged as a viable solution to these challenges.
-By partitioning a neural network into $f_{\text{edge}}$ and $f_{\text{cloud}}$, the edge device processes raw inputs into feature representations via $f_{\text{edge}}$, which are then uploaded to the cloud for further inference by $f_{\text{cloud}}$.
-This approach balances real-time performance and computational efficiency.
-In recent years, this technique has garnered attention in practical applications, such as unmanned aerial vehicles (UAVs) (Qu et al., 2023) and IoT systems (Shlezinger and Bajić, 2022), as well as in solutions like Apple’s private cloud computing (PCC) (Engineering and Architecture, 2024).
-With advancements in 5G, IoT, and AI, collaborative inference holds promise for broader adoption and future impact.
-However, recent studies have revealed significant privacy risks in collaborative inference, where input data on edge devices can be reconstructed through model inversion attacks (MIAs) (He et al., 2019; Yang et al., 2022; Yin et al., 2023), as shown in Figure 1.
-MIAs are fine-grained privacy attacks that leverage the output information of AI models and deep learning generation techniques to reconstruct input data (He et al., 2019; Yang et al., 2019), recover representative training data (Zhang et al., 2020; Liu et al., 2024), or expose sensitive information (Fredrikson et al., 2014; Mehnaz et al., 2022).
-These risks are particularly severe in collaborative inference scenarios.
-The root cause is that, compared to centralized model deployment, edge devices typically employ shallow neural networks to compute and upload feature representations.
-These features contain far more information than the final inference results, enabling adversaries to reconstruct the input data (Azizian and Bajić, 2024).
-For instance, in facial recognition applications, while adversaries in centralized deployments might only recover training data through the model’s API (Yuan et al., 2023; Struppek et al., 2022), those in collaborative inference can directly reconstruct the raw facial data collected by edge devices from the transmitted features.
-[…]
-Firstly, existing defense techniques in collaborative inference lack a thorough and unified analysis of the fundamental nature of MIA.
-Current mainstream defense methods can be categorized into three types based on their technical approaches: perturbation-based (He et al., 2020; Wang et al., 2022), information bottleneck (IB)-based (Wang et al., 2024; Peng et al., 2022), and neural network depth & information bottleneck (NND & IB)-based defense strategies (Azizian and Bajić, 2024; Ding et al., 2024).
-Perturbation-based strategies aim to increase the difficulty of attacks by introducing perturbations to the gradients or output information of the trained model.
-IB-based strategies focus on approximating the reduction of mutual information $I(x,z)$ between the input data $x$ and the feature $z$, thus minimizing redundant information about $x$ in $z$ (Wang et al., 2021).
-NND & IB-based strategies leverage the observation that increasing the depth of the edge network can induce an information bottleneck effect, thus enhancing defense.
-These methods provide theoretically interpretable defense mechanisms, with the introduction of information theory concepts offering notable contributions to defense design.
-However, existing work lacks a comprehensive understanding of MIA, relies on weak assumptions about adversarial capabilities, and fails to establish a clear criterion for evaluating the complexity of implementing MIA.
-As a result, defenses against MIAs still have room for improvement, including addressing potential future advancements.
-Secondly, the trade-off between usability, privacy, and deployability in edge models remains a challenging and underexplored area in current research.
-Encryption-based defense strategies receive limited attention in collaborative inference due to resource constraints and the need for multitask processing in real time (Mireshghallah et al., 2021).
-Additionally, encryption does not guarantee complete data security, as modern side-channel attacks leveraging deep learning can recover encrypted data (Hu et al., 2023).
-The three main defense strategies discussed above also face issues of imbalance.
-Perturbation-based methods often compromise usability, whereas IB-based defenses aim to reduce mutual information $I(x,z)$, providing theoretical protection against simpler attacks such as MLE-based MIA.
-However, accurately computing $I(x,z)$ is challenging, as it relies on approximations through sufficient data and specific algorithms, making it less effective against more advanced attacks such as Gen-based MIA.
-Adversarial training has been introduced as a component to address the aforementioned issues, but its strong specificity and limited interpretability make it difficult to prevent future advanced attacks while significantly increasing the difficulty of training edge models.
-The NND & IB-based approach aims to balance privacy and usability by increasing model depth, but this approach significantly raises the resource demands on edge devices, including memory, storage, and bandwidth, thereby reducing deployability.
-As a result, such methods are often less favored in practical applications.
+Firstly, existing research lacks a comprehensive analysis of the fundamental factors influencing MIA effectiveness in CI.
+Current attack methods focus on improving data reconstruction under specific assumptions, leveraging innovative techniques to enhance reconstruction fidelity and generalization.
+In recent years, defenses have increasingly recognized the role of mutual information in MIAs and have shifted toward designing optimal information bottleneck (IB) approaches (Wang et al., 2024; Peng et al., 2022) to counter MIAs.
+However, these methods remain a methodological extension based on existing findings rather than advancing a fundamental theoretical understanding of MIAs.
+Consequently, their improvements are inherently constrained by existing technical limitations.
+Identifying the key factors that control MIA effectiveness and developing a systematic criterion for assessing attack difficulty would significantly deepen our theoretical insights and offer broader theoretical guidance for designing more effective defense strategies.
+Secondly, the trade-off between usability, privacy, and deployability in edge models is an underexplored and challenging area in current research.
+Encryption-based defenses receive little attention in CI due to resource constraints, real-time demands (Mireshghallah et al., 2021), and vulnerability to side-channel attacks (Hu et al., 2024).
+Mainstream MIA defense strategies in CI, constrained by a limited understanding of MIA effectiveness factors, primarily focus on reducing the mutual information $I(x,z)$ between inputs $x$ and extracted features $z$ to counter MIAs (Wang et al., 2024; Peng et al., 2022; Azizian and Bajić, 2024).
+However, mutual information estimation is inherently challenging due to its reliance on accurate probability density estimation, large-scale training data, and the computational overhead associated with optimization techniques (Wang et al., 2024; Duan et al., 2023), making the usability-privacy trade-off inefficient and difficult to guarantee.
+Moreover, some approaches compromise usability (He et al., 2020), increase computational and storage burdens on edge devices (Wang et al., 2024), and complicate further fine-tuning and optimization.
+Consequently, such methods are often less favored in practical applications.
 Thus, there is a pressing need for a novel approach grounded in an in-depth analysis of attack theories to provide interpretable and effective privacy protection while maintaining the usability and deployability of edge models.
-This need gives rise to three key technical challenges.
-Challenge 1: How to establish an experimentally validated criterion for assessing MIA difficulty in collaborative inference?
-Prior studies have interpreted MIA from an information-theoretic perspective, but challenges in computing mutual information and conducting an in-depth analysis of attack principles have hindered the establishment of a verifiable criterion, which is essential for guiding defense design and evaluating methods.
-Challenge 2: How to effectively and interpretably remove redundant information from transmitted features while preserving edge model usability?
-Although information bottleneck theory provides an interpretable defensive effect, mutual information is challenging to calculate, and current approximation methods are suboptimal.
-Challenge 3: How to balance usability and privacy while ensuring practical deployability?
-NND & IB-based methods significantly increase computational overhead on edge devices, making them challenging for practical application.
-Moreover, when applied to edge models that perform only simple feature extraction, these methods also struggle to resist advanced Gen-based MIAs.
-To overcome the aforementioned challenges, we established a criterion $D_{mia}$ for evaluating the difficulty of MIAs in CI and proposed a lightweight privacy-preserving strategy called Information Sifting Funnel (SiftFunnel), incorporating three key features.
-(1) SiftFunnel provides interpretable defense principles based on $D_{mia}$ We identified factors influencing attack difficulty in collaborative inference, including $I(x,z)$ and the effective information mean $\delta(z)$, i.e., the average number of non-zero elements in $z$.
-By analyzing these factors, we clarified defense design directions and developed a framework to address potential future threats.
-Using the Mutual Information Neural Estimator (MINE) (Belghazi et al., 2018), we quantified the changes in $I(x,z)$ at different network depths, validating the criterion.
-(2) SiftFunnel effectively reduces redundant information in transmitted features while preserving edge model usability.
-Guided by the proposed criterion, we enhanced the model’s loss function to include distance-correlation and Pearson-correlation measures, alongside an $l_{1}$-norm constraint on $z$.
-These refinements reduce $H(z)$, $\delta(z)$ while increasing $H(x|z)$.
-To further filter redundant information, we integrated Squeeze-and-Excitation (SE) Modules (Hu et al., 2018), Triple Attention Mechanisms (Misra et al., 2021), and Convolutional Block Attention Modules (CBAM) (Woo et al., 2018) into the edge model to focus on task-relevant features.
-Usability was preserved through a combination of label-smoothing and KL-divergence loss.
-(3) SiftFunnel balances usability and privacy while light-weighting the edge model.
-A funnel-shaped edge model was designed to progressively reduce transmitted channels, improving efficiency and reducing redundant information while minimizing memory and storage demands.
-To maintain usability, an upsampling module was added on the cloud side.
-Based on prior evaluation metrics, this paper incorporates $I(x,z)$, quantified using MINE, and $\delta(z)$ as key metrics to validate the theoretical analysis and assess the effectiveness of defense methods.
-To evaluate deployability, we consider the edge model’s parameter count, memory usage, storage consumption, and inference latency.
-Through extensive experiments, we demonstrate the validity of the proposed difficulty criterion and the effectiveness of SiftFunnel in balancing usability, deployability, and data privacy compared to existing methods.
+This need gives rise to three challenges.
+Challenge 1 (Difficulty assessment): How to establish a verifiable criterion for assessing MIA difficulty in CI?
+The fundamental factors determining MIA effectiveness remain poorly understood, making it difficult to develop a reliable assessment criterion.
+While MI’s role in MIAs is recognized, its accurate estimation for validation remains challenging.
+Challenge 2 (Methodological design): How to effectively and interpretably remove redundant information from transmitted features while preserving edge model usability?
+Although IB theory provides a theoretically sound approach, its practical implementation is limited by the technical constraints of MI estimation.
+Current methods are computationally expensive and lack flexibility, making fine-tuning and further optimization difficult after initial deployment.
+Challenge 3 (Practical constraint): How to balance usability and privacy while ensuring practical deployability?
+Certain defense strategies (Azizian and Bajić, 2024) inevitably increase computational and storage burdens on edge devices.
+Furthermore, when applied to edge models designed for simple feature extraction, these methods struggle to counter advanced Gen-based MIAs effectively.
+To overcome the above challenges, we proposed a criterion $D_{mia}$ to evaluate MIA difficulty in CI and proposed a lightweight privacy-preserving strategy called Information Sifting Funnel (SiftFunnel), featuring 3 key aspects.
+(1) Criterion basis: SiftFunnel offers interpretable defense principles based on $D_{mia}$ We analyze MIAs using formulaic principles, information theory, and Fano’s inequality, and derive an evaluation formula involving 4 critical factors: mutual information $I(x,z)$, conditional entropy $H(x\mid z)$, feature entropy $H(z)$, and effective information mean $\delta(z)$ (the average number of non-zero elements in $z$).
+Their impact on MIA performance is validated via the Mutual Information Neural Estimator (MINE) (Belghazi et al., 2018).
+(2) Usability-privacy trade-off: SiftFunnel effectively reduces redundant information while preserving edge model usability.
+Aligning with $D_{mia}$ to strengthen privacy protection, we aim to reduce $H(z)$ and $\delta(z)$ while increasing $H(x\mid z)$.
+Specifically, we incorporate distance correlation, Pearson correlation constraints, and an $l_{1}$-norm constraint on $z$ to enforce these objectives.
+To maintain usability, we integrate label smoothing and KL-divergence loss into the training process.
+This optimization enables the edge model to minify redundant information while ensuring the extraction of task-relevant features.
+(3) Lightweight design: SiftFunnel balances usability and privacy while light-weighting the edge model.
+We adopt a funnel-shaped edge model to reduce computational and storage overhead.
+Attention blocks enhance feature selection, while an upsampling module on the cloud side compensates for lost details, preserving inference accuracy.
+This design, guided by $D_{mia}$, also improves privacy by limiting feature $\delta(z)$ exposure.
+Except for previous evaluation metrics, we incorporate MI and $\delta(z)$ as metrics to assess defense effectiveness.
+Deployability is measured by the edge model’s parameter count and inference latency.
+Systematic experiments confirm that SiftFunnel effectively balances usability, deployability, and privacy compared to existing methods, achieving superior MIA resistance while reducing mutual information and storage, with an average accuracy trade-off of 3%.
 Our contributions are summarized as follows.
-  • We establish the first difficulty criterion for implementing MIA in collaborative inference, providing guidance for defense design and preventing future advanced MIAs.
-  • We propose SiftFunnel, a privacy-preserving scheme for collaborative inference that constrains redundant transmitted features through linear and nonlinear correlation, ensuring model usability with label smoothing and loss function regularization.
-  • SiftFunnel adopts a funnel-shaped structure with integrated attention mechanisms to balance usability and privacy while maintaining the deployability of edge models.
-  • We conduct systematic experiments validating the proposed criterion and demonstrating SiftFunnel’s advantages in balancing usability, privacy, and deployability.
+  • We propose the first MIA difficulty criterion in CI, identifying $I(x,z)$, $H(x|z)$, $H(z)$ and $\delta(z)$ as factors impacting attack effectiveness, deepening insights into MIAs.
+  • We propose SiftFunnel, a privacy-preserving scheme for CI that constrains redundant transmitted features through linear and nonlinear correlation, ensuring model usability with label smoothing and loss function regularization.
+  • SiftFunnel adopts a funnel-shaped structure with attention blocks to balance usability and privacy while ensuring the lightweight design of edge models.
+  • We conduct systematic experiments validating the proposed criterion and demonstrating the effectiveness of SiftFunnel in balancing usability, privacy, and deployability.
 2. Background and Related Work
 2.1. Collaborative inference
-The rapid development of artificial intelligence has enabled deep learning models to achieve exceptional performance in image processing and other tasks.
-However, the computational demands of these models often exceed the capabilities of resource-constrained edge devices, which face limitations in processing power, storage, and energy efficiency.
-To address these challenges, collaborative inference (Shlezinger et al., 2021) has emerged as an effective solution.
+Collaborative inference (Shlezinger et al., 2021) has emerged as an effective solution.
 This approach divides the inference process into two stages: one part is performed locally on the edge device (which may consist of a single device or a chain of interconnected devices), while the remaining computation is uploaded to the cloud.
@@
 where $f$ is the full trained model, $x$ is the input, $f_{e}$ represents the model deployed on the edge, and $f_{c}$ corresponds to the portion deployed in the cloud.
-Collaborative inference typically begins with training $f$ on centralized datasets in the cloud, followed by partitioning and deploying the $f_{e}$ to the edge.
-Alternatively, methodologies such as Split Learning  (Zhu et al., 2023) enable edge devices to participate in model training locally.
+Collaborative inference typically begins with training $f$ on centralized datasets in the cloud, then partitioning and deploying the $f_{e}$ to the edge.
+Alternatively, approaches like Split Learning  (Zhu et al., 2023) allow edge devices to participate in model training locally.
 Regardless of the training approach, the inference process remains consistent.
-As depicted in Figure 1, the edge-deployed model first computes an intermediate representation $feature=f_{e}(x)$, which is then transmitted to the cloud for completion of the inference process $result=f_{c}(feature)$.
-The final output is either sent back to the edge device or retained in the cloud for further analysis.
-Compared to the traditional approach of uploading raw data to the cloud for inference, collaborative inference offers distinct advantages.
-First, it reduces communication overhead by allowing the transmitted features to be processed, potentially making them smaller in size than the original data.
-Second, it improves privacy protection by limiting the exposure of sensitive information by transmitting processed features instead of raw data, such as physiological signals or facial images.
-These advantages make collaborative inference a suitable solution for applications in healthcare, facial recognition, and industrial IoT, where efficiency and privacy are critical requirements.
-However, existing research highlights that collaborative inference models are susceptible to Model Inversion Attacks (MIA), where adversaries attempt to reconstruct the original input data from transmitted feature.
+As depicted in Figure 1, the edge model computes an intermediate representation, which is transmitted to the cloud for completion of the inference process.
+The final output is either returned to the edge device or retained in the cloud for further analysis.
+However, existing research (He et al., 2019) highlights that edge models are susceptible to MIAs, where adversaries reconstruct input data from transmitted features.
 Notably, in edge-based inference, computations can be distributed across multiple devices in a chained architecture (He et al., 2019), expressed as $f=f_{e1}\circ f_{e2}...\circ f_{c}$.
-Studies show that deeper neural networks exhibit greater resistance to data reconstruction, making devices positioned closer to the cloud less vulnerable.
-In contrast, the initial edge device, which typically employs a shallow neural network for feature computation and transmission, generates representations that retain more information from the original input, increasing its susceptibility to MIA.
-Given this vulnerability, this paper focuses on the MIA threat to the initial edge device, which uses a shallow neural network while the rest resides in the cloud.
-This common scenario in real-world applications poses a significant security risk due to the heightened likelihood of input data reconstruction, underscoring the need to address this challenge.
+Studies (He et al., 2020; Ho et al., 2024) show that deeper neural networks closer to the cloud transmit more complex, decision-focused features, making them more resistant to reconstruction.
+In contrast, the initial edge device generates features retaining more input information, making it highly susceptible to MIAs.
+This paper focuses on the MIA threat to the initial edge device, which employs a shallow neural network while the remaining computation occurs in the cloud.
+This common scenario in real-world applications presents a significant security risk due to the increased likelihood of input data reconstruction, highlighting the need for effective countermeasures.
 2.2. Model inversion attack
-Model Inversion Attack (MIA) is a fine-grained privacy attack that uses the output of a target model and adversarial knowledge, such as white-box access or training data distribution, to reconstruct training data (Zhang et al., 2020), sensitive attributes (Mehnaz et al., 2022), or input data (Yang et al., 2019).
-Initially proposed by Fredrikson et al.
-(Fredrikson et al., 2015), MIA was applied to linear regression models, successfully reconstructing genomic data from warfarin dosage predictions and patient attributes, thereby revealing the privacy risks.
-Subsequent studies (Fredrikson et al., 2015) extended this concept to deep learning, showing that training data from deep neural networks could also be reconstructed.
-Current MIA techniques in collaborative inference can be broadly categorized into Maximum Likelihood Estimation-based (MLE-based) MIAs and Generative Model-based (Gen-based) MIAs, which are further analyzed in the following sections.
+Current MIA techniques in CI can be broadly categorized into MLE-based MIAs and Gen-based MIAs, with related research progress described as follows.
 MLE-based MIAs.
-MLE-based MIA originated from the 2015 study by Fredrikson et al.
-(Fredrikson et al., 2015), which reconstructed training data from a multilayer perceptron (MLP).
-This method minimizes the difference between target and output class confidence scores as a loss function and employs gradient descent to optimize the input, enabling the reconstruction of the original training data.
-Subsequent research enhanced this approach by incorporating techniques like Total Variation (TV) (Rudin et al., 1992) loss as a regularization term to improve optimization, with these improved methods later referred to as rMLE-based MIA by He et al.
-(He et al., 2019) In collaborative inference scenarios, MLE-based methods are particularly effective for reconstructing data from shallow neural networks deployed on edge devices.
-Unlike the prediction vectors produced by full neural networks, the richer feature representations and lower complexity of shallow networks provide adversaries with more detailed and easily optimizable gradient information.
+MLE-based MIA was first introduced by Fredrikson et al.
+(Fredrikson et al., 2015), who reconstructed training data from a multilayer perceptron (MLP) by minimizing the difference between target and confidence scores using gradient descent.
+Subsequent research (He et al., 2019) incorporated Total Variation (TV) (Rudin et al., 1992) loss as a regularization term to improve optimization, with these advanced techniques referred to as rMLE-based MIA.
+In this paper, both are collectively referred to as MLE-based MIA.
+In CI scenarios, MLE-based methods are particularly effective for reconstructing data from shallow neural networks, as the richer feature representations and lower complexity provide adversaries with detailed gradient information that is easier to optimize compared to the prediction vectors of deeper networks.
 Gen-based MIAs.
-This method originated from the 2019 work of Yang et al.
+This approach was introduced by Yang et al.
 (Yang et al., 2019), who reconstructed input data in facial recognition systems by using generative deep learning models $g$ to approximate the inverse mapping function of the target model $f^{-1}$, such that $g\approx f^{-1}$.
-By inputting prediction vectors into $g$, they successfully generated facial input images.
+By feeding prediction vectors into $g$, they successfully generated facial input images.
 Building on this approach, He et al.
-(He et al., 2019) extended it to collaborative inference scenarios, demonstrating that input data from edge models is more vulnerable to reconstruction, thereby highlighting the risks associated with collaborative inference.
+(He et al., 2019) extended it to collaborative inference, demonstrating that input data from edge models is more vulnerable to reconstruction, thereby highlighting the risks.
 Unlike MLE-based methods, Gen-based MIA does not require white-box access to the edge model.
-Instead, it trains the attack model by calculating losses based on transmitted features, enabling it to approximate the inverse mapping of non-linear computations.
-Furthermore, as edge models in collaborative inference are typically shallow neural networks, their outputs often retain redundant information from the original input and features, maintaining a strong correlation between inputs and outputs and facilitating the reconstruction process.
+Instead, it trains an attack model by calculating losses based on transmitted features and auxiliary data, enabling it to approximate the inverse mapping of non-linear computations.
+Furthermore, as edge models produce rich and redundant outputs, this facilitates the training of attack models and the reconstruction process, as analyzed in detail in Section 4.
 2.3. Related defense work
-To protect the input data collected by edge devices in collaborative inference from potential leakage, He et al.
-(He et al., 2019, 2020) highlighted the necessity of developing defense strategies while identifying these threats.
-Building on a detailed analysis of attack methods, subsequent defense techniques have gradually evolved in recent years.
-These defenses can be broadly categorized into three types: perturbation-based defenses, information bottleneck-based defenses, and defenses combining neural network depth with information bottleneck principles.
-The following sections provide a detailed overview of each of these defense strategies.
+To safeguard input data collected by edge devices in CI from potential leakage, He et al.
+(He et al., 2019, 2020) emphasized the importance of developing effective defense strategies while identifying these vulnerabilities.
+Existing defenses can be classified into three main categories: perturbation-based, IB-based, and neural network depth (NND) & IB-based approaches.
+The following sections outline these strategies.
 Perturbation-based defense.
-This approach introduces carefully designed noise perturbations into the gradients and output features of the edge model during training to reduce the likelihood of adversaries extracting meaningful information and defend against MIA.
+This approach introduces carefully designed noise perturbations into the gradients for training or trained edge model output features to reduce the likelihood of adversaries extracting meaningful information and defending against MIA.
 He et al.
@@
 (Wang et al., 2022) extended this strategy by dynamically adjusting the depth of the edge model while adding differential privacy noise to transmitted features.
-Although these methods enhance defenses, they significantly impact model usability, often reducing accuracy by approximately 10%, and remain less effective against Gen-based MIA.
-This is because sufficiently expressive decoders in attack models can adapt to small noise or feature dropout, leveraging reconstruction capabilities similar to denoising auto-encoders (DAE) (Gondara, 2016), potentially increasing attack robustness.
+Although these methods enhance defenses, they significantly impact model usability, typically reducing accuracy by approximately 10%, and remain less effective against Gen-based MIA.
+The root cause is that the decoder in Gen-based MIAs is sufficiently expressive to adapt to noise or feature dropout, similar to the training of a denoising auto-encoder (DAE) (Gondara, 2016), potentially increasing attack robustness.
 In centralized MIA defenses, Struppek et al.
-(Struppek et al., 2023) proposed perturbing training labels by setting the label smoothing factor to a negative value, which significantly reduced the effectiveness of MLE-based MIA and offers an alternative perspective for designing defense strategies.
+(Struppek et al., 2023) proposed perturbing training labels with a negative label smoothing factor, significantly reducing MIA effectiveness and offering a novel perspective for designing defense strategies.
 IB-based defense.
-This approach enhances defenses against inference attacks by limiting the mutual information between input data $x$ and intermediate features $z$ while preserving the mutual information between features and outputs $y$, thereby maintaining predictive performance (Wang et al., 2021).
+This approach enhances defenses against inference attacks by limiting the mutual information between input data $x$ and intermediate features $z$ while preserving the mutual information between $z$ and outputs $y$, thus maintaining inference performance (Wang et al., 2021).
 The objective of this defense can be formally expressed as follows:
@@
 Initially, Wang et al.
-(Wang et al., 2021) introduced information bottleneck theory into MIA defenses, but their method was not optimized for collaborative inference and required substantial computational resources when applied to intermediate features.
+(Wang et al., 2021) introduced information bottleneck theory into MIA defenses, but their method was not optimized for CI and required substantial computational resources for intermediate feature processing.
 To address this, Peng et al.
-(Peng et al., 2022) proposed BiDO (Bilateral Dependency Optimization), which optimizes the correlation between inputs, features, and outputs using dependency algorithms.
-BiDO reduces the correlation between inputs and features while strengthening the correlation between features and outputs, balancing defense and performance.
+(Peng et al., 2022) proposed BiDO (Bilateral Dependency Optimization), which uses dependency algorithms to optimize correlations among inputs, features, and outputs.
+BiDO reduces input-feature correlation while enhancing feature-output correlation, achieving a balance between defense and performance.
 Additionally, Wang et al.
 (Wang et al., 2024) and Duan et al.
-(Duan et al., 2023) explored direct mutual information estimation between inputs and features using methods based on mutual information definitions and the CLUB (Contrastive Log Upper Bound) (Cheng et al., 2020) technique, incorporating the estimates as constraints in the optimization objective.
+(Duan et al., 2023) explored mutual information estimation between $x$ and $z$ using methods based on mutual information definition and the CLUB (Contrastive Log Upper Bound) (Cheng et al., 2020) technique, incorporating the estimates as constraints in the training loss function.
 IB-based defenses provide interpretable protection mechanisms and valuable design insights for future research.
-However, these methods face limitations: BiDO relies on kernel functions to constrain nonlinear correlations, and its performance is highly sensitive to hyper-parameter choices, such as kernel bandwidth.
-Poor parameter selection can result in suboptimal optimization.
-Methods based on mutual information definitions are constrained by sample scale to estimate $I(x,z)$.
-NND & IB-based defenses.
-Building on the natural advantages of deep neural networks in resisting inference attacks and incorporating information bottleneck theory, this approach designs targeted loss functions to further enhance defense effectiveness.
-Building on this, Ding et al.
-(Ding et al., 2024) proposed increasing network depth while compressing newly added hidden layers to restrict the transmission of redundant information.
+However, these methods face limitations: BiDO relies on kernel functions for non-linear correlation constraints, with performance highly sensitive to hyper-parameters like kernel bandwidth; poor parameter selection can lead to suboptimal results.
+Methods based on mutual information estimation are constrained by sample scale to compute.
+NND & IB-based defense.
+Leveraging deeper networks and IB theory, this approach enhances defense effectiveness with targeted loss functions.
+Ding et al.
+(Ding et al., 2024) proposed increasing network depth while compressing newly added hidden layers to restrict redundant information transmission.
 Subsequently, Azizian et al.
-(Azizian and Bajić, 2024) proposed an improved defense method by introducing encoders at the edge and decoders at the cloud, forming an auto-encoder (AE) structure.
-The encoder at the edge increases network depth and reduces feature dimensionality, achieving the information bottleneck effect, while the decoder at the cloud restores information to maintain predictive performance.
-Additionally, an $l_{1}$-norm constraint constraint is applied during feature transmission to reduce attack risks, and adversarial training is used to enhance defense resilience against inference attacks.
-However, this class of methods has certain limitations.
-First, increasing the depth of neural networks and adding additional modules inevitably raises computational overhead and hardware requirements on edge devices, limiting the method’s applicability.
-Second, the introduction of adversarial training not only obscures certain aspects of the original defense strategy’s effectiveness but also introduces additional challenges.
-It is less effective against future, more advanced attacks, increases the difficulty of fine-tuning edge models, and may degrade inference performance (Zhang et al., 2019).
-Additionally, adversaries can adapt their attack models using similar adversarial strategies, reducing the overall effectiveness of the defense (Jin et al., 2024).
-Thus, while these methods enhance defense performance, their implementation requires careful consideration of computational resources, model adaptability, and suitability for specific application scenarios.
+(Azizian and Bajić, 2024) further improved defenses by introducing an auto-encoder (AE) structure, with encoders at the edge and decoders in the cloud.
+The edge encoder increases network depth and reduces feature dimensionality to achieve the IB effect, while the cloud decoder restores information to maintain predictive performance while applying an $l_{1}$-norm constraint to reduce attack risks.
+Adversarial training (Ding et al., 2024) is also incorporated to strengthen robustness against MIAs.
+However, this method has certain limitations.
+Increasing network depth and adding modules raise computational and hardware demands on edge devices, limiting applicability.
+Adversarial training, while strengthening defenses, introduces additional challenges—it is less effective against advanced attacks, complicates fine-tuning, and degrades inference performance (Zhang et al., 2019).
+Additionally, adversaries can adapt their attack models using similar adversarial strategies, weakening defense effectiveness (Jin et al., 2024).
+While these methods improve defenses, their implementation requires careful consideration of computational resources, model adaptability, and application suitability.
+SiftFunnel enhances computational efficiency by deepening edge model structures and adopting a lightweight design, exhibiting IB-like characteristics after training.
+Thus, our method falls under the NND & IB-based defense strategy.
 3. Threat Model
@@
   2. (2)
-    Black-box Scenario: The adversary has only black-box access to the model, restricting their capability to observing the intermediate features.
-Additionally, their understanding of the edge model’s architecture is restricted to prior knowledge or incomplete assumptions.
+    Black-box Scenario: The adversary has only black-box access to the model, restricting their capability to observe the intermediate features.
+Their understanding of the edge model’s architecture is also restricted to prior knowledge or incomplete assumptions.
   3. (3)
-    Gray-box Scenario: The adversary has partial knowledge of the edge model’s architecture but remains unaware of defensive modifications, such as increased depth or additional hidden layers.
-The gray-box scenario enables the adversary to leverage their partial understanding of the edge model’s architecture to enhance the capability of their inversion model.
-For instance, if the defender partitions the first three convolutional layers of a ResNet-18 model (He et al., 2016) and further deepens the network, while the adversary is only aware of the original ResNet-18 architecture, the adversary can only construct an approximate attack model based on partial knowledge of the initial layers.
-However, in the gray-box scenario, the adversary can address this limitation by utilizing their understanding of the edge model’s architecture and reasoning about its mapping capabilities to refine their inversion model.
+    Gray-box Scenario: The adversary has partial knowledge of the edge model’s architecture but remains unaware of defensive modifications, such as increased hidden layers.
+The gray-box scenario enables the adversary to leverage their partial understanding of the edge model’s architecture to enhance the capability of their attack model.
+For instance, if the defender partitions the first three convolutional layers of a ResNet-18 model (He et al., 2016) and further deepens the network, while the adversary is only aware of the original ResNet-18 architecture, they can only construct an approximate attack model based on partial knowledge of the initial three layers.
+However, adversaries can address this limitation in this scenario by utilizing their understanding of the edge model’s architecture and reasoning about its mapping capabilities to refine their attack model.
 Knowledge shared by adversaries in different scenarios.
-First, in attack scenarios, the adversary is assumed to understand the edge model’s core task.
-To enhance their capability, the attack model in Gen-based MIA is trained on data distributions closely matching the target distribution, as detailed in Section 6.
-Second, the adversary has access to lossless transmission features and their dimensional distribution, which are essential for attacks, as the features represent high-dimensional input data and the distribution informs the design of the attack model.
-This study focuses on image classification tasks, though the proposed defense methods extend to other tasks where the edge model primarily performs feature extraction.
-Tasks like segmentation, detection, and classification rely on convolutional layers for shallow feature extraction, with task-specific variations having minimal impact on reconstruction objectives.
-Prior work (Azizian and Bajić, 2024) has also demonstrated input reconstruction effectiveness in object detection using Gen-based MIA.
-Moreover, image data, due to its high dimensionality, is especially vulnerable to MIA attacks, as shallow convolutional layers retain redundant information (Zhang et al., 2020; He et al., 2016).
-This makes defenses in image-based tasks more demonstrable.
-In contrast, textual data, with its strong contextual dependencies and syntactic complexity, has shown limited success in existing researches for achieving high-quality MIA reconstructions (Dibbo, 2023; Zhang et al., 2022).
-This work focuses on image data to explore privacy-preserving techniques, leveraging its interpretability and the intuitive nature of defense effectiveness to validate the proposed methods.
+The adversary is assumed to understand the edge model’s task and enhances their attack by training Gen-based MIA on a closely matching data distribution (Section 6.1).
+Additionally, the adversary has access to lossless transmission features and their dimensional distribution, which facilitate optimizing the MIA loss function and training the attack model.
+This study focuses on image classification but emphasizes that the defense methods apply to other tasks where the edge model primarily handles feature extraction.
+Since segmentation, detection, and classification also rely on edge-processed features, MIA and defenses are task-agnostic.
+Prior work (Azizian and Bajić, 2024) has demonstrated input reconstruction and defense effectiveness in object detection using Gen-based MIA.
+Moreover, image data is particularly vulnerable to MIAs due to its high dimensionality, as shallow convolutional layers retain more redundant information (Zhang et al., 2020; He et al., 2016).
+This increases the challenge of protecting image data, making it an ideal test case for privacy defenses.
+In contrast, textual data, with its strong contextual dependencies and syntactic complexity, has shown limited success in achieving high-quality MIA reconstructions (Dibbo, 2023; Zhang et al., 2022).
+This study focuses on image data due to its interpretability, making it suitable for evaluating privacy defenses.
 3.2. Attack strategy of the adversary
-In the white-box attack scenario, the adversary can perform attacks using the rMLE-based MIA method proposed in (He et al., 2019, 2020).
+In the white-box attack scenario, the adversary can perform attacks using the MLE-based MIA method proposed in (He et al., 2019, 2020).
 To ensure the experiments are both comparative and representative, hyper-parameter adjustments were made to achieve performance closer to optimal.
@@
 4. Theoretical Analysis of MIA in CI
-This section analyzes key factors affecting MIA effectiveness in collaborative inference, introduces a formulaic representation of its principles, and establishes a difficulty criterion from an information-theoretic perspective.
-We validate this criterion through experiments and use it to guide the design of SiftFunnel.
 4.1. Principles analysis of MIA
-MIAs are categorized into MLE-based and Gen-based methods, both aiming to reconstruct input data using the target model’s output.
-MLE-based MIA calculates the loss between transmitted features and the edge model’s output, then iteratively applies gradient descent to optimize the input, achieving reconstruction of the original data.
+MIAs are categorized into MLE-based and Gen-based methods, both aiming to reconstruct input data using the output of the edge model.
+MLE-based MIA calculates the loss between transmitted features and the output, then iteratively applies gradient descent to optimize the input, achieving reconstruction of the original data.
 The attack principle is formulated as follows:
@@
 According to (He et al., 2019), the loss function combines the $l_{2}$ distance between $f_{\text{edge }}\left(x^{(k)}\right)$ and $z$ with a TV regularization term for $x$, forming the basis of rMLE-based MIA.
-The method relies on gradient descent, where the loss values depend on output features and gradients are computed using the target model’s parameters.
+The method relies on gradient descent, where the loss values depend on output features, and gradients are computed using the target model’s parameters.
 To achieve effective attacks, sufficient output feature information $\delta(z)$ and a well-chosen loss function are essential to ensure optimization space.
@@
 For instance, if two similar images produce nearly indistinguishable output features, reconstruction becomes challenging.
-[…]
+When dealing with complex edge model mappings or models protected by defense techniques, directly computing gradients in MLE-based MIA can significantly hinder optimization performance.
+To address this, He et al.
+(He et al., 2019, 2020), building on the method from (Yang et al., 2019), developed a reverse mapping network for the edge model, which is categorized as Gen-based MIA.
+The key idea is to use a generator to approximate the reverse mapping of the edge model.
+Different methods adopt various strategies to train the generator.
+For instance, Yang et al.
+(Yang et al., 2019) trained a decoder using auxiliary task-related datasets, while Yin et al.
+(Yin et al., 2023) utilized intercepted feature data for training.
+The principle of this attack is summarized as follows.
 (4) $\begin{gathered}G=\arg\min_{G}\mathbb{E}_{x^{\prime}\sim p\left(x^{\prime}\right)}\mathcal{L}\left(G\left(f_{\text{edge }}\left(x^{\prime}\right)\right),x^{\prime}\right),\\ x^{*}=G(z)\end{gathered}$
 Where $x^{\prime}$ represents samples from an auxiliary data distribution $p\left(x^{\prime}\right)$.
-The loss function, typically MSE, measures the distance between $x^{\prime}$ and the reconstructed data $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ or $z=f_{\text{edge }}\left(x^{\prime}\right)$.
+The loss function, typically mean square error, measures the distance between $x^{\prime}$ and the reconstructed data $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ or $z=f_{\text{edge }}\left(x^{\prime}\right)$.
 The trained generator $G$ must achieve the desired mapping $G\approx{f_{\text{edge}}}^{-1}$ to reconstruct $x^{*}$.
@@
 The rationale is that if the information in $f_{\text{edge }}\left(x^{\prime}\right)$ is insufficient or cannot independently represent $x^{\prime}$’s features, the loss in (4) will struggle to converge as $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ remains ambiguous, limiting its training.
-These conditions must be met simultaneously.
-For instance, even with random feature drops, if the remaining features effectively distinguish inputs, the attack can still achieve satisfactory reconstruction (He et al., 2020).
+Among them, the maximum separability is more decisive.
+For instance, even with random feature drops, if the remaining features distinguish inputs well, the attack can still reconstruct effectively (He et al., 2020).
 Such frameworks resemble DAE (Gondara, 2016), where the decoder can adapt to noise or feature drops.
@@
 First, it avoids relying on model parameters for gradient computation, eliminating challenges from non-convex optimization.
-Second, it depends only on edge model output, offering greater flexibility.
+Second, it depends only on the output of the edge model, offering greater flexibility.
 Furthermore, existing defenses struggle to effectively remove redundant feature information, allowing Gen-based MIA to overcome defenses if the generator is capable of learning the edge model’s reverse mapping.
 4.2. Difficulty criterion for MIA in CI
-Above, we summarized the factors influencing MIA effectiveness through a formulaic representation of its principles.
-However, this alone does not provide a concrete basis for evaluating the difficulty of implementing MIA.
-To address this, we will formalize and analyze these factors from an information-theoretic perspective to establish a clear criterion.
+We summarized the factors influencing MIA effectiveness using a formulaic representation.
+To further evaluate the difficulty of MIA implementation, we will formalize and analyze these factors from an information-theoretic perspective to establish a clear criterion.
 First, the mutual information between the input $x$ and the feature $z$, $I(x;z)$, can be expressed as:
@@
 Since there is no uncertainty in the mapping from $x$ to $z$ through $f_{\text{edge}}$, $H(z\mid x)$ is zero.
-From Equation (5), we observe that complex mappings and information loss during neural network transmission reduce $I(x;z)$ and increase $H(x|z)$, making non-convex optimization more difficult.
-Consequently, this increases the challenge for MLE-based MIA to reconstruct $x$ from $z$.
-An inappropriate loss function or insufficient output information can further increase $H(x|z)$ and decrease $H(z)$.
-It is important to note that a low $\delta(z)$ does not necessarily imply a low $H(z)$.
-For instance, even if 80% of $z$’s elements are randomly dropped, $H(z)$ could remain high, as small amounts of information can still exhibit high entropy (He et al., 2020).
-Therefore, the performance of MLE-based MIA is directly influenced by $H(x|z)$, $H(z)$, and $\delta(z)$.
-To further illustrate the influence of mutual information and entropy on the difficulty of MIA, we can derive the following relationship for any $x\to z\to x^{*}$ using Fano’s inequality (Arrow, 1969; Xu et al., 2024):
+Factors influencing MLE-based MIA.
+From Equation (5), we observe that complex mappings and information loss during neural network forward propagation reduce $I(x;z)$ and increase $H(x|z)$, making non-convex optimization with backpropagation more difficult.
+Consequently, reconstructing $x$ from $z$ becomes more difficult for MLE-based MIA.
+These methods primarily compare $z$ with generated features, but a poorly trained or protected model can produce $z$ with high $\delta(z)$ yet low $H(z)$, indicating suboptimal separability.
+This weakens attack effectiveness despite the apparent information sufficiency.
+Thus, MLE-based MIA difficulty is more influenced by $I(x;z)$, $H(x|z)$, and feature entropy $H(z)$.
+We can demonstrate the influence of mutual information and entropy on MLE-based MIA using the following relationship for any $x\to z\to x^{*}$, derived from Fano’s inequality (Arrow, 1969; Xu et al., 2025):
 (6) $H\left(P_{e}\right)+P_{e}\log|\mathbb{X}|\geq H(x\mid x^{*})\geq H(x\mid z)$
 where $P_{e}=Prob\{x\neq x^{*}\}$, $x^{*}$ is the reconstruction by MIA, and $\mathbb{X}$ denotes the value space of $x$.
-Since $\mathbb{X}\geq 2$ and $H\left(P_{e}\right)\leq 1$. Therefore inequality (6) can be written as
+Since $\mathbb{X}\geq 2$ and $H\left(P_{e}\right)\leq 1$. Therefore, inequality (6) can be written as
 (7) $P_{e}\geq\frac{H(x\mid z)-1}{\log|\mathbb{X}|}=\frac{H(x)-I(x,z)-1}{\log|\mathbb{X}|}=\frac{H(x)-H(z)-1}{\log|\mathbb{X}|}$
 Since $H(x)$ and $|\mathbb{X}|$ are influenced by training data, they are typically assumed to be objective and invariant.
-Therefore, through inequality (7), we observe that increasing the lower bound of $P_{e}$, and thus the difficulty of MIA, is influenced by $H(x|z)$, $H(z)$, and $I(x,z)$.
+Therefore, we observe that increasing the lower bound of $P_{e}$, and thus the difficulty of MIA, is influenced by $H(x\mid z)$, $H(z)$, and $I(x,z)$.
+Factors influencing Gen-based MIA.
 From Equation (4), Gen-based MIA requires sufficient input information $\delta(z)$ and high $H(z)$ to ensure maximum separability and enable loss convergence, allowing the generator to simulate reverse mapping effectively.
-Notably, $\delta(z)$ and $H(z)$ are not directly proportional, as high average information does not always guarantee separability.
 A larger $\delta(z)$ not only facilitates generator training but also directly impacts the reconstruction of $x^{*}$.
-The relationship in Inequality (7) is also applicable here.
-The validity of related research can also be explained by both.
-For example, (Yang et al., 2019) reduces $\delta(z)$ by pruning output information, disrupting training and reconstruction, while (Zhang et al., 2023) increases entropy via nonlinear functions to enhance generator performance.
-Thus, $H(z)$ and $\delta(z)$ are critical factors influencing Gen-based MIA.
-The analysis indicates that the difficulty of implementing both types of attacks is inversely proportional to mutual information, entropy, and the amount of information.
-The difficulty of MIA can be formulated as follows, where $k_{1}$ and $k_{2}$ are proportionality constants.
+At the same time, the relationship in Inequality (7) also holds here (Zhang et al., 2023).
+Therefore, in addition to the factors influencing MLE-based MIA, Gen-based MIA is further effectively affected by $\delta(z)$.
+Criterion formula.
+The analysis indicates that the difficulty of implementing both types of MIA is inversely proportional to the mutual information, entropy, and output information.
+Therefore, the difficulty criterion for MIA can be formulated as follows, where $k_{1}$ and $k_{2}$ are proportionality constants:
 (8) $D_{\text{mia}}\propto\left(\frac{k_{1}}{\{I(x;z);H(x\mid z);H(z)\}}+\frac{k_{2}}{\delta(z)}\right)$
 Based on this criterion, existing defenses (Wang et al., 2021; Peng et al., 2022; Duan et al., 2023) using the information bottleneck theory can be understood as approximating a reduction in $I(x;z)$ through various algorithms to increase the difficulty of MIA.
-Perturbation-based defenses (He et al., 2020) aim to directly increase $H(x|z)$.
+Perturbation-based defenses (He et al., 2020) aim to directly increase $H(x\mid z)$.
 The defensive effect of deeper neural networks (Ding et al., 2024; Azizian and Bajić, 2024) can be explained by their ability to reduce $I(x;z)$ and $\delta(z)$.
-This is primarily because, without skip connections (Hao et al., 2025), mutual information decreases progressively during the forward propagation in neural networks.
-In contrast, residual networks with skip connections add the output of one layer to its input, which may prevent $\delta(z)$ from decreasing and, in some cases, even cause it to increase.
-4.3. Experimental Demonstration of the Criterion and Guidance for Defense Design
-To validate $D_{mia}$ and the theoretical analysis, we designed experiments using a CNN that progressively reduces spatial size while increasing channel size, based on the framework in (Yang et al., 2019; He et al., 2016), and ResNet-18 with skip connections, quantifying each block.
-We trained MINE (Belghazi et al., 2018) on the same dataset used to train the full model $f$ to estimate the lower bound of mutual information.
-The core idea of MINE is to optimize a neural network estimator using the Donsker-Varadhan representation (DV bound) (Donsker and Varadhan, 1983) as the loss function, providing an approximate estimate of the mutual information between random variables.
+This is mainly because, without skip connections (Hao et al., 2025), mutual information progressively decreases during forward propagation in neural networks.
+In contrast, residual networks with skip connections preserve or even enhance $\delta(z)$ by adding layer outputs to inputs, but the decline in mutual information still poses challenges for MIA (see section 4.3 for experimental cases).
+4.3. Demonstration and Guidance for Defense
+To validate $D_{mia}$ and the theoretical analysis, we designed experiments using a CNN that progressively reduces the spatial size while increasing the channel size, based on the framework in (Yang et al., 2019; He et al., 2016), and ResNet-18 with skip connections, quantifying each block.
+We trained MINE (Belghazi et al., 2018) on the same dataset used to train the full model $f$ to estimate the lower bound of $I(x,z)$.
+The core idea of MINE is to optimize a neural network estimator using the Donsker-Varadhan (DV) representation (Donsker and Varadhan, 1983) as the loss function, providing an approximate estimate of the mutual information between random variables.
 Mutual information is defined as follows:
 (9) $I(X;Z)=\mathbb{E}_{p(x,z)}\left[\log\frac{p(x,z)}{p(x)p(z)}\right]$
-Where $p(x,z)$ represents the joint distribution, existing defense methods approximate it, but limited batch sizes lead to inaccuracies and variations in mutual information estimates.
+Where $p(x,z)$ represents the joint distribution, existing defense methods approximate it, but limited batch sizes lead to inaccuracies and variations in estimates.
 For high-dimensional data (e.g., images), accurate estimation requires many samples.
-To avoid direct distribution estimation, the Donsker-Varadhan representation reformulates mutual information as an optimization problem:
+To avoid direct distribution estimation, the DV representation reformulates mutual information as an optimization problem:
 (10) $I(X;Z)\geq\sup_{T}\mathbb{E}_{p(x,z)}[T(x,z)]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T(x,z)}\right]$
-Where $T$ is a learnable function, this formulation enables mutual information to be estimated by optimizing $T$ to obtain a lower bound.
+Where $T$ is a learnable function, this formulation enables $I(X;Z)$ to be estimated by optimizing $T$ to obtain a lower bound.
 MINE achieves this by parameterizing $T$ with a neural network $T_{\theta}$ and maximizing the DV bound through gradient descent to approximate mutual information:
 (11) $\hat{I}(X;Z)=\mathbb{E}_{p(x,z)}\left[T_{\theta}(x,z)\right]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T_{\theta}(x,z)}\right]$
-As shown in Figure 2, we trained target models on the CIFAR-10 (Krizhevsky et al., 2009) and used Equation (11) to train MINE for the corresponding blocks.
-Figures 2(a, b) show that increasing network depth reduces both mutual information and effective information, leading to a decline in attack performance.
-Figures 2(c, d) illustrate that while skip connections within two layers do not reduce effective information, they slightly lower mutual information, moderately affecting attack performance.
+As shown in Figure 2, we trained the target models on the CIFAR-10 (Krizhevsky et al., 2009) and used Equation (11) to train MINE for the corresponding blocks.
+Figures 2(a, b) show that increasing network depth reduces both mutual information and effective information, leading to a decrease in attack performance.
+Figures 2(c, d) illustrate that while skip connections within two layers do not reduce $\delta(z)$, they slightly lower $I(x,z)$, moderately affecting attack performance.
+Previous research (Hao et al., 2025) identified vulnerabilities caused by skip connections in centralized MIA.
+Our study further highlights their impact on MIA effectiveness in CI.
 Figure 2. The impact of neural network depth on $I(x,z)$, $\delta(z)$, as well as MLE & Gen-based MIA.
@@
 2
-Previous research (Hao et al., 2025) identified vulnerabilities caused by skip connections in centralized MIA.
-Our study further highlights their impact on MIA effectiveness in collaborative inference.
-Transmission between layers alters feature channels and dimensions, requiring a shortcut layer, which leads to a more pronounced decline in mutual information compared to direct stacking within layers.
+However, Transmission between blocks in ResNet alters feature channels and dimensions, requiring a shortcut layer, leading to a more pronounced decline in mutual information compared to direct stacking within layers.
 This, in turn, degrades MIA performance and aligns with our criterion.
+Figure 3. Overview of SiftFunnel. Specifically, the design is divided into two aspects: the architecture and the loss function for the edge and cloud model.
+[…]
+3
 Notably, while mutual information at ResNet’s fourth layer is higher than that at CNN’s second layer, MLE-based MIA does not perform better.
-This is because these methods rely heavily on gradient descent, making them sensitive to suboptimal optimizers, hyper-parameter settings, and deeper network structures, which hinder gradient calculations for optimizing $x$.
+This stems from the fact that these methods rely heavily on gradient descent, making them sensitive to suboptimal optimizers, hyper-parameter settings, and deeper network structures, which hinder gradient calculations for optimizing $x$.
 The lack of flexible and efficient methods for MLE-based MIA remains a gap in current research, often leading to their exclusion in defense evaluations—a topic warranting further exploration (Ding et al., 2024).
-In our experiments, we still evaluate MLE-based MIA against various defenses and provide evidence demonstrating the effectiveness of our method in defending against such attacks.
-In conclusion, through theoretical analysis and experimental validation, we established $D_{mia}$ and determined that effective defense strategies must impact mutual information, entropy, and $\delta(z)$.
-The following considerations should guide defense design:
+In our experiment and Section 6, we still evaluate MLE-based MIA against various defenses and provide evidence demonstrating the effectiveness of our method in defending against such attacks.
+In conclusion, through theoretical analysis and experimental validation, we proposed $D_{mia}$ and determined that effective defense strategies must impact mutual information, entropy, and $\delta(z)$.
+The following considerations can guide defense design:
     Since $I(x,z)$ is challenging to compute and current algorithms lack flexibility and efficiency, defenses should focus on indirectly influencing $H(x\mid z)$ and $H(z)$.
@@
 5. SiftFunnel: Information Sifting Funnel
-Building on the theoretical analysis, criterion, and experiments in Section 4, this section introduces SiftFunnel and its adherence to $D_{mia}$ in edge model and loss function design.
-We first outline the core principles and objectives of SiftFunnel.
-Then, we detail the edge model’s structure, key modules, and role in the forward propagation of information.
-Finally, we explain the loss function’s impact on $H(x\mid z)$, $H(z)$ and $\delta(z)$, highlighting its ability to extract useful information while minimizing redundancy.
 5.1. Overview of SiftFunnel
-Guided by $D_{{mia}}$, SiftFunnel aims to balance usability and privacy in edge models by increasing the uncertainty of inferring $x$ from $z$ while reducing $z$’s information content and separability.
-To this end, as shown in Figure 3, we introduce two key improvements: First, the edge model is enhanced by modifying the final hidden layer into a funnel-shaped structure and incorporating attention mechanism modules between layers.
+Guided by $D_{{mia}}$, SiftFunnel balances usability and privacy by increasing uncertainty in inferring $x$ from $z$ while limiting redundant information and separability.
+As shown in Figure 3, SiftFunnel introduces two key improvements: First, the edge model modifies the final hidden layer into a funnel-shaped structure and incorporates attention modules.
 Second, the loss function is optimized to ensure that model training balances usability with privacy protection.
@@
 This reduces computation, lightens the edge device load, and limits redundant information.
-Attention mechanisms and multi-receptive field convolutions enhance task-relevant information, while a feature channel supplementation module in the cloud model ensures task accuracy.
-With the improved edge model, designing an effective loss function is essential.
-To reduce redundancy and mutual information between $z$ and $x$, we use Distance Correlation (Li et al., 2012) for nonlinear correlations and the Pearson Coefficient (Cohen et al., 2009) with $l_{1}$-norm for linear correlations.
-Additionally, label smoothing and KL divergence (Donsker and Varadhan, 1983) ensure model accuracy, enhancing usability while reducing the separability of $z$.
+Attention mechanisms and multi-receptive field convolutions enhance task-relevant information, while a cloud-based upsampling module preserves accuracy.
+To reduce redundancy, $I(x,z)$ and increase the uncertainty of $z$ to $x$, we use Distance Correlation (Li et al., 2012) for nonlinear correlations and the Pearson Coefficient (Cohen et al., 2009) with $l_{1}$-norm for linear correlations.
+Additionally, Label Smoothing (LS) and Kullback-Leibler (KL) divergence (Donsker and Varadhan, 1983) are used during training to improve model usability while reducing the separability of representations in $z$.
 5.2. The Design of Shallow Edge Model
 Information Funnel.
-Deepening the neural network enhances feature extraction and reduces $I(x,z)$, as confirmed in the Section 4.
-Previous work (Ho et al., 2024) also demonstrates through Fisher information (Rissanen, 1996) that deeper networks improve feature relevance to tasks.
+Deepening the neural network enhances feature extraction and reduces $I(x,z)$, as confirmed in Section 4.
+Previous work (Ho et al., 2024) shows through Fisher information (Rissanen, 1996) that deeper networks improve task relevance and resist MIA.
 Inspired by this, we redesign the final layer of the edge model with an Information Funnel module.
@@
 Adjusting channels based on spatial size is necessary because spatial information loss is irreversible, whereas channel information loss can be compensated in the cloud model.
-While this compensation increases $I(z,y)$, it does not reverse the reduction in $I(x,z)$, ensuring attack difficulty remains high, as demonstrated in Section 6.3.
+This compensation increases $I(z,y)$ without reversing the reduction in $I(x,z)$, maintaining high attack difficulty as shown in Section 6.3.
 So, we add the channel compensation module in the cloud model to guarantee usability.
-Notably, assumption about extracting features from non-independent transmission modules is extremely strong and idealized, requiring the adversary to have full white-box knowledge of the cloud model as if they were the developer.
-Figure 3. Overview of SiftFunnel. Specifically, the design is divided into two aspects: the architecture and the loss function for the edge and cloud model.
-[…]
-3
+Importantly, assumptions about extracting features from non-independent transmission modules are idealized, requiring adversaries to have full white-box knowledge of the cloud model as if they were the developer.
 The Information Funnel’s channel structure controls output but does not limit redundancy or enhance task-relevant information transmission, affecting usability.
@@
 Attention Block.
-To enhance the output of task-relevant information from the Information Funnel, attention mechanism modules are added both before and after it.
-The first module combines an SE block (Hu et al., 2018) and Triplet Attention (Misra et al., 2021).
-The SE block evaluates the contribution of information within each channel, suppressing or nullifying channels with low task relevance.
-Triplet Attention reduces spatial information while maintaining smooth and continuous outputs, preserving fine-grained features suitable for further processing.
-The second attention mechanism integrates SE and CBAM modules (Woo et al., 2018).
-Compared to Triplet Attention, CBAM is a more lightweight module that sequentially calculates channel and spatial weights to enhance or suppress input features, improving network performance.
-Since CBAM’s channel and spatial attention are independently designed, certain regions are suppressed redundantly, making the output features more sparse.
-This is ideal for final output processing, effectively increasing $H(x|z)$ while suppressing $H(z)$ and $\delta(z)$.
+To enhance the output of task-relevant information from the Information Funnel, attention modules are incorporated both before and after it.
+The first block combines an SE block (Hu et al., 2018) and Triplet Attention (Misra et al., 2021).
+The SE block evaluates the contribution of information within each channel, suppressing channels with low task relevance.
+Triplet Attention reduces spatial information while maintaining smooth, continuous outputs, preserving fine-grained features for further processing.
+The second block integrates SE and CBAM modules (Woo et al., 2018).
+Compared to Triplet Attention, CBAM is a lightweight module that sequentially calculates channel and spatial weights to enhance or suppress input features.
+Since CBAM’s channel and spatial attention are independently designed, certain weights are suppressed redundantly, resulting in sparser outputs.
+This sparsity is beneficial for output processing, effectively increasing $H(x\mid z)$ while reducing $H(z)$ and $\delta(z)$.
 5.3. The Design of Loss Function
-To further enhance model usability and data privacy, a appropriately designed loss function is essential to guide the training process.
-To reasonably increase $H(x\mid z)$ and suppress $H(z)$ within the edge model structure—thereby increasing the uncertainty of inferring $x$ from $z$—loss constraints must be applied to $x$ and $z$.
-The key challenge is that, without intervention, neural networks may retain irrelevant input information in their output during forward propagation.
-This information, which does not significantly impact decision-making, reduces the difficulty of attacks by maintaining $I(x,z)$.
-Some defense approaches utilize Constrained Covariance (COCO) (Gretton et al., 2005b) to estimate correlation, focusing primarily on the linear relationship between $x$ and $z$.
-However, since neural networks predominantly perform nonlinear operations, studies have shown that this constraint is not only ineffective but may also lead to reduced accuracy (Peng et al., 2022).
-Alternatively, some approaches employ the Hilbert-Schmidt Independency Criterion (HSIC) (Gretton et al., 2005a) to estimate correlation.
+To further enhance model usability and data privacy, appropriately designed loss constraints must be applied to $x$ and $z$ to reasonably increase $H(x\mid z)$ and suppress $H(z)$ within the edge model structure.
+Without intervention, neural networks may retain irrelevant input information in their output during forward propagation.
+This information, which does not significantly impact decision-making, reduces the difficulty of MIAs by maintaining the mutual information and entropy.
+The key challenge is how to achieve this effectively and flexibly.
+Some defense approaches utilize Constrained Covariance (COCO) (Gretton and Herbrich, 2005) to estimate correlation, focusing primarily on the linear relationship between $x$ and $z$.
+However, since neural networks predominantly perform nonlinear operations, studies have shown that this constraint is not only ineffective but may also lead to reduced usability (Peng et al., 2022).
+Alternatively, some approaches employ the Hilbert-Schmidt Independency Criterion (HSIC) (Gretton et al., 2005) to estimate nonlinear correlation.
 However, kernel-based calculations in HSIC rely heavily on the appropriate selection of $\sigma$.
-Improper $\sigma$ value can introduce bias in nonlinear correlation estimation, limiting the generalizability of this method and reducing its effectiveness across various scenarios.
+Improper $\sigma$ value can introduce bias in correlation estimation, limiting the generalizability of this method and reducing its effectiveness across various scenarios.
 Nonlinear correlation constraint.
 We observe that shallow neural network features retain significant similarity to the input and are derived through nonlinear operations.
-To address this, we use distance correlation to measure the nonlinear relationship between $x$ and $z$, incorporating it as a loss constraint.
-This helps remove redundant input information from the output, effectively increasing $H(x\mid z)$.
+To address this, we use distance correlation to measure the non-linear relationship between $x$ and $z$, incorporating it as a loss constraint.
+This helps remove redundant input information from the output and effectively increase $H(x\mid z)$.
 The calculation of distance correlation is as follows:
 (12) $\mathcal{L}_{\text{d Cor}}=\mathrm{d}\operatorname{Cor}(x,z)=\frac{\mathrm{d}\operatorname{Cov}(x,z)}{\sqrt{\mathrm{d}\operatorname{Var}(x)\cdot\mathrm{d}\operatorname{Var}(z)}}$
-Where $\mathrm{d}\operatorname{Cov}(x,z)$ represents the distance covariance between $x$ and $z$, and the denominator is the product of the distance variances for the two datasets.
+Where $\mathrm{d}\operatorname{Cov}(x,z)$ represents the distance covariance between $x$ and $z$, and the denominator is the product of the distance variances for the two samples.
 To compute the numerator, a distance matrix $A$ is first calculated, where each element is the Euclidean distance between samples, defined as $A=\left\|x_{i}-x_{j}\right\|_{2}$.
@@
 With the centered distance matrix, the distance covariance is computed as $\mathrm{d}\operatorname{Cov}(x,z)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c}B_{ij}^{c}$, where $B$ represents the matrix of features $z$, and $n$ denotes the batch size.
-The distance variances is defined as $\mathrm{d}\operatorname{Var}(x)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c^{2}}$, which normalizes the distance covariance.
+The distance variance is defined as $\mathrm{d}\operatorname{Var}(x)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c^{2}}$, which normalizes the distance covariance.
 The value range of distance correlation is $[0,1]$, where $1$ indicates perfect correlation.
@@
 Moreover, as a distribution-free measure, it does not rely on specific data distributions, making it effective for non-Gaussian distributions, nonlinear relationships, and complex data types such as images and text.
-For non-stationary and multimodal data, distance correlation provides stable dependency measurements, allowing the constraint to flexibly adapt to different models and datasets.
+For non-stationary and multimodal data, distance correlation provides stable dependency measurements, allowing the constraint to flexibly adapt to different models and data.
 Linear correlation constraint.
-In addition to constraining the loss between $x$ and $z$, it is necessary to design loss constraints specifically for the $z$.
-The primary issue lies in the $z$ containing repetitive or similar information, which is redundant for decision-making and provide additional input for Gen-based MIAs, strengthening local constraints in the input space and improving reconstruction accuracy.
-To address this, we propose using the Pearson correlation coefficient as a loss constraint to reduce the linear correlation within features $z$, effectively eliminating redundant information.
+In addition to constraining the correlation between $x$ and $z$, it is necessary to design loss constraints specifically for the $z$.
+The primary issue lies in the $z$ containing repetitive or similar information, which is redundant for decision-making and provides additional input for Gen-based MIAs, strengthening local constraints in the input space and improving reconstruction accuracy.
+To address this, we propose using the Pearson correlation coefficient as a loss constraint to reduce the linear correlation between $x$ and $z$, effectively eliminating redundant information and reducing $\delta(z)$.
 The calculation is as follows:
@@
 This avoids negative values, which could lead to suboptimal updates during gradient descent.
-Minimizing the $\mathcal{L}_{\text{Pearson }}$ in the presence of negative values might push the correlation toward complete negative correlation, introducing unnecessary feature redundancy or unsuitable learning patterns, ultimately preventing convergence.
+with negative values might drive correlations toward complete negative correlation, introducing unnecessary redundancy or unsuitable learning patterns and hindering convergence.
 Therefore, the goal is to minimize the coefficient toward zero, as lower values indicate reduced linear correlation.
@@
 Additionally, to prevent extreme values in the feature output weights during the optimization of the above constraints, we include the $l_{1}$-norm as a loss term.
-This term is added specifically to suppress weights, but excessive emphasis on it could interfere with normal model training.
+This term is added specifically to suppress weights, but excessive emphasis could interfere with normal model training.
 Therefore, the coefficient $\tau$ for this loss term is set below 1e-2 to balance its influence.
 Label smoothing.
-Finally, to ensure model accuracy, we draw inspiration from (Struppek et al., 2023) and apply label smoothing to refine the target loss, as shown in the formula (14).
+Finally, to ensure model usability, we draw inspiration from (Struppek et al., 2023) and apply LS to refine the target loss, as shown in formula (14).
 Additionally, KL divergence is used instead of cross-entropy (CE) as the primary loss function for optimization.
@@
 $\alpha$ is smooth factor and $K$ is the total number of categories.
-This smoothing mechanism ensures that the loss value computed by cross-entropy emphasizes enhancing inter-class feature distinctions, while KL divergence as a loss function supports smoother and more generalized feature extraction between classes (Struppek et al., 2023).
-Within-class feature extraction focuses on capturing universal characteristics, ensuring the model maintains a certain level of generalization and robustness for the target task.
-Moreover, this optimization reduces maximum feature separability, increasing the difficulty of attacks, as demonstrated in Section 6.
-In summary, the proposed loss function first applies nonlinear correlation constraints between input $x$ and features $z$ to increase $H(x\mid z)$ and suppress $H(z)$.
-Second, it imposes linear correlation and extreme value constraints on $z$ to further enhance $H(x\mid z)$ and suppress $\delta(z)$.
-Finally, label smoothing is applied to the target loss to maintain accuracy while slightly suppressing $H(z)$.
-The final formulation of this optimization is as follows, with specific parameter settings referenced in Section 6.
+CE enhances inter-class feature distinctions, while KL divergence supports smoother and more generalized feature extraction between classes (Struppek et al., 2023).
+Within-class feature extraction captures universal characteristics, ensuring generalization and robustness for the target task.
+This optimization reduces feature separability, increasing the difficulty of attack, as demonstrated in Section 6.3.
+In summary, the proposed loss function first applies nonlinear correlation constraints to increase $H(x|z)$ and suppress $H(z)$.
+Next, it applies linear correlation and extreme value constraints on $z$ to further enhance $H(x\mid z)$ and suppress $\delta(z)$.
+Finally, LS and KL divergence maintain usability while slightly suppressing $H(z)$.
+The final formulation of this optimization is as follows, with specific parameter settings.
+The final formulation of this optimization is as follows, with specific parameter settings referenced in Section 6.1.
+SiftFunnel and MINE are available at https://github.com/SiftFunnel/SiftFunnel.
 (15) $\min\mathcal{L}_{\theta}(x,z,\hat{y})=$
@@
 6. Experiments
-In this section, we systematically evaluate the resistance to MIA and compare our method with existing approaches.
-The results are analyzed from three perspectives: the impact of defense methods on $I(x\mid z)$ and $\delta(z)$, the influence of datasets and model architectures, and ablation studies.
-The specific experimental setup and evaluation metrics are shown below.
+In this section, we systematically evaluate and compare the ability of SiftFunnel and existing approaches to resist MIA. The specific experimental setup and evaluation metrics are shown below.
 6.1. Experimental Setup
@@
   • CIFAR-10 (Krizhevsky et al., 2009).
-The CIFAR-10 dataset consists of 60,000 RGB images categorized into ten classes (e.g., airplane, ship, etc.) with an original resolution of $32\times 32$.
-To make the results clearer and more intuitive, we used an upscaled resolution of $64\times 64$.
+The CIFAR-10 dataset consists of 60,000 RGB images categorized into ten classes with an original resolution of $32\times 32$.
+To make the results clearer and more intuitive, we used an unscaled resolution of $64\times 64$.
   • FaceScrub (Ng and Winkler, 2014).
@@
   • CelebA (Liu et al., 2015).
-CelebA is a dataset with 202,599 images of 10,177 celebrities from the Internet.
+CelebA is a dataset with 202,599 images of 10,177 celebrities.
 We used the same crop as (Yang et al., 2019) to remove the background of images in this dataset other than faces to reduce the impact on the experiment.
-In order to eliminate individual overlap, we removed a total of 6,878 images of 296 individuals from CelebA and similarly resized the images to $64\times 64$.
+To eliminate individual overlap, we removed a total of 6,878 images of 296 individuals and similarly resized the images to $64\times 64$.
   • ChestX-ray (Sait et al., 2020).
 This dataset is a curated collection of COVID-19 chest X-ray images compiled from 15 publicly available datasets.
-It contains 1,281 COVID-19 X-rays, 3,270 normal X-rays, 1,656 viral pneumonia X-rays, and 3,001 bacterial pneumonia X-rays.
-For our analysis, we converted the images to gray scale with a resolution of $128\times 128$.
+It contains 1,281 COVID-19 X-rays, 3,270 normal X-rays, 1,656 viral pneumonia X-rays, and 3,001 bacterial pneumonia X-rays with a resolution of $128\times 128$.
 Table 1. Data allocation of the target model and attack model.
-Classifier Attack Model
+Target model Attack Model
 Task Data Auxiliary Data
@@
 ChestX-ray (4 classes) Train: 66.6%, Test: 16.7% 16.7% (4 classes of ChestX-ray)
-2) Attack Method.
-The selection of attack methods has been discussed in Section 3.2.
-Specifically, for MLE-based MIA, this paper has adjusted to the hyperparameter values that yield the best results.
-For Gen-based MIA, the attack model in this paper is designed as the inverse architecture of the edge model.
-Moreover, under the gray-box scenario assumption, the attack model can enhance the expressive power of the reverse mapping as the edge model architecture changes.
-The attack model is also trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ of 0.5, and a ReduceLROnPlateaud scheduler with a factor of 0.5 and patience of 20.
-The training conditions for the attack models in this paper have been done with strong assumptions and only the attack models where the test Mean Square Error (MSE) reaches the minimum value are saved to highlight the effectiveness of the defense efforts.
-3) Target Edge Model and Implementation Details.
-We employ the same comprehensive model architecture as detailed in (Yang et al., 2019), comprising four Convolutional Neural Network (CNN) blocks succeeded by two fully connected layers.
-This model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ parameter of 0.5, and a ReduceLROnPlateaud scheduler with a reduction factor of 0.5 and a patience setting of 25.
-Each CNN block is composed of a convolutional layer, a batch normalization layer, a max-pooling layer, and a ReLU activation function.
-The target edge model encompasses the initial two CNN blocks.
-For the training of the edge model protected by SiftFunnel, we have specified the parameters $\alpha$ to 3.5, $\epsilon$ to 0.35, $\beta$ to 0.8, and $\gamma$ to 0.6 in Equation (15).
-In addition to this, to argue the impact of the model architecture for the defense, we use different CNNs as follows: (1) VGG16 adapted from (Simonyan, 2014); (2) ResNet-18 adapted from (He et al., 2016), the learning rates are 3e-4.
-The division positions of resnet and vgg are not fixed.
-For detailed accuracy and how to divide them, see Section 6.3 for analysis.
-All experiments were performed on two RTX 4090 GPUs and an Intel Core i9-14900KF ×$\times 32$ CPU.
-4) Comparison of Defense Methods.
-We will compare three existing categories of defense methods.
-For Perturbation-based defenses, we select Adding noise & Dropping features (He et al., 2020), DPSGD (Abadi et al., 2016), and negative label smoothing (NLS) (Struppek et al., 2023).
-For IB-based defenses, we choose MID (Wang et al., 2021), BiDO (with HSIC) (Peng et al., 2022), VIB (Wang et al., 2024), and InfoSCISSORS (Duan et al., 2023).
-For NND & IB-based defenses, we opt for AE-based protection (Azizian and Bajić, 2024).
-Our method falls into the NND & IB-based defense category, which also requires analyzing the defensive effect in a gray-box scenario.
-Furthermore, all comparison methods have been selected and set with appropriate hyper-parameters for comparison, with detailed settings found in Section 6.3 analysis.
-Lastly, all methods are trained without adversarial training, focusing solely on the effects of the original defense methods.
-6.2. Evaluation Metrics
-We have selected 7 evaluation metrics to assess usability, privacy, and deployability.
-For usability assessment, we employ Test Accuracy (Test ACC), which evaluates the final performance of the model protected by defense techniques on the test dataset.
-Regarding privacy assessment, we initially utilize MSE, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) to evaluate the effectiveness of MIA on reconstructing the input data of the protected model (Yin et al., 2023).
-Here, a higher MSE and lower PSNR and SSIM (ranging from 0 to 1) indicate better defense effectiveness.
-Compared to the former two, SSIM focuses more on the perceptual quality of the image, encompassing luminance, contrast, and structural information.
-In addition to these, we also employ MINE (Belghazi et al., 2018), as mentioned in Section  4.3, to assess Mutual Information (MI) (ranging from 0 to 1) and the Effective Information Mean $\delta(z)$, which is the mean of the non-zero information output by the edge model on the test dataset.
-The specific training methods and model architecture for MINE are referenced from xxx.
-For deployability, we consider the Edge model’s parameter count $|\theta_{\text{edge}}|$ and Inference Latency.
-The model parameter count directly determines memory or GPU memory occupation, while inference latency is tested on both CPU and GPU.
-6.3. Experimental Results
 Table 2.
@@
 Blue highlights show unprotected baseline performance, red indicates increased edge device load, yellow indicates reduced load, and gray shows gray-box defense performance.
-MLE-based Attack (He et al., 2016) Gen-based Attack (He et al., 2020)
+MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
 Method Class Method Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
@@
 4
-Table 3.
-Quantitative evaluation of the resistance to MIA across different architectures on CIFAR-10.
-The edge model divisions for CNN and ResNet-18 follow the configurations in Tables 2 and 4, while for VGG16, the edge model is set at the first ReLU layer after a max-pooling operation.
-MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
-Architecture Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
-0.1166 14.1112 0.5419
-CNN 85.49% (-2.79%) 0.6792 6.4510 0.0051 0.0639 16.7314 0.7647 0.0167 1,006 14,911
-0.1957 11.8562 0.0793
-ResNet-18 88.43% (-2.00%) 0.6538 6.6166 0.0059 0.2790 10.3230 0.3001 0.4772 4,028 78,864
-0.0827 15.5928 0.6559
-VGG16 88.89% (-2.18%) 0.7801 5.8450 0.0005 0.0833 15.5916 0.7181 0.0128 3,587 52,086
+2) Attack Method.
+The selection of attack methods has been discussed in Section 3.2.
+Specifically, for MLE-based MIA, this paper has adjusted to the hyper-parameter values that yield the best results.
+For Gen-based MIA, the attack model in this paper is designed as the inverse architecture of the edge model.
+Moreover, under the gray-box scenario assumption, the attack model can enhance the expressive power of the reverse mapping as the edge model architecture changes.
+The attack model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ of 0.5, and a ReduceLROnPlateaud scheduler with a factor of 0.5 and patience of 20.
+The attack model training is based on strong assumptions, as shown in Table 1, and only the models with the minimum test MSE are saved to highlight the defense’s effectiveness.
+3) Target Edge Model and Implementation Details.
+We employ the same comprehensive model architecture as detailed in (Yang et al., 2019), comprising four CNN blocks succeeded by two fully connected layers.
+This model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ parameter of 0.5, and a ReduceLROnPlateaud scheduler with a reduction factor of 0.5 and patience of 25.
+Each CNN block is composed of a convolutional layer, a batch normalization layer, a max-pooling layer, and a ReLU activation function.
+The target edge model encompasses the initial two CNN blocks.
+For the training of the edge model protected by SiftFunnel, we have specified the parameters $\lambda_{1}$ to 3.5, $\alpha$ to 0.35, $\lambda_{2}$ to 0.8, and $\lambda_{3}$ to 0.6 in Equation (15).
+In addition, to argue the impact of the model architecture for the defense, we use different CNNs as follows: (1) VGG16 adapted from (Simonyan, 2014); (2) ResNet-18 adapted from (He et al., 2016), the learning rates are 3e-4.
+The division positions of ResNet and VGG16 are not fixed.
+For detailed accuracy and how to divide them, see Section 6.3 for analysis.
+All experiments were performed on two RTX 4090 GPUs and an Intel Core i9-14900KF ×$\times 32$ CPU.
+4) Comparison of Defense Methods.
+We aim to compare three existing categories of defense methods.
+For Perturbation-based defenses, we select Adding noise & Dropping features (He et al., 2020), DPSGD (Abadi et al., 2016), and negative label smoothing (NLS) (Struppek et al., 2023).
+For IB-based defenses, we choose MID (Wang et al., 2021), BiDO (with HSIC) (Peng et al., 2022), VIB (Wang et al., 2024), and InfoSCISSORS (Duan et al., 2023).
+For NND & IB-based defenses, we opt for AE-based protection (Azizian and Bajić, 2024).
+Our method falls into the NND & IB-based defense category, which also requires analyzing the defensive effect in a gray-box scenario.
+Furthermore, all comparison methods have been selected and set with appropriate hyper-parameters for comparison, with detailed settings found in Section 6.3.
+Lastly, all methods are trained without adversarial training, focusing solely on the effects of the original defense methods.
+6.2. Evaluation Metrics
+We selected 7 evaluation metrics to assess the aspects of usability, privacy, and deployability.
+For usability assessment, we employ Test Accuracy (Test ACC), which evaluates the final performance of the model protected by defense techniques on the test dataset.
+Regarding privacy assessment, we initially utilize MSE, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) to evaluate the effectiveness of MIA on reconstructing the input data of the protected model (Yin et al., 2023).
+Here, a higher MSE and lower PSNR and SSIM (ranging from 0 to 1) indicate better defense effectiveness.
+In addition, we employ MINE (Belghazi et al., 2018), as mentioned in Section  4.3, to assess Mutual Information (MI) (ranging from 0 to 1) and the Effective Information Mean $\delta(z)$, which is the mean of the non-zero output by the edge model on the test dataset.
+The specific training methods and model architecture for MINE are referenced from https://github.com/SiftFunnel/SiftFunnel.
+For deployability, we consider the Edge model’s parameter count $|\theta_{\text{edge}}|$ and Inference Latency.
+The model parameter count directly determines memory or GPU memory occupation, while inference latency is tested on both CPU and GPU.
 Table 3.
@@
 The meanings of color-coded regions are consistent with those in Table 2.
-MLE-based Attack (He et al., 2016) Gen-based Attack (He et al., 2020)
+MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
 Method Class Method Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
@@
 NND & IB-based SiftFunnel ($\lambda_{1}=5.0,\lambda_{2}=0.1,\lambda_{3}=0.1$) 85.39% (-2.92%) 0.7264 6.1593 0.0060 0.0296 20.0576 0.8323 0.0104 637 14,911
+6.3. Experimental Results
+The results are analyzed from 4 perspectives: the impact of defense methods on $I(x,z)$ and $\delta(z)$, the influence of datasets and model architectures, and ablation studies.
 1) The effect of various defenses on the MI and $\delta(z)$.
-the effectiveness of defenses is reflected by MI and $\delta(z)$, aligning with the criterion $D_{\text{mia}}$ and the theoretical analysis presented in Section 4.
-Specifically, unprotected edge models show significant vulnerability to input reconstruction under both types of MIAs, with MI at 0.0566, MSE below 0.01, and SSIM exceeding 0.98.
-When protected by Perturbation-based defenses, all methods except NLS exhibit over a 10% drop in Test ACC.
-Moreover, these methods achieve only an 11% reduction in MI at most, demonstrating clear weaknesses against both MIAs.
-NLS shows a slight increase in mutual information for shallow neural networks, leading to weaker resistance to both types of attacks.
-Under IB-based defenses, usability sees a slight decline, but all methods except MID provide strong resistance against MLE-based MIA, with MSE values exceeding 0.5.
-Among these techniques, only BiDO significantly impacts MI and $\delta(z)$ , providing effective defense, but its parameter selection and kernel width configuration (as shown in the table) are challenging, limiting its generalizability—a limitation further evidenced in subsequent experiments.
-This experiment adheres to the InfoSCISSORS setup, providing CLUB with a complete dataset and sufficient training iterations to ensure convergence.
-While CLUB effectively defends against MLE-based MIA, it struggles under the experimental conditions to resist Gen-based MIA and results in a 4.83% accuracy loss.
-Under AE-based defenses, target model usability improves, likely because the original model lacked sufficient capacity for optimal task classification, and the added model complexity enhanced performance.
-Additionally, MI indicates notable reductions in mutual information, showcasing reasonable defense effectiveness.
-However, in gray-box scenarios, AE-based defenses show reduced resistance, and the added complexity imposes a eightfold storage burden on edge models, as revealed by $|\theta_{\text{edge}}|$.
-SiftFunnel achieves the best defense performance against both MIAs while maintaining an accuracy loss below 3%.
-For MLE-based MIA, it achieves an MSE of 0.6792, and for Gen-based MIA, an MSE above 0.1.
-In gray-box scenarios, it maintains an MSE of 0.06 and an SSIM of 0.7647, ensuring robust resistance.
-Furthermore, MI and $\delta(z)$ demonstrate that SiftFunnel exerts optimal influence on the key elements of $D_{\text{mia}}$.
-Additionally, $|\theta_{\text{edge}}|$ reveals that SiftFunnel reduces the edge model’s load by nearly 20 times.
-2) The effect of model architecture.
-Changing the model architecture does not affect the parameter settings of the proposed method, which remain consistent with Section 6.1.
-We adapt the model to a ResNet-18 with skip connections, using the first residual block as the edge model.
-This setup is challenging for defense techniques, as shallow networks are simple and do not alter spatial dimensions.
-Table 4 shows minimal MI impact in unprotected models, leading to weak resistance against MIAs, with MLE-based MIA achieving an MSE of $10^{-4}$ and Gen-based MIA $10^{-3}$.
-Under these conditions, existing perturbation-based defenses fail to balance usability and privacy effectively, resulting in poor defense performance.
-Among IB-based defenses, only VIB demonstrates resistance to MLE-based MIA by leveraging the estimation of $I(x,z)$ as a loss to optimize the edge model.
-However, its effectiveness remains limited, particularly against Gen-based MIA.
-For BiDO, despite following the kernel bandwidth estimation methods from (Peng et al., 2022) and performing extensive experimental tuning, it fails to provide a configuration that balances usability and privacy, ultimately leading to suboptimal defense outcomes.
-InfoSCISSORS demonstrates some impact on MI under this configuration, but the increase in $\delta(z)$ undermines its defense effectiveness against Gen-based MIA, showing only marginal improvements.
-AE-based defenses exhibit certain resistance to attacks on low-dimensional features in earlier stages, but under the ResNet configuration, their performance diminishes.
-While this method reduces $\delta(z)$ by 60%, it has minimal impact on MI, with only a 0.04% decrease.
-Consequently, AE-based defenses struggle to counter both black-box and gray-box Gen-based MIAs effectively.
+As shown in Table 2, defense effectiveness is reflected by MI and $\delta(z)$, aligning with the criterion $D_{\text{mia}}$.
+Specifically, unprotected edge models are highly vulnerable to MIAs, with MI at 0.0566, MSE below 0.01, and SSIM exceeding 0.98.
+Perturbation-based defenses reduce MI by at most 11% but cause over a 10% drop in Test ACC, with NLS even increasing MI in shallow networks, weakening resistance to attacks.
+UIB-based defenses slightly reduce usability but effectively counter MLE-based MIAs, except for MID, which shows weaker resistance.
+Among them, BiDO significantly impacts MI and $\delta(z)$, providing effective defense, but its parameter selection and kernel width configuration (as shown in the table) limits generalizability, as confirmed in subsequent experiments.
+This experiment follows the InfoSCISSORS setup, ensuring CLUB receives a complete dataset and sufficient training for convergence.
+While CLUB effectively defends against MLE-based MIAs, it struggles against Gen-based MIAs and incurs a 4.83% accuracy loss.
+AE-based defenses improve model usability, likely due to enhanced model capacity, while MI reductions confirm reasonable defense effectiveness.
+However, gray-box scenarios weaken AE-based defenses, and their added complexity increases storage costs by 8×8$\times$8 ×, as seen in $|\theta_{\text{edge}}|$.
+SiftFunnel achieves the best defense performance against both MIAs while keeping accuracy loss below 3%.
+It achieves MSE of 0.6792 (MLE-based MIA) and above 0.1 (Gen-based MIA).
+In gray-box scenarios, it maintains MSE of 0.06 and SSIM of 0.7647, ensuring robust resistance.
+Furthermore, MI and $\delta(z)$ confirm its optimal impact on the key factors of $D_{\text{mia}}$.
+Additionally, $|\theta_{\text{edge}}|$ reveals a nearly 20×20$\times$20 × reduction in edge model load, highlighting SiftFunnel’s efficiency.
 Table 5.
@@
 SiftFunnel 88.43% (-2.00%) 0.2790 10.3230 0.3001 0.4772 4,028
-In contrast, SiftFunnel reduces MI by 50% and lowers $\delta(z)$ to the millesimal scale, achieving the best resistance against MIAs while maintaining usability and deployability.
-Notably, it achieves a minimum SSIM of 0.0793, highlighting its superior performance in balancing privacy and usability.
+2) The effect of model architecture.
+Modifying the model architecture does not alter the parameter settings of the proposed method, which remain consistent with Section 6.1.
+We adapt the model to ResNet-18 with skip connections, using the first residual block as the edge model.
+This setup poses challenges for defense techniques, as shallow networks retain simple structures and do not modify spatial dimensions.
+Table 3 shows that unprotected models exhibit minimal MI impact, resulting in weak resistance to MIAs, with MLE-based MIA achieving an MSE of $10^{-4}$ and Gen-based MIA $10^{-3}$.
+Under these conditions, perturbation-based defenses fail to balance usability and privacy, leading to poor defense performance.
+Among IB-based defenses, only VIB demonstrates resistance to MLE-based MIA by optimizing the edge model through $I(x,z)$ estimation as a loss function, but its effectiveness remains limited, particularly against Gen-based MIA.
+For BiDO, despite applying kernel bandwidth estimation methods from (Peng et al., 2022) and extensive experimental tuning, it fails to achieve a configuration that effectively balances usability and privacy, resulting in suboptimal defense performance.
+InfoSCISSORS impacts MI under this configuration, but the increase in $\delta(z)$ undermines its effectiveness against Gen-based MIAs, leading to only marginal improvements.
+AE-based defenses show some resistance to attacks on low-dimensional features in earlier stages, but their effectiveness diminishes under the ResNet configuration.
+While they reduce $\delta(z)$ by 60%, they have minimal impact on MI, achieving only a 0.04% reduction.
+Consequently, AE-based defenses struggle to counter both black-box and gray-box Gen-based MIAs effectively.
+Table 6.
+Quantitative evaluation of the resistance to MIA across different architectures on CIFAR-10.
+The edge model divisions for CNN and ResNet-18 follow the configurations in Tables 2 and 3, while for VGG16, the edge model is set at the first ReLU layer after a max-pooling operation.
+MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
+Architecture Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
+0.1166 14.1112 0.5419
+CNN 85.49% (-2.79%) 0.6792 6.4510 0.0051 0.0639 16.7314 0.7647 0.0167 1,006 14,911
+0.1957 11.8562 0.0793
+ResNet-18 88.43% (-2.00%) 0.6538 6.6166 0.0059 0.2790 10.3230 0.3001 0.4772 4,028 78,864
+0.0827 15.5928 0.6559
+VGG16 88.89% (-2.18%) 0.7801 5.8450 0.0005 0.0833 15.5916 0.7181 0.0128 3,587 52,086
+In contrast, SiftFunnel reduces MI by 50% and lowers $\delta(z)$ to the millesimal scale, delivering the strongest resistance against MIAs while preserving usability and deployability.
+Notably, it achieves a minimum SSIM of 0.0793, underscoring its superior ability to balance privacy and usability.
+The different protection performances of the three model architectures with the same dataset are shown in Table 6.
 3) The effect of data type.
@@
 Specifically, the division points were set after the second residual block for ResNet and after the first ReLU layer following the second max-pooling layer for VGG.
-The results, summarized in the table 5, show that the defense performance remains unaffected by changes in data types.
+The results, summarized in Table 4 and Figure 5, show that the defense performance remains unaffected by changes in data types.
 SiftFunnel continues to be the best defense method, offering an optimal balance between usability, privacy, and deployability.
+As shown in Figure 5, the VGG16 model protected by SiftFunnel demonstrates better resistance to Gen-based MIA in terms of the visual effects of usability and privacy on facial and X-ray data, with PSNR values consistently around 15.
 4) Ablation Study.
-To validate the effectiveness of SiftFunnel’s model structure and loss function design, as well as its defense capability, we conducted ablation experiments using Gen-based MIA.
+To assess the effectiveness of SiftFunnel’s model structure, loss function design, and defense capability, we conducted ablation experiments using Gen-based MIA.
 As shown in Table 5, we separately removed the Information Funnel and Attention Blocks from the model structure.
-The results indicate that removing the Funnel had minimal impact on accuracy, with only a 0.1% decrease, while removing the attention mechanism led to a significant increase in MI compared to SiftFunnel.
-For the loss function, we evaluated the impact of individually removing KL divergence, linear and nonlinear constraints, and the $l_{1}$-norm.
-The results demonstrate that KL divergence and $l_{1}$-norm moderately improved defense, whereas nonlinear constraints played a critical role; removing them increased the MSE to 0.0268.
-Additionally, we verified the hypothesis from Section 5.2, showing that even when adversaries reconstructed input using features from the cloud’s compensation module, SiftFunnel outperformed existing defenses in protecting transmitted features.
+The results indicate that removing the Funnel had minimal impact on accuracy, reducing it by only 0.1%, whereas removing the attention mechanism significantly increased MI, weakening privacy protection.
+For the loss function, we examined the effect of removing KL divergence, linear and nonlinear constraints, and the $l_{1}$-norm.
+The results show that KL divergence and $l_{1}$-norm provided moderate defense improvements, while nonlinear constraints were critical—removing them increased MSE to 0.0268.
+Additionally, we validated the hypothesis from Section 5.2, demonstrating that even when adversaries reconstructed inputs using cloud-compensated features, SiftFunnel still outperformed existing defenses in protecting transmitted features.
+Figure 5.
+Quantitative and visual evaluation of privacy and usability against Gen-based MIA in the VGG16 edge model on the ChestX-ray ($\lambda_{1}=3.5,\lambda_{2}=1.2,\lambda_{3}=0.8$) and FaceScrub ($\lambda_{1}=4.5,\lambda_{2}=0.4,\lambda_{3}=0.1$).
+[…]
+5
 We also evaluated the processing speed of the edge model on both CPU and GPU under this configuration.
@@
 While SiftFunnel showed slightly higher latency on GPU, it required significantly less memory, enabling more parallel processing of inputs under the same resource constraints.
-7. Discussion
-Systematic experiments validated both the criterion and the effectiveness of the method.
-However, our analysis revealed that achieving optimal resistance against attacks remains challenging for current defense methods, including SiftFunnel, particularly when applied to shallow edge neural networks.
-When edge models perform only basic feature processing, even SiftFunnel can still allow Gen-based MIA to reconstruct blurred feature structures, while other methods are more vulnerable to input reconstruction.
-Enhancing privacy protection for extremely shallow edge neural networks presents a promising direction for future research.
-Additionally, using MI as an evaluation metric, we observed that some defense methods result in only a 0.5% reduction in MI between inputs and features, yet current MLE-based MIAs still fail to bypass these defenses.
-This is directly linked to the lack of advancement in existing attack methods.
-Based on our experimental findings, we suggest that further exploration of more powerful MIAs in white-box scenarios could be a potential research direction.
 7. Conclusion
-The increasing complexity of neural networks and inference tasks, alongside the demand for efficiency and real-time feedback, presents significant challenges for edge devices with limited resources.
-Collaborative inference offers a solution by delegating feature extraction to edge devices and offloading subsequent tasks to the cloud, but this process leaves transmitted features vulnerable to MIAs.
-To address these vulnerabilities, we established a first criterion $D_{\text{mia}}$ for evaluating MIA difficulty and proposed SiftFunnel, a privacy-preserving framework designed to limit redundant information and enhance usability.
-By integrating linear and non-linear correlation constraints, label smoothing, and a funnel-shaped edge model with attention mechanisms, SiftFunnel achieves robust privacy protection while maintaining usability and deployability.
-Experimental results demonstrate its effectiveness, achieving superior defense against MIAs with minimal accuracy loss and a balanced trade-off among usability, privacy, and practicality.
+While CI enhances efficiency, it also exposes transmitted features to MIAs, posing severe privacy risks.
+Despite extensive research, a fundamental understanding of the factors determining MIA effectiveness remains lacking, and existing defenses struggle to balance usability, privacy, and deployability, facing theoretical, methodological, and practical constraints.
+To address these challenges, we proposed the first criterion for evaluating MIA difficulty in CI, identifying MI, entropy, and effective information volume as key influencing factors.
+Based on this, we proposed SiftFunnel, a privacy-preserving framework designed to limit redundant information while maintaining usability and deployability.
+By integrating linear and nonlinear correlation constraints, label smoothing, and a funnel-shaped edge model with attention mechanisms, SiftFunnel achieved robust privacy protection with an average accuracy loss of around 3%, and revealed a nearly 20×20$\times$20 × reduction in edge burdens.
+Experimental results demonstrated that SiftFunnel offers superior privacy protection across various evaluation metrics and achieves an optimal balance among usability, privacy, and practicality.

### v5

@@
 How Breakable Is Privacy: Probing and Resisting Model Inversion Attacks in Collaborative Inference
-Abstract.
+Abstract
 Collaborative inference (CI) improves computational efficiency for edge devices by transmitting intermediate features to cloud models.
@@
 Experiments show that, compared to state-of-the-art defense, SiftFunnel increases reconstruction error by $\sim$30%, lowers mutual and effective information metrics by $\geq$50%, and reduces edge burdens by almost 20×20$\times$20 ×, while maintaining comparable usability.
-Collaborative inference, model inversion attacks, deep learning security and privacy, information theory
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference on Computer and Communications Security; October 13–17, 2025; xx, xxisbn: 978-1-4503-XXXX-X/25/10ccs: Security and privacy Domain-specific security and privacy architectures
-1. Introduction
+1 Introduction
 The rapid advancement of deep learning has led to increasingly complex neural networks that achieve remarkable performance in tasks like image recognition and object detection.
-However, deploying such models on edge devices remains challenging due to limited computational resources (Huckelberry et al., 2024).
-Applications like autonomous driving, facial recognition, IoT surveillance, and remote healthcare require real-time processing, yet the limited computational capacity of edge devices makes standalone deep learning inference impractical (Shlezinger and Bajić, 2022).
+However, deploying such models on edge devices remains challenging due to limited computational resources [1].
+Applications like autonomous driving, facial recognition, IoT surveillance, and remote healthcare require real-time processing, yet the limited computational capacity of edge devices makes standalone deep learning inference impractical [2].
 Cloud computing provides substantial computational power but faces latency, bandwidth, and privacy concerns when transmitting raw data.
-Collaborative inference (CI) (Shlezinger et al., 2021) mitigates these issues by partitioning neural networks into an edge model $f_{\text{edge}}$ and a cloud model $f_{\text{cloud}}$, where $f_{\text{edge}}$ extracts feature representations for $f_{\text{cloud}}$ to inference.
-This method enhances efficiency while maintaining real-time performance and is widely applied in UAVs (Qu et al., 2024), IoT systems (Shlezinger and Bajić, 2022), and private cloud computing (PCC) (Engineering and Architecture, 2024).
+Collaborative inference (CI) [3] mitigates these issues by partitioning neural networks into an edge model $f_{\text{edge}}$ and a cloud model $f_{\text{cloud}}$, where $f_{\text{edge}}$ extracts feature representations for $f_{\text{cloud}}$ to inference.
+This method enhances efficiency while maintaining real-time performance and is widely applied in UAVs [4], IoT systems [2], and private cloud computing (PCC) [5].
 With the rise of 5G, IoT, and AI, CI is poised for broader adoption.
-However, recent studies have revealed significant privacy risks in CI, where input data on edge devices can be reconstructed by model inversion attacks (MIAs) (He et al., 2019; Yang et al., 2022; Yin et al., 2023), as shown in Figure 1.
-Unlike centralized deployments, where adversaries recover training data through model APIs (Yuan et al., 2023; Struppek et al., 2022), CI often relies on shallow neural networks at the edge, producing feature representations that retain significant mutual information (MI) with input data (Azizian and Bajić, 2024).
-This exposes raw inputs to adversaries, who can reconstruct sensitive data, such as facial images directly from transmitted features (Ding et al., 2024).
-MIAs in CI can be categorized into maximum likelihood estimati-on-based (MLE-based) MIAs (He et al., 2019, 2020) and generative model-based (Gen-based) MIAs (Yin et al., 2023; He et al., 2020).
+However, recent studies have revealed significant privacy risks in CI, where input data on edge devices can be reconstructed by model inversion attacks (MIAs) [6, 7, 8], as shown in Figure 1.
+Unlike centralized deployments, where adversaries recover training data through model APIs [41, 53], CI often relies on shallow neural networks at the edge, producing feature representations that retain significant mutual information (MI) with input data [15].
+This exposes raw inputs to adversaries, who can reconstruct sensitive data, such as facial images directly from transmitted features [20].
+MIAs in CI can be categorized into maximum likelihood estimati-on-based (MLE-based) MIAs [6, 16] and generative model-based (Gen-based) MIAs [8, 16].
 MLE-based attacks optimize input data to minimize the loss between the edge model’s output and transmitted features but require white-box access, limiting practicality.
@@
 These privacy risks have driven extensive research on attacks and defenses, but existing methods still face notable limitations.
-Figure 1. Collaborative inference and the threat of model inversion attacks.
-[…]
-Collaborative inference and the threat of model inversion attacks.
+Figure 1: Collaborative inference and the threat of model inversion attacks.
 Firstly, existing research lacks a comprehensive analysis of the fundamental factors influencing MIA effectiveness in CI.
 Current attack methods focus on improving data reconstruction under specific assumptions, leveraging innovative techniques to enhance reconstruction fidelity and generalization.
-In recent years, defenses have increasingly recognized the role of mutual information in MIAs and have shifted toward designing optimal information bottleneck (IB) approaches (Wang et al., 2024; Peng et al., 2022) to counter MIAs.
+In recent years, defenses have increasingly recognized the role of mutual information in MIAs and have shifted toward designing optimal information bottleneck (IB) approaches [19, 18] to counter MIAs.
 However, these methods remain a methodological extension based on existing findings rather than advancing a fundamental theoretical understanding of MIAs.
@@
 Secondly, the trade-off between usability, privacy, and deployability in edge models is an underexplored and challenging area in current research.
-Encryption-based defenses receive little attention in CI due to resource constraints, real-time demands (Mireshghallah et al., 2021), and vulnerability to side-channel attacks (Hu et al., 2024).
-Mainstream MIA defense strategies in CI, constrained by a limited understanding of MIA effectiveness factors, primarily focus on reducing the mutual information $I(x,z)$ between inputs $x$ and extracted features $z$ to counter MIAs (Wang et al., 2024; Peng et al., 2022; Azizian and Bajić, 2024).
-However, mutual information estimation is inherently challenging due to its reliance on accurate probability density estimation, large-scale training data, and the computational overhead associated with optimization techniques (Wang et al., 2024; Duan et al., 2023), making the usability-privacy trade-off inefficient and difficult to guarantee.
-Moreover, some approaches compromise usability (He et al., 2020), increase computational and storage burdens on edge devices (Wang et al., 2024), and complicate further fine-tuning and optimization.
+Encryption-based defenses receive little attention in CI due to resource constraints, real-time demands [22], and vulnerability to side-channel attacks [23].
+Mainstream MIA defense strategies in CI, constrained by a limited understanding of MIA effectiveness factors, primarily focus on reducing the mutual information $I(x,z)$ between inputs $x$ and extracted features $z$ to counter MIAs [19, 18, 15].
+However, mutual information estimation is inherently challenging due to its reliance on accurate probability density estimation, large-scale training data, and the computational overhead associated with optimization techniques [19, 34], making the usability-privacy trade-off inefficient and difficult to guarantee.
+Moreover, some approaches compromise usability [16], increase computational and storage burdens on edge devices [19], and complicate further fine-tuning and optimization.
 Consequently, such methods are often less favored in practical applications.
@@
 Challenge 3 (Practical constraint): How to balance usability and privacy while ensuring practical deployability?
-Certain defense strategies (Azizian and Bajić, 2024) inevitably increase computational and storage burdens on edge devices.
+Certain defense strategies [15] inevitably increase computational and storage burdens on edge devices.
 Furthermore, when applied to edge models designed for simple feature extraction, these methods struggle to counter advanced Gen-based MIAs effectively.
@@
 (1) Criterion basis: SiftFunnel offers interpretable defense principles based on $D_{mia}$ We analyze MIAs using formulaic principles, information theory, and Fano’s inequality, and derive an evaluation formula involving 4 critical factors: mutual information $I(x,z)$, conditional entropy $H(x\mid z)$, feature entropy $H(z)$, and effective information mean $\delta(z)$ (the average number of non-zero elements in $z$).
-Their impact on MIA performance is validated via the Mutual Information Neural Estimator (MINE) (Belghazi et al., 2018).
+Their impact on MIA performance is validated via the Mutual Information Neural Estimator (MINE) [24].
 (2) Usability-privacy trade-off: SiftFunnel effectively reduces redundant information while preserving edge model usability.
@@
 Our contributions are summarized as follows.
-  • We propose the first MIA difficulty criterion in CI, identifying $I(x,z)$, $H(x|z)$, $H(z)$ and $\delta(z)$ as factors impacting attack effectiveness, deepening insights into MIAs.
-  • We propose SiftFunnel, a privacy-preserving scheme for CI that constrains redundant transmitted features through linear and nonlinear correlation, ensuring model usability with label smoothing and loss function regularization.
+  • We propose the first MIA difficulty criterion in CI, identifying $I(x,z)$, $H(x|z)$, $H(z)$ and $\delta(z)$ as factors impacting attack effectiveness, deepening insights.
+  • We propose SiftFunnel, a privacy-preserving scheme for CI that constrains redundant features via linear and nonlinear correlation, ensuring model usability with label smoothing and loss function regularization.
   • SiftFunnel adopts a funnel-shaped structure with attention blocks to balance usability and privacy while ensuring the lightweight design of edge models.
-  • We conduct systematic experiments validating the proposed criterion and demonstrating the effectiveness of SiftFunnel in balancing usability, privacy, and deployability.
-2. Background and Related Work
-2.1. Collaborative inference
-Collaborative inference (Shlezinger et al., 2021) has emerged as an effective solution.
+  • We conduct systematic experiments to validate the proposed criterion and demonstrate SiftFunnel’s effectiveness in balancing usability, privacy, and deployability, using additional metrics including mutual information, $\delta(z)$, and parameter count for evaluation.
+2 Background and Related Work
+2.1 Collaborative inference
+Collaborative inference [3] has emerged as an effective solution.
 This approach divides the inference process into two stages: one part is performed locally on the edge device (which may consist of a single device or a chain of interconnected devices), while the remaining computation is uploaded to the cloud.
 Formally, this can be expressed as
-(1) $f(x)=f_{e}(x)\circ f_{c}(f_{e}(x))$
+$f(x)=f_{e}(x)\circ f_{c}(f_{e}(x))$ (1)
 where $f$ is the full trained model, $x$ is the input, $f_{e}$ represents the model deployed on the edge, and $f_{c}$ corresponds to the portion deployed in the cloud.
-Collaborative inference typically begins with training $f$ on centralized datasets in the cloud, then partitioning and deploying the $f_{e}$ to the edge.
-Alternatively, approaches like Split Learning  (Zhu et al., 2023) allow edge devices to participate in model training locally.
+CI typically begins with training $f$ on centralized datasets in the cloud, then partitioning and deploying the $f_{e}$ to the edge.
+Therefore, it is imperative to note that CI represents a scenario that has advanced to the inference application phase.
+Alternatively, approaches like Split Learning  [28] allow edge devices to participate in model training locally.
 Regardless of the training approach, the inference process remains consistent.
@@
 The final output is either returned to the edge device or retained in the cloud for further analysis.
-However, existing research (He et al., 2019) highlights that edge models are susceptible to MIAs, where adversaries reconstruct input data from transmitted features.
-Notably, in edge-based inference, computations can be distributed across multiple devices in a chained architecture (He et al., 2019), expressed as $f=f_{e1}\circ f_{e2}...\circ f_{c}$.
-Studies (He et al., 2020; Ho et al., 2024) show that deeper neural networks closer to the cloud transmit more complex, decision-focused features, making them more resistant to reconstruction.
+However, existing research [6] highlights that edge models are susceptible to MIAs, where adversaries reconstruct input data from transmitted features.
+Notably, in edge-based inference, computations can be distributed across multiple devices in a chained architecture [6], expressed as $f=f_{e1}\circ f_{e2}...\circ f_{c}$.
+Studies [16, 49] show that deeper neural networks closer to the cloud transmit more complex, decision-focused features, making them more resistant to reconstruction.
 In contrast, the initial edge device generates features retaining more input information, making it highly susceptible to MIAs.
@@
 This common scenario in real-world applications presents a significant security risk due to the increased likelihood of input data reconstruction, highlighting the need for effective countermeasures.
-2.2. Model inversion attack
-Current MIA techniques in CI can be broadly categorized into MLE-based MIAs and Gen-based MIAs, with related research progress described as follows.
+2.2 Model inversion attack
+Current MIAs in CI can be broadly categorized into MLE-based MIAs and Gen-based MIAs, with related research progress described as follows.
 MLE-based MIAs.
 MLE-based MIA was first introduced by Fredrikson et al.
-(Fredrikson et al., 2015), who reconstructed training data from a multilayer perceptron (MLP) by minimizing the difference between target and confidence scores using gradient descent.
-Subsequent research (He et al., 2019) incorporated Total Variation (TV) (Rudin et al., 1992) loss as a regularization term to improve optimization, with these advanced techniques referred to as rMLE-based MIA.
+[29], who reconstructed training data from a multilayer perceptron (MLP) by minimizing the difference between target and confidence scores using gradient descent.
+Subsequent research [6] incorporated Total Variation (TV) [30] loss as a regularization term to improve optimization, with these advanced techniques referred to as rMLE-based MIA.
 In this paper, both are collectively referred to as MLE-based MIA.
@@
 This approach was introduced by Yang et al.
-(Yang et al., 2019), who reconstructed input data in facial recognition systems by using generative deep learning models $g$ to approximate the inverse mapping function of the target model $f^{-1}$, such that $g\approx f^{-1}$.
+[10], who reconstructed input data in facial recognition systems by using generative deep learning models $g$ to approximate the inverse mapping function of the target model $f^{-1}$, such that $g\approx f^{-1}$.
 By feeding prediction vectors into $g$, they successfully generated facial input images.
 Building on this approach, He et al.
-(He et al., 2019) extended it to collaborative inference, demonstrating that input data from edge models is more vulnerable to reconstruction, thereby highlighting the risks.
+[6] extended it to collaborative inference, demonstrating that input data from edge models is more vulnerable to reconstruction, thereby highlighting the risks.
 Unlike MLE-based methods, Gen-based MIA does not require white-box access to the edge model.
@@
 Furthermore, as edge models produce rich and redundant outputs, this facilitates the training of attack models and the reconstruction process, as analyzed in detail in Section 4.
-2.3. Related defense work
+Facts and Motivation.
+Although MIA may be confused with “Membership Inference Attack” [9], its meaning is clearly defined in this work and widely adopted in the model inversion literature.
+Moreover, the feasibility of MIA and their threat to CI have been well established.
+While encryption may protect data in transit, it does not prevent input reconstruction once intermediate features are exposed.
+Such exposure can occur locally or via compromised cloud infrastructure.
+Moreover, encryption adds substantial overhead to edge systems, limiting its practicality.
+These concerns highlight the need to analyze attack factors and design lightweight, deployable defenses.
+2.3 Related defense work
 To safeguard input data collected by edge devices in CI from potential leakage, He et al.
-(He et al., 2019, 2020) emphasized the importance of developing effective defense strategies while identifying these vulnerabilities.
+[6, 16] emphasized the importance of developing effective defense strategies while identifying these vulnerabilities.
 Existing defenses can be classified into three main categories: perturbation-based, IB-based, and neural network depth (NND) & IB-based approaches.
@@
 He et al.
-(He et al., 2020) initially proposed using Gaussian noise and random feature dropout to limit the effective information accessible to attackers, and later, Wang et al.
-(Wang et al., 2022) extended this strategy by dynamically adjusting the depth of the edge model while adding differential privacy noise to transmitted features.
+[16] initially proposed using Gaussian noise and random feature dropout to limit the effective information accessible to attackers, and later, Wang et al.
+[17] extended this strategy by dynamically adjusting the depth of the edge model while adding differential privacy noise to transmitted features.
 Although these methods enhance defenses, they significantly impact model usability, typically reducing accuracy by approximately 10%, and remain less effective against Gen-based MIA.
-The root cause is that the decoder in Gen-based MIAs is sufficiently expressive to adapt to noise or feature dropout, similar to the training of a denoising auto-encoder (DAE) (Gondara, 2016), potentially increasing attack robustness.
+The root cause is that the decoder in Gen-based MIAs is sufficiently expressive to adapt to noise or feature dropout, similar to the training of a denoising auto-encoder (DAE) [32], potentially increasing attack robustness.
 In centralized MIA defenses, Struppek et al.
-(Struppek et al., 2023) proposed perturbing training labels with a negative label smoothing factor, significantly reducing MIA effectiveness and offering a novel perspective for designing defense strategies.
+[33] proposed perturbing training labels with a negative label smoothing factor, significantly reducing MIA effectiveness and offering a novel perspective for designing defense strategies.
 IB-based defense.
-This approach enhances defenses against inference attacks by limiting the mutual information between input data $x$ and intermediate features $z$ while preserving the mutual information between $z$ and outputs $y$, thus maintaining inference performance (Wang et al., 2021).
-The objective of this defense can be formally expressed as follows:
-(2) $\min_{\theta}-{I}(z,y)+{I}(x,z)$
+This approach enhances defenses against inference attacks by limiting the mutual information between input data $x$ and intermediate features $z$ while preserving the mutual information between $z$ and outputs $y$, thus maintaining inference performance [21].
+The objective of this defense can be formally expressed as $\min_{\theta}-{I}(z,y)+{I}(x,z)$.
 Initially, Wang et al.
-(Wang et al., 2021) introduced information bottleneck theory into MIA defenses, but their method was not optimized for CI and required substantial computational resources for intermediate feature processing.
+[21] introduced information bottleneck theory into MIA defenses, but their method was not optimized for CI and required substantial computational resources for intermediate feature processing.
 To address this, Peng et al.
-(Peng et al., 2022) proposed BiDO (Bilateral Dependency Optimization), which uses dependency algorithms to optimize correlations among inputs, features, and outputs.
+[18] proposed BiDO (Bilateral Dependency Optimization), which uses dependency algorithms to optimize correlations among inputs, features, and outputs.
 BiDO reduces input-feature correlation while enhancing feature-output correlation, achieving a balance between defense and performance.
 Additionally, Wang et al.
-(Wang et al., 2024) and Duan et al.
-(Duan et al., 2023) explored mutual information estimation between $x$ and $z$ using methods based on mutual information definition and the CLUB (Contrastive Log Upper Bound) (Cheng et al., 2020) technique, incorporating the estimates as constraints in the training loss function.
+[19] and Duan et al.
+[34] explored mutual information estimation between $x$ and $z$ using methods based on mutual information definition and the CLUB (Contrastive Log Upper Bound) [35] technique, incorporating the estimates as constraints in the training loss function.
 IB-based defenses provide interpretable protection mechanisms and valuable design insights for future research.
@@
 Ding et al.
-(Ding et al., 2024) proposed increasing network depth while compressing newly added hidden layers to restrict redundant information transmission.
+[20] proposed increasing network depth while compressing newly added hidden layers to restrict redundant information transmission.
 Subsequently, Azizian et al.
-(Azizian and Bajić, 2024) further improved defenses by introducing an auto-encoder (AE) structure, with encoders at the edge and decoders in the cloud.
+[15] further improved defenses by introducing an auto-encoder (AE) structure, with encoders at the edge and decoders in the cloud.
 The edge encoder increases network depth and reduces feature dimensionality to achieve the IB effect, while the cloud decoder restores information to maintain predictive performance while applying an $l_{1}$-norm constraint to reduce attack risks.
-Adversarial training (Ding et al., 2024) is also incorporated to strengthen robustness against MIAs.
+Adversarial training [20] is also incorporated to strengthen robustness against MIAs.
 However, this method has certain limitations.
 Increasing network depth and adding modules raise computational and hardware demands on edge devices, limiting applicability.
-Adversarial training, while strengthening defenses, introduces additional challenges—it is less effective against advanced attacks, complicates fine-tuning, and degrades inference performance (Zhang et al., 2019).
-Additionally, adversaries can adapt their attack models using similar adversarial strategies, weakening defense effectiveness (Jin et al., 2024).
+Adversarial training, while strengthening defenses, introduces additional challenges—it is less effective against advanced attacks, complicates fine-tuning, and degrades inference performance [36].
+Additionally, adversaries can adapt their attack models using similar adversarial strategies, weakening defense effectiveness [37].
 While these methods improve defenses, their implementation requires careful consideration of computational resources, model adaptability, and application suitability.
@@
 Thus, our method falls under the NND & IB-based defense strategy.
-3. Threat Model
-3.1. Knowledge background of adversary
+3 Threat Model
+3.1 Knowledge background of adversary
 Attack scenarios. Based on the adversary’s level of knowledge, attack scenarios can be categorized into the following three types:
-  1. (1)
+  1. 1.
     White-box Scenario: The adversary has complete white-box knowledge of the edge model, including its architecture, parameters, and gradient information.
-  2. (2)
+  2. 2.
     Black-box Scenario: The adversary has only black-box access to the model, restricting their capability to observe the intermediate features.
 Their understanding of the edge model’s architecture is also restricted to prior knowledge or incomplete assumptions.
-  3. (3)
+  3. 3.
     Gray-box Scenario: The adversary has partial knowledge of the edge model’s architecture but remains unaware of defensive modifications, such as increased hidden layers.
 The gray-box scenario enables the adversary to leverage their partial understanding of the edge model’s architecture to enhance the capability of their attack model.
-For instance, if the defender partitions the first three convolutional layers of a ResNet-18 model (He et al., 2016) and further deepens the network, while the adversary is only aware of the original ResNet-18 architecture, they can only construct an approximate attack model based on partial knowledge of the initial three layers.
+For instance, if the defender partitions the first three convolutional layers of a ResNet-18 model [38] and further deepens the network, while the adversary is only aware of the original ResNet-18 architecture, they can only construct an approximate attack model based on partial knowledge of the initial three layers.
 However, adversaries can address this limitation in this scenario by utilizing their understanding of the edge model’s architecture and reasoning about its mapping capabilities to refine their attack model.
@@
 Since segmentation, detection, and classification also rely on edge-processed features, MIA and defenses are task-agnostic.
-Prior work (Azizian and Bajić, 2024) has demonstrated input reconstruction and defense effectiveness in object detection using Gen-based MIA.
-Moreover, image data is particularly vulnerable to MIAs due to its high dimensionality, as shallow convolutional layers retain more redundant information (Zhang et al., 2020; He et al., 2016).
+Prior work [15] has demonstrated input reconstruction and defense effectiveness in object detection using Gen-based MIA.
+Moreover, image data is particularly vulnerable to MIAs due to its high dimensionality, as shallow convolutional layers retain more redundant information [11, 38].
 This increases the challenge of protecting image data, making it an ideal test case for privacy defenses.
-In contrast, textual data, with its strong contextual dependencies and syntactic complexity, has shown limited success in achieving high-quality MIA reconstructions (Dibbo, 2023; Zhang et al., 2022).
+In contrast, textual data, with its strong contextual dependencies and syntactic complexity, has shown limited success in achieving high-quality MIA reconstructions [39, 40].
 This study focuses on image data due to its interpretability, making it suitable for evaluating privacy defenses.
-3.2. Attack strategy of the adversary
-In the white-box attack scenario, the adversary can perform attacks using the MLE-based MIA method proposed in (He et al., 2019, 2020).
+3.2 Attack strategy of the adversary
+In the white-box attack scenario, the adversary can perform attacks using the MLE-based MIA method proposed in [6, 16].
 To ensure the experiments are both comparative and representative, hyper-parameter adjustments were made to achieve performance closer to optimal.
-In the black-box attack scenario, the adversary can employ methods from (Yang et al., 2019; He et al., 2019; Yin et al., 2023), relying solely on output features for reverse mapping without access to the edge model’s internal details.
+In the black-box attack scenario, the adversary can employ methods from [10, 6, 8], relying solely on output features for reverse mapping without access to the edge model’s internal details.
 When protected by NND & IB-based defenses, the adversary lacks knowledge of architectural changes and must design the attack model based on prior assumptions.
@@
 Detailed designs and training procedures for the attack models are provided in Section 6.
-4. Theoretical Analysis of MIA in CI
-4.1. Principles analysis of MIA
+4 Theoretical Analysis of MIA in CI
+4.1 Principles analysis of MIA
 MIAs are categorized into MLE-based and Gen-based methods, both aiming to reconstruct input data using the output of the edge model.
@@
 The attack principle is formulated as follows:
-(3) $\begin{gathered}x^{(k+1)}=x^{(k)}-\eta\nabla_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right),\\ x^{*}=\arg\min_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right)\end{gathered}$
+$\begin{gathered}x^{(k+1)}=x^{(k)}-\eta\nabla_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right),\\ x^{*}=\arg\min_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right)\end{gathered}$ (2)
 Where $x^{(k)}$ represents the input value at the $k$-th iteration, typically initialized as Gaussian noise, zeros, or ones.
@@
 $f_{\text{edge }}\left(\cdot\right)$ refers to the target edge model, $\nabla_{x}\mathcal{L}$ is the gradient calculated from the loss function and edge model parameters to optimize $x$, and $\eta$ is the step size for each update.
-According to (He et al., 2019), the loss function combines the $l_{2}$ distance between $f_{\text{edge }}\left(x^{(k)}\right)$ and $z$ with a TV regularization term for $x$, forming the basis of rMLE-based MIA.
+According to [6], the loss function combines the $l_{2}$ distance between $f_{\text{edge }}\left(x^{(k)}\right)$ and $z$ with a TV regularization term for $x$, forming the basis of rMLE-based MIA.
 The method relies on gradient descent, where the loss values depend on output features, and gradients are computed using the target model’s parameters.
 To achieve effective attacks, sufficient output feature information $\delta(z)$ and a well-chosen loss function are essential to ensure optimization space.
-Additionally, white-box knowledge of the edge model is required, as complex model mappings and limited feature information increase the risk of gradient descent converging to local minima (Struppek et al., 2022; Zhang et al., 2020).
+Additionally, white-box knowledge of the edge model is required, as complex model mappings and limited feature information increase the risk of gradient descent converging to local minima [53, 11].
 For instance, if two similar images produce nearly indistinguishable output features, reconstruction becomes challenging.
@@
 To address this, He et al.
-(He et al., 2019, 2020), building on the method from (Yang et al., 2019), developed a reverse mapping network for the edge model, which is categorized as Gen-based MIA.
+[6, 16], building on the method from [10], developed a reverse mapping network for the edge model, which is categorized as Gen-based MIA.
 The key idea is to use a generator to approximate the reverse mapping of the edge model.
@@
 For instance, Yang et al.
-(Yang et al., 2019) trained a decoder using auxiliary task-related datasets, while Yin et al.
-(Yin et al., 2023) utilized intercepted feature data for training.
+[10] trained a decoder using auxiliary task-related datasets, while Yin et al.
+[8] utilized intercepted feature data for training.
 The principle of this attack is summarized as follows.
-(4) $\begin{gathered}G=\arg\min_{G}\mathbb{E}_{x^{\prime}\sim p\left(x^{\prime}\right)}\mathcal{L}\left(G\left(f_{\text{edge }}\left(x^{\prime}\right)\right),x^{\prime}\right),\\ x^{*}=G(z)\end{gathered}$
+$\begin{gathered}G=\arg\min_{G}\mathbb{E}_{x^{\prime}\sim p\left(x^{\prime}\right)}\mathcal{L}\left(G\left(f_{\text{edge }}\left(x^{\prime}\right)\right),x^{\prime}\right),\\ x^{*}=G(z)\end{gathered}$ (3)
 Where $x^{\prime}$ represents samples from an auxiliary data distribution $p\left(x^{\prime}\right)$.
@@
 The above formula highlights the conditions required for effective attacks:
-  1. (1)
-    A well-matched training dataset aligned with the target task, ideally with a distribution similar to the target dataset (Yang et al., 2019).
-  2. (2)
+  1. 1.
+    A well-matched training dataset aligned with the target task, ideally with a distribution similar to the target dataset [10].
+  2. 2.
     Sufficient and highly separable input information for the generator.
@@
 Among them, the maximum separability is more decisive.
-For instance, even with random feature drops, if the remaining features distinguish inputs well, the attack can still reconstruct effectively (He et al., 2020).
-Such frameworks resemble DAE (Gondara, 2016), where the decoder can adapt to noise or feature drops.
+For instance, even with random feature drops, if the remaining features distinguish inputs well, the attack can still reconstruct effectively [16].
+Such frameworks resemble DAE [32], where the decoder can adapt to noise or feature drops.
 Gen-based MIA demonstrates superior reconstruction performance compared to MLE-based MIA in defense scenarios.
@@
 Furthermore, existing defenses struggle to effectively remove redundant feature information, allowing Gen-based MIA to overcome defenses if the generator is capable of learning the edge model’s reverse mapping.
-4.2. Difficulty criterion for MIA in CI
+4.2 Difficulty criterion for MIA in CI
 We summarized the factors influencing MIA effectiveness using a formulaic representation.
@@
 First, the mutual information between the input $x$ and the feature $z$, $I(x;z)$, can be expressed as:
-(5) $I(x;z)=H(x)-H(x\mid z)=H(z)-H(z\mid x)=H(z)$
+$I(x;z)=H(x)-H(x\mid z)=H(z)-H(z\mid x)=H(z)$ (4)
 Where $H(x\mid z)$ represents the conditional entropy of input $x$ given $z$, and $H(z\mid x)$ represents the reverse.
 Since there is no uncertainty in the mapping from $x$ to $z$ through $f_{\text{edge}}$, $H(z\mid x)$ is zero.
+This assumption holds primarily because the CI scenario involves deterministic inference, where model parameters are fixed and the feature extractor contains no stochastic components.
+Even if $x$ is perturbed by noise, the mapping $f_{\text{edge}}(x)$ still produces a unique and deterministic output $z$, ensuring that $H(z\mid x)=0$.
 Factors influencing MLE-based MIA.
@@
 This weakens attack effectiveness despite the apparent information sufficiency.
-Thus, MLE-based MIA difficulty is more influenced by $I(x;z)$, $H(x|z)$, and feature entropy $H(z)$.
-We can demonstrate the influence of mutual information and entropy on MLE-based MIA using the following relationship for any $x\to z\to x^{*}$, derived from Fano’s inequality (Arrow, 1969; Xu et al., 2025):
-(6) $H\left(P_{e}\right)+P_{e}\log|\mathbb{X}|\geq H(x\mid x^{*})\geq H(x\mid z)$
-where $P_{e}=Prob\{x\neq x^{*}\}$, $x^{*}$ is the reconstruction by MIA, and $\mathbb{X}$ denotes the value space of $x$.
-Since $\mathbb{X}\geq 2$ and $H\left(P_{e}\right)\leq 1$. Therefore, inequality (6) can be written as
-(7) $P_{e}\geq\frac{H(x\mid z)-1}{\log|\mathbb{X}|}=\frac{H(x)-I(x,z)-1}{\log|\mathbb{X}|}=\frac{H(x)-H(z)-1}{\log|\mathbb{X}|}$
-Since $H(x)$ and $|\mathbb{X}|$ are influenced by training data, they are typically assumed to be objective and invariant.
-Therefore, we observe that increasing the lower bound of $P_{e}$, and thus the difficulty of MIA, is influenced by $H(x\mid z)$, $H(z)$, and $I(x,z)$.
+Thus, MLE-based MIA difficulty is more influenced by $I(x;z)$, $H(x|z)$, and $H(z)$.
+To theoretically assess the difficulty of recovering the input $x$, we consider the reconstruction error probability $P_{e}$.
+In information theory, Fano’s inequality provides a fundamental lower bound on the reconstruction error in terms of conditional entropy.
+Modeling the MIA process as a Markov chain $x\to z\to x^{*}$, where $x^{*}$ is the reconstruction from $z$, Fano’s inequality [43, 42] impiles:
+$H\left(P_{e}\right)+P_{e}\log|\mathbb{X}|\geq H(x\mid x^{*})\geq H(x\mid z)$ (5)
+where $\mathbb{X}$ denotes the value space of $x$.
+Since $|\mathbb{X}|\geq 2$ and $H(P_{e})\leq 1$, we can derive a simplified lower bound:
+$P_{e}\geq\frac{H(x\mid z)-1}{\log|\mathbb{X}|}$ $=\frac{H(x)-I(x,z)-1}{\log|\mathbb{X}|}$ (6)
+$=\frac{H(x)-H(z)-1}{\log|\mathbb{X}|}$
+Here, $H(x)$ and $|\mathbb{X}|$ are considered intrinsic to the dataset and task, and typically remain fixed.
+Therefore, increasing the lower bound of $P_{e}$, and hence the difficulty of model inversion, is primarily influenced by $H(x\mid z)$, $H(z)$, and $I(x;z)$.
+While Fano’s inequality traditionally assumes exact recovery of $x$, it remains applicable under relaxed semantic criteria, provided that the error probability $P_{e}$ is redefined accordingly.
+The inequality merely requires a binary distinction between success and failure, allowing $P_{e}$ to be interpreted in terms of semantic equivalence or perceptual similarity in the context of model inversion attacks.
 Factors influencing Gen-based MIA.
@@
 A larger $\delta(z)$ not only facilitates generator training but also directly impacts the reconstruction of $x^{*}$.
-At the same time, the relationship in Inequality (7) also holds here (Zhang et al., 2023).
+At the same time, the relationship in Inequality (7) also holds here [31].
 Therefore, in addition to the factors influencing MLE-based MIA, Gen-based MIA is further effectively affected by $\delta(z)$.
@@
 Therefore, the difficulty criterion for MIA can be formulated as follows, where $k_{1}$ and $k_{2}$ are proportionality constants:
-(8) $D_{\text{mia}}\propto\left(\frac{k_{1}}{\{I(x;z);H(x\mid z);H(z)\}}+\frac{k_{2}}{\delta(z)}\right)$
-Based on this criterion, existing defenses (Wang et al., 2021; Peng et al., 2022; Duan et al., 2023) using the information bottleneck theory can be understood as approximating a reduction in $I(x;z)$ through various algorithms to increase the difficulty of MIA.
-Perturbation-based defenses (He et al., 2020) aim to directly increase $H(x\mid z)$.
-The defensive effect of deeper neural networks (Ding et al., 2024; Azizian and Bajić, 2024) can be explained by their ability to reduce $I(x;z)$ and $\delta(z)$.
-This is mainly because, without skip connections (Hao et al., 2025), mutual information progressively decreases during forward propagation in neural networks.
+$D_{\text{mia}}\propto\left(\frac{k_{1}}{\{I(x;z);H(x\mid z);H(z)\}}+\frac{k_{2}}{\delta(z)}\right)$ (7)
+Based on this criterion, existing defenses [21, 18, 34] using the information bottleneck theory can be understood as approximating a reduction in $I(x;z)$ through various algorithms to increase the difficulty of MIA.
+Perturbation-based defenses [16] aim to directly increase $H(x\mid z)$.
+The defensive effect of deeper neural networks [20, 15] can be explained by their ability to reduce $I(x;z)$ and $\delta(z)$.
+This is mainly because, without skip connections [44], mutual information progressively decreases during forward propagation in neural networks.
 In contrast, residual networks with skip connections preserve or even enhance $\delta(z)$ by adding layer outputs to inputs, but the decline in mutual information still poses challenges for MIA (see section 4.3 for experimental cases).
-4.3. Demonstration and Guidance for Defense
-To validate $D_{mia}$ and the theoretical analysis, we designed experiments using a CNN that progressively reduces the spatial size while increasing the channel size, based on the framework in (Yang et al., 2019; He et al., 2016), and ResNet-18 with skip connections, quantifying each block.
-We trained MINE (Belghazi et al., 2018) on the same dataset used to train the full model $f$ to estimate the lower bound of $I(x,z)$.
-The core idea of MINE is to optimize a neural network estimator using the Donsker-Varadhan (DV) representation (Donsker and Varadhan, 1983) as the loss function, providing an approximate estimate of the mutual information between random variables.
+4.3 Demonstration and Guidance for Defense
+To validate $D_{mia}$ and the theoretical analysis, we designed experiments using a CNN that progressively reduces the spatial size while increasing the channel size, based on the framework in [10, 38], and ResNet-18 with skip connections, quantifying each block.
+We trained MINE [24] on the same dataset used to train the full model $f$ to estimate the lower bound of $I(x,z)$.
+The core idea of MINE is to optimize a neural network estimator using the Donsker-Varadhan (DV) representation [45] as the loss function, providing an approximate estimate of the mutual information between random variables.
 Mutual information is defined as follows:
-(9) $I(X;Z)=\mathbb{E}_{p(x,z)}\left[\log\frac{p(x,z)}{p(x)p(z)}\right]$
+$I(X;Z)=\mathbb{E}_{p(x,z)}\left[\log\frac{p(x,z)}{p(x)p(z)}\right]$ (8)
 Where $p(x,z)$ represents the joint distribution, existing defense methods approximate it, but limited batch sizes lead to inaccuracies and variations in estimates.
@@
 To avoid direct distribution estimation, the DV representation reformulates mutual information as an optimization problem:
-(10) $I(X;Z)\geq\sup_{T}\mathbb{E}_{p(x,z)}[T(x,z)]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T(x,z)}\right]$
+$I(X;Z)\geq\sup_{T}\mathbb{E}_{p(x,z)}[T(x,z)]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T(x,z)}\right]$ (9)
 Where $T$ is a learnable function, this formulation enables $I(X;Z)$ to be estimated by optimizing $T$ to obtain a lower bound.
 MINE achieves this by parameterizing $T$ with a neural network $T_{\theta}$ and maximizing the DV bound through gradient descent to approximate mutual information:
-(11) $\hat{I}(X;Z)=\mathbb{E}_{p(x,z)}\left[T_{\theta}(x,z)\right]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T_{\theta}(x,z)}\right]$
-As shown in Figure 2, we trained the target models on the CIFAR-10 (Krizhevsky et al., 2009) and used Equation (11) to train MINE for the corresponding blocks.
+$\hat{I}(X;Z)=\mathbb{E}_{p(x,z)}\left[T_{\theta}(x,z)\right]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T_{\theta}(x,z)}\right]$ (10)
+As shown in Figure 2, we trained the target models on the CIFAR-10 [46] and used Equation (11) to train MINE for the corresponding blocks.
 Figures 2(a, b) show that increasing network depth reduces both mutual information and effective information, leading to a decrease in attack performance.
 Figures 2(c, d) illustrate that while skip connections within two layers do not reduce $\delta(z)$, they slightly lower $I(x,z)$, moderately affecting attack performance.
-Previous research (Hao et al., 2025) identified vulnerabilities caused by skip connections in centralized MIA.
+Previous research [44] identified vulnerabilities caused by skip connections in centralized MIA.
 Our study further highlights their impact on MIA effectiveness in CI.
-Figure 2. The impact of neural network depth on $I(x,z)$, $\delta(z)$, as well as MLE & Gen-based MIA.
-[…]
-2
+Figure 2: The impact of neural network depth on $I(x,z)$, $\delta(z)$, as well as MLE & Gen-based MIA.
 However, Transmission between blocks in ResNet alters feature channels and dimensions, requiring a shortcut layer, leading to a more pronounced decline in mutual information compared to direct stacking within layers.
 This, in turn, degrades MIA performance and aligns with our criterion.
-Figure 3. Overview of SiftFunnel. Specifically, the design is divided into two aspects: the architecture and the loss function for the edge and cloud model.
-[…]
-3
+Figure 3: Overview of SiftFunnel. Specifically, the design is divided into two aspects: the architecture and the loss function for the edge and cloud model.
 Notably, while mutual information at ResNet’s fourth layer is higher than that at CNN’s second layer, MLE-based MIA does not perform better.
 This stems from the fact that these methods rely heavily on gradient descent, making them sensitive to suboptimal optimizers, hyper-parameter settings, and deeper network structures, which hinder gradient calculations for optimizing $x$.
-The lack of flexible and efficient methods for MLE-based MIA remains a gap in current research, often leading to their exclusion in defense evaluations—a topic warranting further exploration (Ding et al., 2024).
+The lack of flexible and efficient methods for MLE-based MIA remains a gap in current research, often leading to their exclusion in defense evaluations—a topic warranting further exploration [20].
 In our experiment and Section 6, we still evaluate MLE-based MIA against various defenses and provide evidence demonstrating the effectiveness of our method in defending against such attacks.
@@
     For practical applications, the light-weighting of edge models should balance usability with their impact on the factors outlined in the criterion.
-5. SiftFunnel: Information Sifting Funnel
-5.1. Overview of SiftFunnel
+5 SiftFunnel: Information Sifting Funnel
+5.1 Overview of SiftFunnel
 Guided by $D_{{mia}}$, SiftFunnel balances usability and privacy by increasing uncertainty in inferring $x$ from $z$ while limiting redundant information and separability.
@@
 Attention mechanisms and multi-receptive field convolutions enhance task-relevant information, while a cloud-based upsampling module preserves accuracy.
-To reduce redundancy, $I(x,z)$ and increase the uncertainty of $z$ to $x$, we use Distance Correlation (Li et al., 2012) for nonlinear correlations and the Pearson Coefficient (Cohen et al., 2009) with $l_{1}$-norm for linear correlations.
-Additionally, Label Smoothing (LS) and Kullback-Leibler (KL) divergence (Donsker and Varadhan, 1983) are used during training to improve model usability while reducing the separability of representations in $z$.
-5.2. The Design of Shallow Edge Model
+To reduce redundancy, $I(x,z)$ and increase the uncertainty of $z$ to $x$, we use Distance Correlation [47] for nonlinear correlations and the Pearson Coefficient [48] with $l_{1}$-norm for linear correlations.
+Additionally, Label Smoothing (LS) and Kullback-Leibler (KL) divergence [45] are used during training to improve model usability while reducing the separability of representations in $z$.
+5.2 The Design of Shallow Edge Model
 Information Funnel.
 Deepening the neural network enhances feature extraction and reduces $I(x,z)$, as confirmed in Section 4.
-Previous work (Ho et al., 2024) shows through Fisher information (Rissanen, 1996) that deeper networks improve task relevance and resist MIA.
+Previous work [49] shows through Fisher information [50] that deeper networks improve task relevance and resist MIA.
 Inspired by this, we redesign the final layer of the edge model with an Information Funnel module.
@@
 The Information Funnel’s channel structure controls output but does not limit redundancy or enhance task-relevant information transmission, affecting usability.
-Inspired by Inception networks (Szegedy et al., 2017), we redesign it as a multi-receptive field convolutional framework with three low-channel layers.
+Inspired by Inception networks [51], we redesign it as a multi-receptive field convolutional framework with three low-channel layers.
 A $1*1$ kernel fuses channel information, followed by $3*3$ and $5*5$ kernels to expand the receptive field, extract richer task-relevant features, improve decision-making, and reduce mutual information with $x$, while lowering parameters of the edge model and controlling $\delta(z)$.
@@
 To enhance the output of task-relevant information from the Information Funnel, attention modules are incorporated both before and after it.
-The first block combines an SE block (Hu et al., 2018) and Triplet Attention (Misra et al., 2021).
+The first block combines an SE block [25] and Triplet Attention [26].
 The SE block evaluates the contribution of information within each channel, suppressing channels with low task relevance.
 Triplet Attention reduces spatial information while maintaining smooth, continuous outputs, preserving fine-grained features for further processing.
-The second block integrates SE and CBAM modules (Woo et al., 2018).
+The second block integrates SE and CBAM modules [27].
 Compared to Triplet Attention, CBAM is a lightweight module that sequentially calculates channel and spatial weights to enhance or suppress input features.
@@
 This sparsity is beneficial for output processing, effectively increasing $H(x\mid z)$ while reducing $H(z)$ and $\delta(z)$.
-5.3. The Design of Loss Function
+5.3 The Design of Loss Function
 To further enhance model usability and data privacy, appropriately designed loss constraints must be applied to $x$ and $z$ to reasonably increase $H(x\mid z)$ and suppress $H(z)$ within the edge model structure.
@@
 The key challenge is how to achieve this effectively and flexibly.
-Some defense approaches utilize Constrained Covariance (COCO) (Gretton and Herbrich, 2005) to estimate correlation, focusing primarily on the linear relationship between $x$ and $z$.
-However, since neural networks predominantly perform nonlinear operations, studies have shown that this constraint is not only ineffective but may also lead to reduced usability (Peng et al., 2022).
-Alternatively, some approaches employ the Hilbert-Schmidt Independency Criterion (HSIC) (Gretton et al., 2005) to estimate nonlinear correlation.
+Some defense approaches utilize Constrained Covariance (COCO) [52] to estimate correlation, focusing primarily on the linear relationship between $x$ and $z$.
+However, since neural networks predominantly perform nonlinear operations, studies have shown that this constraint is not only ineffective but may also lead to reduced usability [18].
+Alternatively, some approaches employ the Hilbert-Schmidt Independency Criterion (HSIC) [54] to estimate nonlinear correlation.
 However, kernel-based calculations in HSIC rely heavily on the appropriate selection of $\sigma$.
@@
 The calculation of distance correlation is as follows:
-(12) $\mathcal{L}_{\text{d Cor}}=\mathrm{d}\operatorname{Cor}(x,z)=\frac{\mathrm{d}\operatorname{Cov}(x,z)}{\sqrt{\mathrm{d}\operatorname{Var}(x)\cdot\mathrm{d}\operatorname{Var}(z)}}$
+$\mathcal{L}_{\text{d Cor}}=\mathrm{d}\operatorname{Cor}(x,z)=\frac{\mathrm{d}\operatorname{Cov}(x,z)}{\sqrt{\mathrm{d}\operatorname{Var}(x)\cdot\mathrm{d}\operatorname{Var}(z)}}$ (11)
 Where $\mathrm{d}\operatorname{Cov}(x,z)$ represents the distance covariance between $x$ and $z$, and the denominator is the product of the distance variances for the two samples.
@@
 To remove the influence of sample shifts, the distance matrix is centered as $A_{ij}^{c}=A_{ij}-\overline{A_{i}}-\overline{A_{j}}+\overline{A_{i}}$, where $\overline{A_{i}}$ denotes the mean of the $i$-th column, and similar operations apply to other terms.
-With the centered distance matrix, the distance covariance is computed as $\mathrm{d}\operatorname{Cov}(x,z)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c}B_{ij}^{c}$, where $B$ represents the matrix of features $z$, and $n$ denotes the batch size.
+With the centered distance matrix, the distance covariance is computed as $\mathrm{d}\operatorname{Cov}(x,z)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c}B_{ij}^{c}$, where $B$ represents the matrix of features $z$, and $n$ is the batch size.
 The distance variance is defined as $\mathrm{d}\operatorname{Var}(x)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c^{2}}$, which normalizes the distance covariance.
@@
 The calculation is as follows:
-(13) $\mathcal{L}_{\text{Pearson }}=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\operatorname{pCor}_{c_{1},c_{2}}\right)^{2}=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\frac{\operatorname{Cov}_{c_{1},c_{2}}}{\sigma_{c_{1}}\cdot\sigma_{c_{2}}}\right)^{2}$
+$\mathcal{L}_{\text{Pearson }}$ $=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\operatorname{pCor}_{c_{1},c_{2}}\right)^{2}$ (12)
+$=\frac{1}{C^{2}}\sum_{c_{1}=1}^{C}\sum_{c_{2}=1}^{C}\left(\frac{\operatorname{Cov}_{c_{1},c_{2}}}{\sigma_{c_{1}}\cdot\sigma_{c_{2}}}\right)^{2}$
 Where $\operatorname{Cov}_{c_{1},c_{2}}$ represents the covariance matrix between channels $c_{1}$ and $c_{2}$, calculated as $\operatorname{Cov}_{c_{1},c_{2}}=\frac{1}{\mathrm{H}\cdot\mathrm{W}}\sum_{\mathrm{i}=1}^{\mathrm{H}\cdot\mathrm{W}}\left(z_{c_{1},i}-\mu_{c_{1}}\right)\cdot\left(z_{c_{2},i}-\mu_{c_{2}}\right)$, where $z_{c_{1},i}$ is the $i$-th element of channel $c_{1}$ and $\mu_{c_{1}}$ is its mean.
@@
 Label smoothing.
-Finally, to ensure model usability, we draw inspiration from (Struppek et al., 2023) and apply LS to refine the target loss, as shown in formula (14).
+Finally, to ensure model usability, we draw inspiration from [33] and apply LS to refine the target loss, as shown in formula (14).
 Additionally, KL divergence is used instead of cross-entropy (CE) as the primary loss function for optimization.
-(14) $LS\left(\hat{y}_{c}\right)=1-\alpha+\frac{\alpha}{K},\quad LS\left(\hat{y}_{i}\right)=\frac{\alpha}{K}$
+$LS\left(\hat{y}_{c}\right)=1-\alpha+\frac{\alpha}{K},\quad LS\left(\hat{y}_{i}\right)=\frac{\alpha}{K}$ (13)
 Where $LS\left(\hat{y}_{c}\right)$ refers to the label smoothing element under the target class, $LS\left(\hat{y}_{i}\right)$ is the i-th element of the non-target class.
 $\alpha$ is smooth factor and $K$ is the total number of categories.
-CE enhances inter-class feature distinctions, while KL divergence supports smoother and more generalized feature extraction between classes (Struppek et al., 2023).
-Within-class feature extraction captures universal characteristics, ensuring generalization and robustness for the target task.
-This optimization reduces feature separability, increasing the difficulty of attack, as demonstrated in Section 6.3.
+While the $z$ at this layer have not fully transitioned into decision-level semantics, the combination of architectural constraints and the implicit regularization of $I(x,z)$ imposed by the loss function encourages more task-relevant representations.
+Consequently, positive-factor LS can induce slight intra-class compactness, modestly reducing the feature entropy and hindering inversion attacks [33], as evidenced in Section 6.3.
+Even in cases where LS increases entropy due to insufficient decision-oriented expression, its impact on attack vulnerability remains limited.
+Ultimately, label smoothing is intended to promote generalization, with improved resistance to MIAs appearing only as a byproduct when sufficient constraints are imposed on the feature representations.
 In summary, the proposed loss function first applies nonlinear correlation constraints to increase $H(x|z)$ and suppress $H(z)$.
@@
 SiftFunnel and MINE are available at https://github.com/SiftFunnel/SiftFunnel.
-(15) $\min\mathcal{L}_{\theta}(x,z,\hat{y})=$
+$\min\mathcal{L}_{\theta}(x,z,\hat{y})=$ (14)
 $\lambda_{1}\cdot KL(f(x),LS(\hat{y}))+\lambda_{2}\cdot\mathcal{L}_{\text{d Cor}}+\lambda_{3}\cdot\mathcal{L}_{\text{Pearson}}+\tau\cdot\|z\|_{1}$
-6. Experiments
-In this section, we systematically evaluate and compare the ability of SiftFunnel and existing approaches to resist MIA. The specific experimental setup and evaluation metrics are shown below.
-6.1. Experimental Setup
-1) Datasets. We used three types of image recognition datasets, with processing details outlined as follows. The detailed data allocation is shown in Table 1.
-  • CIFAR-10 (Krizhevsky et al., 2009).
+6 Experiments
+6.1 Experimental Setup
+1) Datasets. We used three types of image recognition datasets, with processing details outlined as follows. The detailed data allocation is shown in Table I.
+  • CIFAR-10 [46].
 The CIFAR-10 dataset consists of 60,000 RGB images categorized into ten classes with an original resolution of $32\times 32$.
 To make the results clearer and more intuitive, we used an unscaled resolution of $64\times 64$.
-  • FaceScrub (Ng and Winkler, 2014).
+  • FaceScrub [55].
 FaceScrub is a URL dataset with 100,000 images of 530 actors, which contains 265 male actors and 265 female actors.
 However, since not every URL was available during the writing period, we downloaded a total of 43,149 images for 530 individuals and resized the images to $64\times 64$.
-  • CelebA (Liu et al., 2015).
+  • CelebA [56].
 CelebA is a dataset with 202,599 images of 10,177 celebrities.
-We used the same crop as (Yang et al., 2019) to remove the background of images in this dataset other than faces to reduce the impact on the experiment.
+We used the same crop as [10] to remove the background of images in this dataset other than faces to reduce the impact on the experiment.
 To eliminate individual overlap, we removed a total of 6,878 images of 296 individuals and similarly resized the images to $64\times 64$.
-  • ChestX-ray (Sait et al., 2020).
+  • ChestX-ray [57].
 This dataset is a curated collection of COVID-19 chest X-ray images compiled from 15 publicly available datasets.
 It contains 1,281 COVID-19 X-rays, 3,270 normal X-rays, 1,656 viral pneumonia X-rays, and 3,001 bacterial pneumonia X-rays with a resolution of $128\times 128$.
-Table 1. Data allocation of the target model and attack model.
+TABLE I: Data allocation of the target model and attack model.
 Target model Attack Model
@@
 ChestX-ray (4 classes) Train: 66.6%, Test: 16.7% 16.7% (4 classes of ChestX-ray)
-Table 2.
-Quantitative evaluation of CNN on CIFAR-10 with different defenses against two MIAs.
+TABLE II: Quantitative evaluation of CNN on CIFAR-10 with different defenses against two MIAs.
 ↑ indicates higher is better, ↓ indicates lower is better.
@@
 Blue highlights show unprotected baseline performance, red indicates increased edge device load, yellow indicates reduced load, and gray shows gray-box defense performance.
-MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
+MLE-based Attack [6] Gen-based Attack [16]
 Method Class Method Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
 Unprotected 88.28% 0.0045 28.3421 0.9883 0.0017 32.5498 0.9987 0.0566 32,549 299,520
-Adding noise (He et al., 2020) ($\sigma$ = 0.8)
+Adding noise [16] ($\sigma$ = 0.8)
 74.85% (-13.43%) 0.0291 20.1511 0.9135 0.0172 22.4249 0.9299 0.0502 65,535
-Dropping (He et al., 2020) (r = 0.8)
+Dropping [16] (r = 0.8)
 21.18% (-67.10%) 0.0127 23.7340 0.9609 0.0135 23.4835 0.9481 0.0501 6,509
-DPSGD (Abadi et al., 2016) (20, 1e-5)
+DPSGD [59] (20, 1e-5)
 61.83% (-26.45%) 0.0218 21.3868 0.9479 0.0016 32.4800 0.994 0.0623 39,829
 Perturbation-based
-NLS (Struppek et al., 2023) ($\alpha$ = -0.05)
-87.60% (-0.68%)
-0.0002 42.2684 0.9996 0.0014 33.3732 0.9949 0.0573 47,183 299,520
-MID (Wang et al., 2021) (1e-2) 81.53% (-6.75%) 1.74e-5 50.4675 0.9999 0.0013 34.5184 0.9967 0.0584 45,198
-BiDO (Peng et al., 2022) (2; 20) 84.12% (-4.16%) 0.6699 6.5111 0.0138 0.035 19.2965 0.8751 0.0429 3,345
-VIB (Wang et al., 2024) 88.21% (-0.07%) 0.5552 7.3264 0.0358 0.0042 28.5054 0.9842 0.0551 18,069
-IB-based InfoSCISSORS (Duan et al., 2023) 83.46% (-4.83%) 0.5715 7.2007 0.1345 0.0062 26.8799 0.9948 0.0549 43,364 299,520
+NLS [33] ($\alpha$ = -0.05)
+85.53% (-2.75%)
+0.0002 42.2684 0.9996 0.0015 33.0003 0.9943 0.0573 47,183 299,520
+MID [21] (1e-2) 81.53% (-6.75%) 1.74e-5 50.4675 0.9999 0.0013 34.5184 0.9967 0.0584 45,198
+BiDO [18] (2; 20) 84.12% (-4.16%) 0.6699 6.5111 0.0138 0.035 19.2965 0.8751 0.0429 3,345
+VIB [19] 88.21% (-0.07%) 0.5552 7.3264 0.0358 0.0042 28.5054 0.9842 0.0551 18,069
+IB-based InfoSCISSORS [34] 83.46% (-4.83%) 0.5715 7.2007 0.1345 0.0062 26.8799 0.9948 0.0549 43,364 299,520
 0.0542 17.4852 0.8151
-AE-based (Azizian and Bajić, 2024) 90.56% (+2.28%) 0.6345 6.7472 0.0051 0.01926 21.9259 0.9295 0.0358 2,617 2,365,056
+AE-based [15] 90.56% (+2.28%) 0.6345 6.7472 0.0051 0.01926 21.9259 0.9295 0.0358 2,617 2,365,056
 0.1166 14.1112 0.5419
 NND & IB-based SiftFunnel 85.49% (-2.79%) 0.6792 6.4510 0.0051 0.0639 16.7314 0.7647 0.0167 1,006 14,911
-Figure 4.
-A visual evaluation of the attack effectiveness of two MIAs on CNN edge models protected by various defense techniques, with NND & IB-based results specifically reflecting the gray-box scenario.
-[…]
-4
+Figure 4: A visual evaluation of the attack effectiveness of two MIAs on CNN edge models protected by various defense techniques, with NND & IB-based results specifically reflecting the gray-box scenario.
 2) Attack Method.
@@
 3) Target Edge Model and Implementation Details.
-We employ the same comprehensive model architecture as detailed in (Yang et al., 2019), comprising four CNN blocks succeeded by two fully connected layers.
+We employ the same comprehensive model architecture as detailed in [10], comprising four CNN blocks succeeded by two fully connected layers.
 This model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ parameter of 0.5, and a ReduceLROnPlateaud scheduler with a reduction factor of 0.5 and patience of 25.
@@
 For the training of the edge model protected by SiftFunnel, we have specified the parameters $\lambda_{1}$ to 3.5, $\alpha$ to 0.35, $\lambda_{2}$ to 0.8, and $\lambda_{3}$ to 0.6 in Equation (15).
-In addition, to argue the impact of the model architecture for the defense, we use different CNNs as follows: (1) VGG16 adapted from (Simonyan, 2014); (2) ResNet-18 adapted from (He et al., 2016), the learning rates are 3e-4.
+In addition, to argue the impact of the model architecture for the defense, we use different CNNs as follows: (1) VGG16 adapted from [58]; (2) ResNet-18 adapted from [38], the learning rates are 3e-4.
 The division positions of ResNet and VGG16 are not fixed.
@@
 We aim to compare three existing categories of defense methods.
-For Perturbation-based defenses, we select Adding noise & Dropping features (He et al., 2020), DPSGD (Abadi et al., 2016), and negative label smoothing (NLS) (Struppek et al., 2023).
-For IB-based defenses, we choose MID (Wang et al., 2021), BiDO (with HSIC) (Peng et al., 2022), VIB (Wang et al., 2024), and InfoSCISSORS (Duan et al., 2023).
-For NND & IB-based defenses, we opt for AE-based protection (Azizian and Bajić, 2024).
+For Perturbation-based defenses, we select Adding noise & Dropping features [16], DPSGD [59], and negative label smoothing (NLS) [33].
+For IB-based defenses, we choose MID [21], BiDO (with HSIC) [18], VIB [19], and InfoSCISSORS [34].
+For NND & IB-based defenses, we opt for AE-based protection [15].
 Our method falls into the NND & IB-based defense category, which also requires analyzing the defensive effect in a gray-box scenario.
@@
 Lastly, all methods are trained without adversarial training, focusing solely on the effects of the original defense methods.
-6.2. Evaluation Metrics
+6.2 Evaluation Metrics
 We selected 7 evaluation metrics to assess the aspects of usability, privacy, and deployability.
 For usability assessment, we employ Test Accuracy (Test ACC), which evaluates the final performance of the model protected by defense techniques on the test dataset.
-Regarding privacy assessment, we initially utilize MSE, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) to evaluate the effectiveness of MIA on reconstructing the input data of the protected model (Yin et al., 2023).
+Regarding privacy assessment, we initially utilize Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) to evaluate the effectiveness of MIA on reconstructing the input data of the protected model [8].
 Here, a higher MSE and lower PSNR and SSIM (ranging from 0 to 1) indicate better defense effectiveness.
-In addition, we employ MINE (Belghazi et al., 2018), as mentioned in Section  4.3, to assess Mutual Information (MI) (ranging from 0 to 1) and the Effective Information Mean $\delta(z)$, which is the mean of the non-zero output by the edge model on the test dataset.
+In addition, we employ MINE [24], as mentioned in Section  4.3, to assess Mutual Information (MI) (ranging from 0 to 1) and the Effective Information Mean $\delta(z)$, which is the mean of the non-zero output by the edge model on the test dataset.
 The specific training methods and model architecture for MINE are referenced from https://github.com/SiftFunnel/SiftFunnel.
 For deployability, we consider the Edge model’s parameter count $|\theta_{\text{edge}}|$ and Inference Latency.
-The model parameter count directly determines memory or GPU memory occupation, while inference latency is tested on both CPU and GPU.
-Table 3.
-Quantitative evaluation of ResNet-18 on CIFAR-10 with different defenses against two MIAs.
+Model parameters directly affect memory usage, and inference latency is measured on both CPU and GPU.
+TABLE III: Quantitative evaluation of ResNet-18 on CIFAR-10 with different defenses against two MIAs.
 The target edge model is the first residual block of ResNet-18.
@@
 Bolded values are the best.
-The meanings of color-coded regions are consistent with those in Table 2.
-MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
+The meanings of color-coded regions are consistent with those in Table II.
+MLE-based Attack [6] Gen-based Attack [16]
 Method Class Method Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
 Unprotected 90.43% 0.0001 71.0996 0.9991 0.0010 34.6599 0.9991 0.9719 182,134 149,824
-Adding noise (He et al., 2020) ($\sigma$ = 2)
+Adding noise [16] ($\sigma$ = 2)
 81.39% (-9.04%)
 0.0035 29.3354 0.9872 0.0054 27.4167 0.9794 0.9619 262,144
-Dropping (He et al., 2020) (r = 0.8)
+Dropping [16] (r = 0.8)
 14.09% (-76.34%)
 0.0069 26.3641 0.9754 0.0077 25.9334 0.9711 0.9514 35,574
-DPSGD (Abadi et al., 2016) (20, 1e-5)
+DPSGD [59] (20, 1e-5)
 61.85% (-28.58%)
-0.1901 11.9821 0.7172 0.0060 26.9751 0.9802 0.9515 178,956
+0.0115 23.9453 0.9628 0.0060 26.9751 0.9802 0.9515 178,956
 Perturbation-based
-NLS (Struppek et al., 2023) ($\alpha$ = -0.05)
-88.57% (-1.86%)
-0.0034 29.5001 0.9894 0.0011 34.4231 0.9961 0.9718 176,456 149,824
-MID (Wang et al., 2021) (1e-2)
+NLS [33] ($\alpha$ = -0.05)
+88.13% (-2.30%)
+0.0034 29.5001 0.9894 0.0008 36.0306 0.9974 0.9721 176,456 149,824
+MID [21] (1e-2)
 79.48% (-10.95%)
 1.42e-7 73.2595 0.9999 0.0010 34.5893 0.9965 0.9729 173,587
-BiDO (Peng et al., 2022) (5; 10)
+BiDO [18] (5; 10)
 90.08% (-0.35%)
 4.38e-7 68.3603 0.9999 0.0020 31.7621 0.9927 0.9647 219,832
-VIB (Wang et al., 2024)
+VIB [19]
 89.18% (-1.25%)
 0.5869 7.0853 0.0216 0.0020 31.7621 0.9928 0.9402 54,800
-IB-based InfoSCISSORS (Duan et al., 2023)
+IB-based InfoSCISSORS [34]
 88.03% (-2.40%)
@@
 0.0143 23.2332 0.9488
-AE-based (Azizian and Bajić, 2024) 89.19% (-1.24%) 0.6543 6.6189 0.1243 0.0094 25.0633 0.9660 0.9397 74,356 279,136
+AE-based [15] 89.19% (-1.24%) 0.6543 6.6189 0.1243 0.0094 25.0633 0.9660 0.9397 74,356 279,136
 0.1957 11.8562 0.0793
 NND & IB-based SiftFunnel 88.43% (-2.00%) 0.6538 6.6166 0.0059 0.2790 10.3230 0.3001 0.4772 4,028 78,864
-Table 4.
-Quantitative evaluation of different defense methods in protecting CNN against MIAs on the FaceScrub dataset.
+TABLE IV: Quantitative evaluation of different defense methods in protecting CNN against MIAs on the FaceScrub dataset.
 Changes in data types alter the inference task, requiring adjustments to the parameters of the defense methods.
-MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
+MLE-based Attack [6] Gen-based Attack [16]
 Method Class Method Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
@@
 Perturbation-based
-NLS (Struppek et al., 2023) ($\alpha$ = -0.2)
-87.97% (-0.34%) 0.0004 38.8904 0.9980 0.0018 32.3274 0.9906 0.0496 44,783 299,520
-MID (Wang et al., 2021) (1e-2) 79.45% (-10.98%) 0.0017 32.3435 0.9897 0.0018 32.1285 0.9901 0.0507 39,277
-BiDO (Peng et al., 2022) (2; 20) 88.02% (-0.29%) 0.0004 39.0838 0.9978 0.0018 32.2048 0.9904 0.0493 44,915
-VIB (Wang et al., 2024) 88.05% (-0.26%) 0.7354 6.1060 0.0032 0.0033 29.5697 0.9823 0.0493 15,649
-IB-based InfoSCISSORS (Duan et al., 2023) 81.61% (-6.70%) 0.7449 6.0504 0.0026 0.0060 26.9803 0.9675 0.0408 57,050 299,520
+NLS [33] ($\alpha$ = -0.2)
+88.03% (-0.28%) 0.0004 38.8904 0.9980 0.0017 32.5480 0.9911 0.0515 44,783 299,520
+MID [21] (1e-2) 79.45% (-10.98%) 0.0017 32.3435 0.9897 0.0018 32.1285 0.9901 0.0507 39,277
+BiDO [18] (2; 20) 88.02% (-0.29%) 0.0004 39.0838 0.9978 0.0018 32.2048 0.9904 0.0493 44,915
+VIB [19] 88.05% (-0.26%) 0.7354 6.1060 0.0032 0.0033 29.5697 0.9823 0.0493 15,649
+IB-based InfoSCISSORS [34] 81.61% (-6.70%) 0.7449 6.0504 0.0026 0.0060 26.9803 0.9675 0.0408 57,050 299,520
 0.0609 16.9260 0.6656
-AE-based (Azizian and Bajić, 2024) 83.45 (-4.86%) 0.6474 6.6594 0.0090 0.0220 21.3444 0.8795 0.0238 585 2,365,056
+AE-based [15] 83.45 (-4.86%) 0.6474 6.6594 0.0090 0.0220 21.3444 0.8795 0.0238 585 2,365,056
 0.0559 17.3018 0.6854
 NND & IB-based SiftFunnel ($\lambda_{1}=5.0,\lambda_{2}=0.1,\lambda_{3}=0.1$) 85.39% (-2.92%) 0.7264 6.1593 0.0060 0.0296 20.0576 0.8323 0.0104 637 14,911
-6.3. Experimental Results
+6.3 Experimental Results
 The results are analyzed from 4 perspectives: the impact of defense methods on $I(x,z)$ and $\delta(z)$, the influence of datasets and model architectures, and ablation studies.
 1) The effect of various defenses on the MI and $\delta(z)$.
-As shown in Table 2, defense effectiveness is reflected by MI and $\delta(z)$, aligning with the criterion $D_{\text{mia}}$.
+As shown in Table II, defense effectiveness is reflected by MI and $\delta(z)$, aligning with the criterion $D_{\text{mia}}$.
 Specifically, unprotected edge models are highly vulnerable to MIAs, with MI at 0.0566, MSE below 0.01, and SSIM exceeding 0.98.
 Perturbation-based defenses reduce MI by at most 11% but cause over a 10% drop in Test ACC, with NLS even increasing MI in shallow networks, weakening resistance to attacks.
-UIB-based defenses slightly reduce usability but effectively counter MLE-based MIAs, except for MID, which shows weaker resistance.
+From the theoretical analysis, it can be seen that DPSGD does not alter the complexity of the mapping from $x$ to $z$, nor does it directly affect the intermediate feature $z$ itself.
+As a result, it reduces model accuracy without effectively mitigating MIA risks in the CI setting.
+IB-based defenses slightly reduce usability but effectively counter MLE-based MIAs, except for MID, which shows weaker resistance.
 Among them, BiDO significantly impacts MI and $\delta(z)$, providing effective defense, but its parameter selection and kernel width configuration (as shown in the table) limits generalizability, as confirmed in subsequent experiments.
@@
 While CLUB effectively defends against MLE-based MIAs, it struggles against Gen-based MIAs and incurs a 4.83% accuracy loss.
-AE-based defenses improve model usability, likely due to enhanced model capacity, while MI reductions confirm reasonable defense effectiveness.
+AE-based defenses improve model usability, due to enhanced model capacity, while MI reductions confirm reasonable defense effectiveness.
 However, gray-box scenarios weaken AE-based defenses, and their added complexity increases storage costs by 8×8$\times$8 ×, as seen in $|\theta_{\text{edge}}|$.
@@
 Additionally, $|\theta_{\text{edge}}|$ reveals a nearly 20×20$\times$20 × reduction in edge model load, highlighting SiftFunnel’s efficiency.
-Table 5.
-SiftFunnel Ablation Study on CIFAR-10.
+TABLE V: SiftFunnel Ablation Study on CIFAR-10.
 The target edge model is the first residual block of ResNet-18, with Gen-based MIA as the attack method.
@@
 This setup poses challenges for defense techniques, as shallow networks retain simple structures and do not modify spatial dimensions.
-Table 3 shows that unprotected models exhibit minimal MI impact, resulting in weak resistance to MIAs, with MLE-based MIA achieving an MSE of $10^{-4}$ and Gen-based MIA $10^{-3}$.
+Table III shows that unprotected models exhibit minimal MI impact, resulting in weak resistance to MIAs, with MLE-based MIA achieving an MSE of $10^{-4}$ and Gen-based MIA $10^{-3}$.
 Under these conditions, perturbation-based defenses fail to balance usability and privacy, leading to poor defense performance.
 Among IB-based defenses, only VIB demonstrates resistance to MLE-based MIA by optimizing the edge model through $I(x,z)$ estimation as a loss function, but its effectiveness remains limited, particularly against Gen-based MIA.
-For BiDO, despite applying kernel bandwidth estimation methods from (Peng et al., 2022) and extensive experimental tuning, it fails to achieve a configuration that effectively balances usability and privacy, resulting in suboptimal defense performance.
+For BiDO, despite applying kernel bandwidth estimation methods from [18] and extensive experimental tuning, it fails to achieve a configuration that effectively balances usability and privacy, resulting in suboptimal defense performance.
 InfoSCISSORS impacts MI under this configuration, but the increase in $\delta(z)$ undermines its effectiveness against Gen-based MIAs, leading to only marginal improvements.
@@
 Consequently, AE-based defenses struggle to counter both black-box and gray-box Gen-based MIAs effectively.
-Table 6.
-Quantitative evaluation of the resistance to MIA across different architectures on CIFAR-10.
-The edge model divisions for CNN and ResNet-18 follow the configurations in Tables 2 and 3, while for VGG16, the edge model is set at the first ReLU layer after a max-pooling operation.
-MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
+TABLE VI: Quantitative evaluation of MIA resistance on CIFAR-10 across different architectures.
+Edge models for CNN and ResNet-18 follow Tables II and III; for VGG16, the split is placed at the first ReLU after max pooling.
+MLE-based Attack [6] Gen-based Attack [16]
 Architecture Test ACC↑ MSE↑ PSNR↓ SSIM↓ MSE↑ PSNR↓ SSIM↓ MI↓ $\delta(z)$ $|\theta_{\text{edge}}|$
@@
 VGG16 88.89% (-2.18%) 0.7801 5.8450 0.0005 0.0833 15.5916 0.7181 0.0128 3,587 52,086
-In contrast, SiftFunnel reduces MI by 50% and lowers $\delta(z)$ to the millesimal scale, delivering the strongest resistance against MIAs while preserving usability and deployability.
-Notably, it achieves a minimum SSIM of 0.0793, underscoring its superior ability to balance privacy and usability.
-The different protection performances of the three model architectures with the same dataset are shown in Table 6.
+In contrast, SiftFunnel reduces MI by 50% and lowers $\delta(z)$ to the millesimal scale, offering the strongest MIA resistance while maintaining usability.
+It achieves a minimum SSIM of 0.0793, highlighting its privacy–utility balance.
+Table VI compares performance across architectures.
 3) The effect of data type.
 Changes in data types lead to shifts in task requirements.
-For instance, in this study, the facial recognition task involves 530 classes, while ChestX-ray classification involves 4 classes, necessitating parameter adjustments as shown in Table 4.
+For instance, in this study, the facial recognition task involves 530 classes, while ChestX-ray classification involves 4 classes, necessitating parameter adjustments as shown in Table IV.
 As analyzed in the previous section on model architectures, to achieve better defense performance, we adjusted the edge model divisions for ResNet-18 and VGG.
 Specifically, the division points were set after the second residual block for ResNet and after the first ReLU layer following the second max-pooling layer for VGG.
-The results, summarized in Table 4 and Figure 5, show that the defense performance remains unaffected by changes in data types.
+The results, summarized in Table IV and Figure 5, show that the defense performance remains unaffected by changes in data types.
 SiftFunnel continues to be the best defense method, offering an optimal balance between usability, privacy, and deployability.
-As shown in Figure 5, the VGG16 model protected by SiftFunnel demonstrates better resistance to Gen-based MIA in terms of the visual effects of usability and privacy on facial and X-ray data, with PSNR values consistently around 15.
+As shown in Figure 5, the VGG16 model with SiftFunnel shows better resistance to Gen-based MIA in terms of the visual effects of usability and privacy on facial and X-ray data, with PSNR values consistently around 15.
 4) Ablation Study.
 To assess the effectiveness of SiftFunnel’s model structure, loss function design, and defense capability, we conducted ablation experiments using Gen-based MIA.
-As shown in Table 5, we separately removed the Information Funnel and Attention Blocks from the model structure.
+As shown in Table V, we separately removed the Information Funnel and Attention Blocks from the model structure.
 The results indicate that removing the Funnel had minimal impact on accuracy, reducing it by only 0.1%, whereas removing the attention mechanism significantly increased MI, weakening privacy protection.
@@
 Additionally, we validated the hypothesis from Section 5.2, demonstrating that even when adversaries reconstructed inputs using cloud-compensated features, SiftFunnel still outperformed existing defenses in protecting transmitted features.
-Figure 5.
-Quantitative and visual evaluation of privacy and usability against Gen-based MIA in the VGG16 edge model on the ChestX-ray ($\lambda_{1}=3.5,\lambda_{2}=1.2,\lambda_{3}=0.8$) and FaceScrub ($\lambda_{1}=4.5,\lambda_{2}=0.4,\lambda_{3}=0.1$).
-[…]
-5
-We also evaluated the processing speed of the edge model on both CPU and GPU under this configuration.
-For single-sample input, the unprotected edge model had a CPU latency of 1.204 ms, the AE-based protected model had 2.812 ms, and SiftFunnel achieved 2.192 ms.
-On GPU, the processing latency was 3.9 ms, 4.3 ms, and 6.9 ms, respectively.
-While SiftFunnel showed slightly higher latency on GPU, it required significantly less memory, enabling more parallel processing of inputs under the same resource constraints.
-7. Conclusion
+Figure 5: Quantitative and visual evaluation of privacy and usability against Gen-based MIA in the VGG16 edge model on the ChestX-ray ($\lambda_{1}=3.5,\lambda_{2}=1.2,\lambda_{3}=0.8$) and FaceScrub ($\lambda_{1}=4.5,\lambda_{2}=0.4,\lambda_{3}=0.1$).
+We also evaluated the edge model’s latency on both CPU and GPU under the same configuration.
+For single-sample input, the unprotected, AE-based, and SiftFunnel-protected models had CPU latencies of 1.204 ms, 2.812 ms, and 2.192 ms, respectively.
+On GPU, their latencies were 3.9 ms, 4.3 ms, and 6.9 ms.
+Although SiftFunnel incurred slightly higher latency on GPU, it consumed significantly less memory, allowing more parallel input processing within the same resource constraints.
+7 Conclusion
 While CI enhances efficiency, it also exposes transmitted features to MIAs, posing severe privacy risks.